### v2

# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Variable Coefficients

Abstract
We consider the adaptive-rank integration of 2D and 3D time-dependent advection-diffusion partial differential equations (PDEs) with variable coefficients. We employ a standard finite-difference method for spatial discretization coupled with diagonally implicit Runge-Kutta temporal schemes. The discrete equation is a generalized Sylvester equation (GSE), which we solve with an adaptive-rank algorithm structured around three key strategies: (i) constructing dimension-wise subspaces based on an extended Krylov strategy, (ii) developing an effective preconditioner for the reduced coefficient matrix, and (iii) efficiently computing the residual of the equation without explicitly reverting to the full-rank form. The low-rank decomposition is performed in 2D using SVD, and with high-order SVD (HOSVD) in 3D to represent the tensor in a compressed Tucker format. The computational complexity of the proposed approach is demonstrated numerically to be comparable to the constant-coefficient case [14], scaling as $\mathcal{O}(N{r^{2}}+{r^{d+1}})$ for $d$-dimensional problems (here, $d=2$ or $3$), with $N$ the resolution in one dimension and $r$ the maximal rank during the Krylov iteration (which we find to be largely independent of $N$). We present numerical examples that illustrate the computational efficacy and complexity of our algorithm.
Keywords: Adaptive-rank, extended Krylov subspaces, implicit Runge-Kutta integrators, advection-diffusion equations, Sylvester preconditioning, matrix equations.
1 Introduction:
We propose an adaptive-rank numerical scheme for approximating solutions of time-dependent variable-coefficient advection-diffusion equations. The goal is to reduce the computational complexity of high-dimensional PDE solvers by exploiting the low-rank structure of the solution, building upon existing literature on low-rank PDE solvers [16, 13] and model order reduction (MOR) for solving high-dimensional linear and nonlinear systems [24, 16, 5, 40].
Two primary strategies in the literature exploit evolving low-rank structures for time-dependent PDEs: the dynamical low-rank (DLR) approximation [21, 27, 8, 11, 26] and the step-and-truncate (SAT) method, which has both explicit [23, 10, 18, 17, 19] and implicit [35, 30, 14, 28, 29] variants. Our study employs SAT, discretizing implicitly in time with high-order accuracy, leading to matrix/tensor differential equations to be solved at every time step. For constant coefficients in 2D, the advection-diffusion PDE becomes a Sylvester equation (SE) of the form:
$A\mathbf{X}_{1}+\mathbf{X}_{1}B^{\top}=\mathbf{X}_{0},$ (1)
where $\mathbf{X}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ represents the solution at time $t_{1}$, $\mathbf{X}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$ is the initial condition at $t_{0}$, and $A$ and $B$ are difference operators arising from the PDE discretization. Here, $N_{1}$ and $N_{2}$ are the 2D grid dimensions.
Krylov subspace methods have proven effective for solving SEs resulting from the discretization of PDEs [24, 39, 35, 31, 6]. In these methods, the solution bases are chosen from a Krylov (or Krylov-like) subspace of $A$ and $B$ in (1), and the original system is Galerkin-projected onto this reduced subspace. To accelerate the convergence of iterative methods, extensions of Krylov subspaces have been developed [12] involving $A^{-1}$, $B^{-1}$, which have been successfully applied to matrix equations [39, 14].
While substantial progress has been made for constant-coefficient problems, PDEs with variable coefficients remain challenging. Discretizing such systems leads to the so-called generalized Sylvester equation (GSE), of the form:
$\sum_{i=1}^{l}A_{i}\mathbf{X}_{1}{B}^{\top}_{i}=\mathbf{X}_{0},$ (2)
where $A_{i}$ and $B_{i}$ represent discretization operators for advection and/or diffusion terms. Several methods have been proposed to solve GSEs either with stationary iterative techniques  [5, 38] or Krylov methods  [5, 39, 32, 42]. Stationary iterative methods leverage low-rank techniques to enhance efficiency, achieving computational complexity of $\mathcal{O}(N)$ when implemented effectively [38]. However, their effectiveness critically depends on whether a dominating operator can be correctly identified to enable a contractive iterative process. For the Krylov approach, recent approaches leverage a Galerkin projection of the matrix equation on the dimension-wise Krylov bases when the right-hand side is low-rank [36, 40]. Compared to a standard Kronecker Krylov iteration, dimension-wise Krylov subspaces are inherently richer and yield superior approximations [32, 42, 24].
Several studies in the literature have considered using atypical extended Krylov subspaces to solve GSEs (see [5, 33, 20] and references therein). In [5], the authors present a series of techniques to tackle the Generalized Lyapunov equation, which is a special case of (2) of the form:
$A\mathbf{X}+\mathbf{X}A^{T}+\sum_{j=1}^{m}M_{j}\mathbf{X}M_{j}^{T}=\mathbf{F}.$ (3)
Among them, they discuss a low-rank projection-based scheme, in which extended Krylov subspaces are constructed using $A$ and standard Krylov subspaces are used for the remaining operators. In [20], a Krylov-based projection technique is developed for Generalized Lyapunov equations with $m=1$. The method builds its approximation subspaces by block augmentations generated from $A$ and $M_{1}$. In [33], a low-rank projection-based technique is employed for systems of disparate dimension sizes, $N_{2}\ll N_{1}$. The authors employ a nested sequence of block rational Krylov subspaces generated from all the operators in the $N_{1}$ direction, and solve the resulting small-sized equation with either matrix-based Conjugate Gradient or a Kronecker formulation (tractable since $N_{2}$ is small). In all of these studies, atypical approximation subspaces are formed by augmenting the basis with multiple applications of the operators in the multi‑term matrix equation, a strategy we adopt in this work.
Another Krylov-based line of inquiry, particularly relevant to our study, employs a basis-update-and-Galerkin (BUG) DLR integrator as a low-rank GMRES preconditioner for the GSE arising from anisotropic diffusion time-dependent problems with variable coefficients [28]. The BUG preconditioner evolves the basis and the coefficient matrix to provide guesses for the basis vectors and coefficient updates of the outer GMRES iteration, demonstrating significant potential in controlling GMRES iteration count and maximal Krylov rank. The approach is compatible with high-order temporal integrators. However, the reference offers no discussion of asymptotic computational complexity.
In this study, we propose to solve the GSE for time-dependent advection-diffusion equations leveraging adaptive-rank approximations of solutions on tensor-product grids with $N$ grid points per dimension, and using extended Krylov subspaces to update the basis vectors as well as the rank (denoted as $r$). Our main goal is to minimize the computational cost of the adaptive rank implementation without sacrificing accuracy. Key innovations of our proposed algorithm include:
  • Generalized Extended Krylov (xKrylov) Adaptive-Rank Integrators with Complexity Reduction for 2D and 3D Problems. We consider matrix solutions with factorizations of the form $X=USV^{\top}$, where basis vectors in $U$ and $V$ span the solution space and determine the rank of the solution. In the 2D setting, we employ a truncated singular value decomposition (SVD) to factorize the solution. In 3D, we use high-order singular value decomposition (HOSVD) to represent the 3D tensor in a compressed Tucker format. Our proposed approach constructs a set of xKrylov subspaces (for each dimension) that effectively capture the evolution of solution basis in an iterative manner. A Galerkin projection is then applied to project the full system to a reduced GSE. We propose a residual-checking mechanism that leverages the low-rank structure of solutions directly, thus avoiding forming the full residual matrix for norm evaluations. For the rest of the paper, we refer to the xKrylov iteration as the outer iteration.
  • Effective Preconditioning Strategy: The effectiveness of these projection-based approaches depends on the availability of an effective solver for the reduced GSE (2) of size up to $r$, the maximum size of dimension-wise xKrylov subspaces. There are two main challenges to solve it: (i) an increased computational load due to the multi-term nature of the equation, leading to rapid increase of xKrylov subspace sizes, and (ii) slow convergence of iterative methods (e.g., GMRES) resulting from algebraic ill-conditioning when solved iteratively. These challenges highlight the need for an effective preconditioner. Inspired by [31], we propose an averaged-coefficient Sylvester (ACS) approximation operator to the GSE as a preconditioner to GMRES. The approximate operator is of the Sylvester form and compact-equivalent to the original GSE, ensuring excellent performance as a GMRES accelerator [1, 2]. The direct inversion of the SE features a favorable computational complexity of $\mathcal{O}(r^{d+1})$, to be contrasted with the $\mathcal{O}(r^{3d})$ for $d$-dimensional problems when directly inverting the Kronecker-product form of the GSE. We refer to the preconditioned-GMRES reduced-GSE solve as the inner iteration.
  • Efficient Computational Complexity: As proposed, our method achieves a computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$ for $d$-dimensional problems, offering a significant advantage over existing state-of-the-art adaptive-rank approaches. The $Nr^{2}$ term corresponds to the outer xKrylov iteration, and the $r^{d+1}$ term to the inner GMRES iteration, assuming an $r$-independent GMRES convergence rate (which we achieve with the ACS preconditioner). In principle, the rank $r$ may scale with $N$ [24], but we do not find this in practice. The favorable scaling properties of our approach will be verified in our numerical experiments.
The rest of this paper is organized as follows. In Section 2, we consider the case of rank-one advection and diffusion coefficients and derive the GSE that results from a finite-difference discretization in space and backward Euler in time. In Section 3, we introduce an adaptive-rank solution strategy for 2D problems, including the coefficient-averaged approximation operator, construction of xKrylov subspaces, preconditioning technique in solving the reduced system, and low-rank computation of the residual norm. The extension of the technique to high-order temporal discretizations with multi-rank coefficients is provided in Appendix C, and for 3D problems in Appendix D. In Section 4, numerical results are presented to demonstrate the efficacy of the proposed algorithms. Finally, we conclude and provide outlook for future research in Section 5.
2 Problem Description and Generalized Sylvester Equation
For simplicity, we focus this section on 2D, with the extension to 3D elaborated in Appendix D. We establish the following notation conventions. Lowercase Greek letters, such as $\phi(x,y)$ or $\sigma(x,y)$, denote continuous scalar fields. Uppercase bold italic Greek letters, such as $\bm{\mathit{\Sigma}}(x,y)$ or $\bm{\mathit{\Phi}}(x,y)$, are used for vector-valued and tensor-valued functions. Uppercase Greek letters, like $\Phi$ or $\Sigma$, represent discrete matrices arising from the discretization of continuous scalar functions or differential operators on a grid. Boldface uppercase letters, such as $\mathbf{F}$ or $\mathbf{S}$, represent unknown matrix solutions or matrix right-hand sides in matrix equations. The identity matrix of size $n$ is denoted by ${I}_{n}$.
We aim to solve the time-dependent advection-diffusion equation:
$\frac{\partial f}{\partial t}=\nabla_{v}\cdot(\bm{\mathit{\Phi}}\cdot\nabla f)-\nabla\cdot(\bm{\mathit{\Sigma}}f),$ (4)
where $(x,y)\in\Omega=[a_{1},b_{1}]\times[a_{2},b_{2}]$ and $\bm{\mathit{\Phi}}\in\mathbb{R}^{2\times 2}$ is an isotropic diffusion tensor:
$\bm{\mathit{\Phi}}(x,y)=\begin{bmatrix}\phi^{x}(x,y)&0\\ 0&\phi^{y}(x,y)\end{bmatrix},$
and $\bm{\mathit{\Sigma}}\in\mathbb{R}^{2}$ is the advection vector, given by:
$\bm{\mathit{\Sigma}}(x,y)=\begin{bmatrix}\sigma^{x}(x,y)\\ \sigma^{y}(x,y)\end{bmatrix}.$
We assume that the diffusion and advection coefficients are separable with finite rank. Specifically, the diagonal diffusion coefficients, $\phi^{x}(x,y)$ and $\phi^{y}(x,y)$, are given by:
$\phi^{x}(x,y)=\sum_{i=1}^{\ell_{x}}\phi^{1,x}_{i}(x)\phi^{2,x}_{i}(y),\quad\phi^{y}(x,y)=\sum_{i=1}^{\ell_{y}}\phi^{1,y}_{i}(x)\phi^{2,y}_{i}(y).$ (5)
Similarly, for the advection coefficients we have:
$\sigma^{x}(x,y)=\sum_{i=1}^{k_{x}}\sigma^{1,x}_{i}(x)\sigma^{2,x}_{i}(y),\quad\sigma^{y}(x,y)=\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\sigma^{2,y}_{i}(y).$ (6)
Substituting these coefficients in their separable forms into Equation (4), we obtain:
$\frac{\partial f}{\partial t}$ $=\sum_{i=1}^{\ell_{x}}\frac{\partial}{\partial x}\left(\phi^{1,x}_{i}(x)\frac{\partial f(x,y)}{\partial x}\right)\phi^{2,x}_{i}(y)+\sum_{i=1}^{\ell_{y}}\phi^{1,y}_{i}(x)\frac{\partial}{\partial y}\left(\phi^{2,y}_{i}(y)\frac{\partial f(x,y)}{\partial y}\right)$ (7)
$-\sum_{i=1}^{k_{x}}\frac{\partial}{\partial x}\left(\sigma^{1,x}_{i}(x)f(x,y)\right)\sigma^{2,x}_{i}(y)-\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\frac{\partial}{\partial y}\left(\sigma^{2,y}_{i}(y)f(x,y)\right),$
where, in each term, the factor independent of the differentiation variable has been pulled out of the derivative. To simplify the presentation, we assume for now rank-one advection and diffusion coefficients:
$\phi^{x}=\phi^{1,x}(x)\phi^{2,x}(y),\quad\phi^{y}=\phi^{1,y}(x)\phi^{2,y}(y),$
$\sigma^{x}=\sigma^{1,x}(x)\sigma^{2,x}(y),\quad\sigma^{y}=\sigma^{1,y}(x)\sigma^{2,y}(y).$
This will be generalized to multiple ranks below. Equation (7) then simplifies to:
$\frac{\partial f}{\partial t}$ $={\frac{\partial}{\partial x}\left(\phi^{1,x}(x)\frac{\partial f(x,y)}{\partial x}\right)\phi^{2,x}(y)+\phi^{1,y}(x)\frac{\partial}{\partial y}\left(\phi^{2,y}(y)\frac{\partial f(x,y)}{\partial y}\right)}$ (8)
${-}{\frac{\partial}{\partial x}\left(\sigma^{1,x}(x)f(x,y)\right)\sigma^{2,x}(y){-}\sigma^{1,y}(x)\frac{\partial}{\partial y}\left(\sigma^{2,y}(y)f(x,y)\right)}.$
We discretize this equation in a two-dimensional tensor-product spatial grid with ${N}_{1}$ and ${N}_{2}$ grid points in the $x$ and $y$ directions, respectively. The grid points in the $x$ direction are denoted as $x_{i}$ for $i=0,\dots,N_{1}-1$, and similarly for the $y$ direction. We consider the point-wise multiplication of the advection and diffusion coefficients in the discrete by setting:
$\Phi_{1}$ $=\text{diag}(\phi^{1,y}(x_{i})),\Sigma_{1}=\text{diag}(\sigma^{1,y}(x_{i})),\quad i=0,\cdots,N_{1}-1,$
$\Phi_{2}$ $=\text{diag}(\phi^{2,x}(y_{j})),\Sigma_{2}=\text{diag}(\sigma^{2,x}(y_{j})),\quad j=0,\cdots,N_{2}-1.$
We approximate the spatial derivative with a second-order finite-difference scheme using a three-point stencil, and obtain the following matrix differential equation for $\mathbf{F}$, for which its $(i,j)$ entry $F_{i,j}$ corresponds to the approximation of $f(x_{i},y_{j})$:
$\frac{\partial\mathbf{F}}{\partial t}={D_{xx}^{\Phi}\mathbf{F}\Phi_{2}^{\top}+\Phi_{1}\mathbf{F}{D_{yy}^{\Phi}}^{\top}}-{D_{x}^{\Sigma}\mathbf{F}\Sigma_{2}^{\top}-\Sigma_{1}\mathbf{F}{D_{y}^{\Sigma}}^{\top}}.$ (9)
Here, $D_{xx}^{\Phi},D_{yy}^{\Phi},D_{x}^{\Sigma},D_{y}^{\Sigma}$ represent compositions of difference operators with pointwise multiplication operators. See Appendices A and B for more details. For notational simplicity, let:
$T_{1}=D_{xx}^{\Phi},\quad T_{2}=D_{yy}^{\Phi},\quad T_{3}=-D_{x}^{\Sigma},\quad T_{4}=-D_{y}^{\Sigma}.$ (10)
Equation (9) can be written compactly as:
$\frac{\partial\mathbf{F}}{\partial t}=\mathscr{L}(\mathbf{F}),$ (11)
where the right-hand side induces the following mapping:
${\mathscr{L}}:\mathbf{F}\mapsto T_{1}\mathbf{F}\Phi_{2}^{\top}+\Phi_{1}\mathbf{F}T_{2}^{\top}+T_{3}\mathbf{F}\Sigma_{2}^{\top}+\Sigma_{1}\mathbf{F}T_{4}^{\top}.$ (12)
We consider first a backward Euler implicit treatment of the temporal component of the matrix differential equation (11), which gives the following discrete GSE:
$\mathbf{F}_{1}-\Delta t\mathscr{L}({\bf F}_{1})=\mathbf{F}_{0},$ (13)
and associated residual equation:
$\mathbf{R_{\mathscr{L}}}=\mathbf{F}_{1}-\Delta t\,\mathscr{L}(\mathbf{F}_{1})-\mathbf{F}_{0}.$ (14)
Iterative solvers update the solution $\mathbf{F}_{1}$ until the residual is sufficiently small under a suitable norm. In this work, we opt for Galerkin-projection-based iterative solvers, where the bases spanning $\mathbf{F}_{1}$ are updated by augmenting Krylov-like subspaces, as shown next.
3 Krylov-Based Adaptive-Rank Solver for Generalized Sylvester Equation
We propose an adaptive-rank algorithm for the GSE (13). We assume a low-rank factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ at time $t^{(1)}=t^{(0)}+\Delta t$ also in the low-rank format, i.e.:
[…]
where $U_{0}\in\mathbb{R}^{N_{1}\times r_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r_{1}}$, and $V_{1}\in\mathbb{R}^{N_{2}\times r_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$ and $S_{1}\in\mathbb{R}^{r_{1}\times r_{1}}$ are diagonal matrices with singular values at the corresponding times. Projection methods involve constructing bases $U_{1}$ and $V_{1}$ and then projecting the original equation onto them to determine the coefficients of $S_{1}$. A key aspect of the proposed adaptive-rank approach is that we never form the full solutions explicitly. Instead, we seek bases in a dimension-by-dimension manner and then use them to perform a Galerkin projection. Key questions for solving the GSE that we address in this study include:
  • How do we choose suitable subspaces effectively to build $U_{1}$ and $V_{1}$ from $U_{0}$ and $V_{0}$?
  • How can we solve the reduced equation for $S_{1}$ efficiently?
  • How can we check the residual of the GSE without forming the matrices explicitly?
Our main innovation leverages an approximate operator $\tilde{\mathscr{L}}$ (which is of the Sylvester form) to the original operator ${\mathscr{L}}$ in (12). This approximation not only allows us to construct dimension-wise Krylov bases $U_{1}$ and $V_{1}$ effectively and efficiently, but also enables effective preconditioning for solving the projected reduced system for $S_{1}$, in the spirit of [31]. In the following, we cover the construction of the approximated operator $\tilde{\mathscr{L}}$ in Section 3.1; we discuss the dimension-wise basis construction in Section 3.2; we present the reduced equation for the matrix of coefficients $S_{1}$ and introduce our ACS preconditioner in Section 3.3, and outline the residual computation in Section 3.4. Finally, Section 3.5 extends the algorithm to the case of multi-rank coefficients.
3.1 Averaged-Coefficient Approximate Operator $\tilde{\mathscr{L}}$
Our approach incorporates ideas from [31] and [14] to construct an approximate operator $\tilde{\mathscr{L}}$ and efficiently solve the time-dependent inhomogeneous advection-diffusion equation using a projection-based adaptive-rank integration framework. The approximated operator $\tilde{\mathscr{L}}$ serves two purposes: (i) constructing the solution bases $U_{1}$ and $V_{1}$ as xKrylov subspaces, and (ii) preconditioning the solution of the coefficient matrix $S_{1}$.
We seek an approximation to ${\mathscr{L}}$ that captures its spectral content and is easier to invert. As proposed in [31], the advection diagonal operators can be approximated by scaling an identity matrix, where the scaling factor is the average of the diagonal entries. In this work, we approximate not only the advection coefficients by a suitable constant multiple of the identity, but extend this strategy to the scalar diffusion coefficients by averaging them across the computational domain. Let $\Phi_{i}\approx{\alpha}_{i}I$ and $\Sigma_{i}\approx\gamma_{i}I$, where $\alpha_{i}$ and $\gamma_{i}$ are the averages of the diagonal entries of $\Phi_{i}$ and $\Sigma_{i}$, respectively. This approximation yields the following mapping:
~${\tilde{\mathscr{L}}}:\mathbf{F}\mapsto\left(\alpha_{2}T_{1}+\gamma_{2}T_{3}\right)\mathbf{F}+\mathbf{F}\left(\alpha_{1}T_{2}+\gamma_{1}T_{4}\right)^{\top}.$ (16)
A backward Euler temporal discretization using (16) leads to the SE:
$(A_{1}+A_{3})\mathbf{F}_{1}+\mathbf{F}_{1}(A_{2}+A_{4})^{\top}=\mathbf{F}_{0},$ (17)
with:
$A_{1}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{2}T_{1}\right)\quad A_{3}=\left(\frac{1}{4}I-\Delta t\,\gamma_{2}T_{3}\right),\quad P_{1}=A_{1}+A_{3},$ (18)
$A_{2}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{1}T_{2}\right)\quad A_{4}=\left(\frac{1}{4}I-\Delta t\,\gamma_{1}T_{4}\right),\quad P_{2}=A_{2}+A_{4}.$ (19)
The solution to (17) is an approximation to that of (13), but it can be efficiently computed using low-rank approximations [14, 39, 37, 36]. It can therefore be exploited as a viable preconditioner for a GMRES iterative solution of the GSE (13). This was done in [31] with $\mathcal{O}(N^{3})$ computational complexity. In this work, we extend this approach by performing a Galerkin projection of both the original operator $\mathscr{L}$ and its approximation $\tilde{\mathscr{L}}$ on the low-rank bases before solving the GSE with GMRES. This reduces the problem to a lower-dimensional subspace, where preconditioning can be applied more effectively on a smaller $d$-dimensional problem with $\mathcal{O}(r^{d+1})$ complexity, resulting in an overall computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$.
3.2 Basis Construction
We begin by defining the xKrylov subspaces used for our basis construction at the $m$-th outer iteration as:
$\kappa_{m}(M_{1},\cdots,M_{l},U)=\text{{\color[rgb]{0,0,0}Range}}\Big{\{}U,\underbrace{M_{1}U,\cdots,M_{l}U}_{\text{1st augmentation}},\ldots,\underbrace{M_{1}^{m+1}U,\cdots,M_{l}^{m+1}U}_{\text{mth augmentation}}\Big{\}}.$ (20)
where $M_{1},\cdots,M_{l}$ are operators of interest and their inverses arising in our GSE, and $l$ is the number of operators used for each basis augmentation step.
The xKrylov basis construction involves several operators from the original GSE (13) and its approximation (17). Following [14], we construct an xKrylov subspace in the $x$-direction for the approximated SE (17) using $P_{1}$ and its inverse. We also include the inverses of the individual operators $A_{1}$ and $A_{3}$, as they may contain information not captured by the inverse of $P_{1}$. Additionally, we consider the action of the diagonal operators $\Phi_{1}$, $\Sigma_{1}$, $\Phi_{2}$, and $\Sigma_{2}$, which are averaged to obtain $\tilde{\mathscr{L}}$. We have found that including all these matrices is effective in capturing the range of the solution. Using these operators, we construct a generalized Krylov subspace in the $x$-direction as follows:
$\kappa_{m}\left(M_{1},\cdots,M_{6},U_{0}\right),\quad\text{where}$ $M_{1}=P_{1},$ $M_{2}=P_{1}^{-1},$ $M_{3}=A_{1}^{-1},$ $M_{4}=A_{3}^{-1},$ $M_{5}=\Phi_{1},$ $M_{6}=\Sigma_{1}.$
Depending on the structure of the GSE (2), $l$ may be large, and the generated subspace vectors at each augmentation may exhibit linear dependencies. To limit subspace inflation, we employ an SVD-based truncation strategy (Algorithm 1), such that only basis vectors corresponding to significant singular values are retained.
The construction of the dimension-wise xKrylov subspaces (20), following [39], is as follows. At each iteration $m$, Algorithm 2 augments the previously constructed basis $\mathcal{V}_{(m-1)}$ by applying powers of operators $M_{1}$ through $M_{l}$ to their corresponding basis blocks $V^{(1)},\dots,V^{(l)}$. Each block matrix $V^{(i)}$ may have a rank of at most $r_{0}$. We apply Algorithm 2 to orthogonalize the subspace and obtain a basis $\mathcal{U}_{(m)}\in\mathbb{R}^{N_{1}\times r^{(m)}_{x}}$ and $\mathcal{V}_{(m)}\in\mathbb{R}^{N_{2}\times r^{(m)}_{y}}$ in the $x$ and $y$ directions, respectively. For notational simplicity, we omit the calligraphic notation and subscript $(m)$ from now on, understanding that $U_{1}\in\mathbb{R}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{R}^{N_{2}\times r_{y}}$ refer to the bases of the Krylov subspaces at the $m$th iteration. The truncation step in line 2 of Algorithm 2 removes any linear dependencies within each block during the power iterations; line 2 applies a global truncation across all newly generated blocks, eliminating redundancies among them. As a result, the basis of the subspace is kept optimal in size.
Input: Krylov Subspace $\kappa_{m}$; Tolerance $\epsilon_{\kappa}$.
Output: Truncated bases ${Q}$.
1 Compute the reduced QR decomposition: $\{Q,R\}=\texttt{QR}\left(\kappa_{m}\right)$;
2 Perform reduced SVD decomposition: $Z_{1}\tilde{R}Z_{2}=\texttt{SVD}(R)$, where $\tilde{R}=\text{diag}(\sigma_{j})$ Identify the last index $r$ such that $\sigma_{r}>\epsilon_{\kappa}$;
3 Update the basis vectors: $Q\leftarrow QZ_{1}(:,1:r)$;
Algorithm 1 SVD-truncated QR, $\mathscr{T}_{\epsilon_{\kappa}}$
Input: Operators $M_{1},\dots,M_{l}$; augmentation basis blocks $\{V^{(1)},\dots,V^{(l)}\}$; previous Krylov basis ${\mathcal{V}}_{(m-1)}$; tolerance $\epsilon_{\kappa}$.
Output: Updated Krylov basis $\mathcal{V}_{(m)}$; basis blocks $\{V^{(1)},\dots,V^{(l)}\}$ for the next iteration.
1
2for $i=1,\dots,l$ do
3        $V^{(i)}\leftarrow M_{i}V^{(i)}$;
4        $\widetilde{V}^{(i)}\leftarrow\text{orthogonalize }V^{(i)}\text{ w.r.t. }V_{(m-1)}$ (e.g., modified Gram-Schmidt);
        $V^{(i)}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}(\widetilde{V}^{(i)})$ ;
         // truncate to $\epsilon_{\kappa}$ for next iteration using Algorithm 1
5
[…]
$V^{\prime}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}(V^{\prime})$ ;
  // final orthogonalization and truncation using Algorithm 1
7 $\mathcal{V}_{(m)}\leftarrow[\mathcal{V}_{(m-1)},V^{\prime}]$;
Algorithm 2 Construction of xKrylov Subspace $\kappa_{m}^{\texttt{trunc}}$
We comment next on the performance of the xKrylov iteration. In principle, the number of xKrylov iterations may depend on the mesh size, $N$. For instance, [24, 4] theoretically show for the 2D and 3D Poisson equations that the dimension of the extended Krylov subspace (denoted by $r^{(m)}$ in the basis construction above) scales as $\mathcal{O}(\kappa^{1/4})\sim\mathcal{O}(N^{1/2})$, with $\kappa\sim N^{2}$ the condition number. While this provides an upper bound of the convergence rate (without considering solution smoothness), our numerical results suggest that it may be pessimistic in practice. In particular, for moderately tight tolerances, we find that the required Krylov subspace dimensions remain independent of grid resolution for a wide range of grid sizes $N$.
3.3 Solving for the Matrix of Coefficients, $S_{1}$: ACS-Preconditioned GMRES
With the orthonormal bases $U_{1}\in\mathbb{R}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{R}^{N_{2}\times r_{y}}$ at hand, we perform a Galerkin projection to derive a reduced GSE for the matrix $S_{1}\in\mathbb{R}^{r_{x}\times r_{y}}$. The reduced GSE for ${\bf S}_{1}$ arises from projecting the residual via the Galerkin condition $U_{1}^{\top}\mathbf{R}_{\mathscr{L}}V_{1}=0$, yielding:
${\bf S}_{1}-\Delta t\left(\tilde{T}_{1}{\bf S}_{1}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}{\bf S}_{1}\tilde{T}_{2}^{\top}+\tilde{T}_{3}{\bf S}_{1}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}{\bf S}_{1}\tilde{T}_{4}^{\top}\right)=\tilde{{\bf B}}_{1},\qquad\text{with}\qquad\tilde{{\bf B}}_{1}=(U_{1}^{\top}U_{0})S_{0}(V_{1}^{\top}V_{0})^{\top}.$ (21)
This matrix-based equation is equivalent to the classical Kronecker formulation:
$\underbrace{\left[I_{r_{x}r_{y}\times r_{x}r_{y}}-\Delta t\left(\tilde{\Phi}_{2}\otimes\tilde{T}_{1}+\tilde{T}_{2}\otimes\tilde{\Phi}_{1}+\tilde{\Sigma}_{2}\otimes\tilde{T}_{3}+\tilde{T}_{4}\otimes\tilde{\Sigma}_{1}\right)\right]}_{\mathcal{A}}\text{vec}({\bf S}_{1})=\text{vec}(\tilde{{\bf B}}_{1}).$ (22)
The projected operators are defined as follows:
$\tilde{T}_{1}$ $=U_{1}^{\top}T_{1}U_{1},\quad\tilde{T}_{3}=U_{1}^{\top}T_{3}U_{1},\quad\tilde{\Phi}_{1}=U_{1}^{\top}\Phi_{1}U_{1},\quad\tilde{\Sigma}_{1}=U_{1}^{\top}\Sigma_{1}U_{1},$
$\tilde{T}_{2}$ $=V_{1}^{\top}T_{2}V_{1},\quad\tilde{T}_{4}=V_{1}^{\top}T_{4}V_{1},\quad\tilde{\Phi}_{2}=V_{1}^{\top}\Phi_{2}V_{1},\quad\tilde{\Sigma}_{2}=V_{1}^{\top}\Sigma_{2}V_{1}.$
Solving the Kronecker formulation $\mathcal{A}\text{vec}({{\bf S}_{1}})=\text{vec}({\tilde{{\bf B}}_{1}})$ in (22) with a direct solver scales as $\mathcal{O}(r^{6})$ in 2D and $\mathcal{O}(r^{9})$ in 3D, where $r=\text{max}(r_{x},r_{y})$, which may be too expensive even for moderate $r$-values. Instead, we solve this equation iteratively using preconditioned GMRES. Instead of adapting GMRES to the matrix equation, we wrap GMRES around the Kronecker form, but exploiting the matrix-equation form (21) for both the $\mathcal{A}$-times-vector operation and the preconditioning step as indicated below, with overall complexity $\mathcal{O}(r^{3})$ in 2D (and $\mathcal{O}(r^{4})$ in 3D). For the GMRES operator-vector product, given a vector $s_{k}=\text{vec}(\mathbf{S}_{k})$, the product of the matrix $\mathcal{A}$ times a vector $s_{k}$ is performed as follows:
  1. 1.
    Given $s_{k}$, reshape to $\mathbf{S}_{k}\in\mathbb{R}^{r_{x}\times r_{y}}$, such that $\text{vec}(\mathbf{S}_{k})=s_{k}$.
  2. 2.
    Compute $\hat{\mathbf{Z}}_{k}\leftarrow\mathbf{S}_{k}-\Delta t\left(\tilde{T}_{1}\mathbf{S}_{k}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}\mathbf{S}_{k}\tilde{T}_{2}^{\top}+\tilde{T}_{3}\mathbf{S}_{k}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}\mathbf{S}_{k}\tilde{T}_{4}^{\top}\right)$, with complexity $\mathcal{O}(r^{3})$.
  3. 3.
    Reshape $\hat{z}_{k}=\text{vec}(\hat{\mathbf{Z}}_{k})$.
For efficiency, GMRES requires preconditioning so that the eigenvalues of the preconditioned matrix are sufficiently clustered to deliver $r$-independent convergence rates  [7]. Here, we employ the ACS preconditioner, with $\mathcal{P}=I_{r_{y}}\otimes\tilde{P}_{1}+\tilde{P}_{2}\otimes I_{r_{x}}$ arising from the Galerkin projection of (17). The ACS preconditioner is expected to deliver optimal rank-independent, superlinear convergence (both observed in our numerical experiments), owing to its connection to the theory of compact-equivalent operators for diffusion-like equations, [^1] The averaged-coefficient advection-diffusion equation is compact-equivalent to the variable-coefficient one. which are known to be effective preconditioners for Krylov methods [1, 2].
We consider left preconditioning as:
$\mathcal{P}^{-1}\mathcal{A}\text{vec}({\bf S}_{1})=\mathcal{P}^{-1}\text{vec}(\tilde{{\bf B}}_{1}).$ (23)
To compute $\mathcal{P}^{-1}\hat{z}_{k}=z_{k}$ for an arbitrary vector $\hat{z}_{k}$ during GMRES, we solve the corresponding SE:
$\tilde{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\tilde{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k},$
where $z_{k}=\text{vec}(\mathbf{Z}_{k})\in\mathbb{R}^{r_{x}r_{y}}$. Specific steps are:
  1. 1.
    Given $\hat{z}_{k}$, reshape to $\hat{\mathbf{Z}}_{k}\in\mathbb{R}^{r_{x}\times r_{y}}$, such that $\text{vec}(\hat{\mathbf{Z}}_{k})=\hat{z}_{k}$.
  2. 2.
    Solve $\mathcal{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\mathcal{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k}$ with complexity $\mathcal{O}(r^{d+1})$, for example with the Bartels-Stewart algorithm [15].
  3. 3.
    Reshape $z_{k}=\text{vec}(\mathbf{Z}_{k})$.
As a result, the $S$-matrix solve scales as $O(r^{d+1})$ overall, a great reduction in complexity compared to the $O(r^{3d})$ scaling of a naïve implementation of GMRES with the Kronecker form.
3.4 Truncation, Efficient Residual Computation
Once the solution ${\bf F}_{1}=U_{1}S_{1}V_{1}^{\top}$ is obtained, we apply truncation to control rank growth and reduce the residual computational load. Since $S_{1}$ resulting from the GMRES solve is not necessarily diagonal, we diagonalize it with SVD [19] [also scaling as $\mathcal{O}(r^{d+1})$] and discard singular values below a threshold $\epsilon$. The SVD truncation gives ${\bf F}_{1}=U_{1,\epsilon}S_{1,\epsilon}V_{1,\epsilon}^{\top}$, with $U_{1,\epsilon}\in\mathbb{R}^{N_{1}\times r_{1}}$, $V_{1,\epsilon}\in\mathbb{R}^{N_{2}\times r_{1}}$, and $S_{1,\epsilon}\in\mathbb{R}^{r_{1}\times r_{1}}$, with $r_{1}\leq\min(r_{x},r_{y})$ (Algorithm 3).
Following [14], we implement a cost-effective way to compute the residual using the identity:
$\|\mathbf{R_{{\mathscr{L}}}}\|$ $=\|\mathbf{F}_{1}-\Delta t\,\mathscr{L}(\mathbf{F}_{1})-\mathbf{F}_{0}\|$
$=\underbrace{\left\|R_{U}\begin{bmatrix}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}\\ \end{bmatrix}R_{V}^{\top}\right\|}_{\in\mathbb{R}^{(r_{x}+5r_{1})\times(r_{y}+5r_{1})}}$
$={\left\|R_{U}\begin{bmatrix}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&{S_{1,\epsilon}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,(I_{\mathcal{R}_{\text{ranks}}}\otimes{S_{1,\epsilon}})\end{bmatrix}R_{V}^{\top}\right\|}\leq\epsilon_{\texttt{tol}},$
where $\mathcal{R}_{\text{ranks}}={\ell_{x}+k_{x}+\ell_{y}+k_{y}}$ is the total rank number of the advection and diffusion coefficients, and $R_{U}$ and $R_{V}$ are the upper triangular matrices arising from the following Q-less QR decomposition:
{_$\{\_\ ,R_{U}\}$ $=\texttt{QR}\left(\begin{bmatrix}{U}_{1}&{U}_{1,\epsilon}&{T_{1}}{U}_{1,\epsilon}&\Phi_{1}{U}_{1,\epsilon}&{T_{3}}{U}_{1,\epsilon}&\Sigma_{1}{U}_{1,\epsilon}\end{bmatrix}\right),$ (24)
{_$\{\_\ ,R_{V}\}$ $=\texttt{QR}\left(\begin{bmatrix}{V}_{1}&{V}_{1,\epsilon}&\Phi_{2}{V}_{1,\epsilon}&{T_{2}}{V}_{1,\epsilon}&\Sigma_{2}{V}_{1,\epsilon}&{T_{4}}{V}_{1,\epsilon}\end{bmatrix}\right).$
The above residual result follows from the property that $U_{1}(U_{1})^{\top}{\bf F}_{0}V_{1}(V_{1})^{\top}={\bf F}_{0}$, with ${\bf F}_{0}=U_{0}S_{0}V_{0}^{\top}$ (where $U_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ and $V_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ have orthonormal columns), since $U_{1}$ and $V_{1}$ contain $U_{0}$ and $V_{0}$, respectively.
Input: Bases $U$, $V$; matrix of coefficients ${\bf S}$; Tolerance $\epsilon$.
Output: Truncated bases $\tilde{U}$, $\tilde{V}$; truncated singular values $\tilde{S}$.
1
2Perform reduced SVD decomposition: $T_{1}\tilde{S}T_{2}=\texttt{SVD}(S)$, where $\tilde{S}=\text{diag}(\sigma_{j})$;
3
4Identify the last index $r_{1}$ such that $\sigma_{r_{1}}/\|S\|>\epsilon$;
5
6Update matrix of singular values: $\tilde{S}\leftarrow\tilde{S}(1:r_{1},1:r_{1})$;
7
8Update left singular vectors: $\tilde{U}\leftarrow UT_{1}(:,1:r_{1})$;
9
10Update the right singular vectors: $\tilde{V}\leftarrow VT_{2}(:,1:r_{1})$;
11
Algorithm 3 Truncated SVD, $\mathcal{T}_{\epsilon}$
3.5 Generalization to Multi-rank Coefficients
We generalize next the adaptive-rank approach to the multi-rank coefficient case in (5) and (6), using backward Euler for now. Discretization of the advection-diffusion equation in this case yields a multi-term GSE of the form:
$\mathbf{F}_{1}-\Delta t\left(\underbrace{\sum_{i=1}^{\ell_{x}}T_{1,i}\mathbf{F}_{1}{\Phi_{i}^{2,x}}^{\top}+\sum_{j=1}^{\ell_{y}}\Phi_{j}^{1,y}\mathbf{F}_{1}T_{2,j}^{\top}+\sum_{k=1}^{k_{x}}T_{3,k}\mathbf{F}_{1}{\Sigma_{k}^{2,x}}^{\top}+\sum_{l=1}^{k_{y}}\Sigma_{l}^{1,y}\mathbf{F}_{1}T_{4,l}^{\top}}_{\mathscr{L}(\mathbf{F}_{1})}\right)={\bf F}_{0}.$ (25)
[…]
${P_{1}}\mathbf{F}_{1}+\mathbf{F}_{1}{P_{2}^{\top}}=\mathbf{F}_{0},$ (26)
where
$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{3,k}}=\sum_{i=1}^{\ell_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\alpha_{x,i}T_{1,i}\right)}+\sum_{k=1}^{k_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\gamma_{x,k}T_{3,k}\right)},$ (27)
$P_{2}$ $=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{4,l}}=\sum_{j=1}^{\ell_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\alpha_{y,j}T_{2,j}\right)}+\sum_{l=1}^{k_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{y,l}T_{4,l}\right)}.$ (28)
Also as before, the bases $U_{1}$ and $V_{1}$ are constructed from orthonormalization of dimension-wise xKrylov from equations (25) and (26) as:
$U_{1}$ $=\kappa^{\texttt{trunc}}_{m}\left(P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}},U_{0}\right),$ (29)
$V_{1}$ $=\kappa^{\texttt{trunc}}_{m}\left(P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}},V_{0}\right).$ (30)
In the multi-rank setting, after a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$, we obtain a GSE similar to (21) for the matrix of coefficients ${\bf S}_{1}$:
${\bf S}_{1}-\Delta_{t}\left(\sum_{i=1}^{\ell_{x}}\tilde{T}_{1,i}{\bf S}_{1}\tilde{\Phi}_{i}^{2,x\top}+\sum_{j=1}^{\ell_{y}}\tilde{\Phi}_{j}^{1,y}{\bf S}_{1}\tilde{T}_{2,j}^{\top}+\sum_{k=1}^{k_{x}}\tilde{T}_{3,k}{\bf S}_{1}\tilde{\Sigma}_{k}^{2,x\top}+\sum_{l=1}^{k_{y}}\tilde{\Sigma}_{l}^{1,y}{\bf S}_{1}\tilde{T}_{4,l}^{\top}\right)=\tilde{{\bf B}}_{1},$ (31)
where the projected operators are analogous to those defined in the rank-one case. We solve (31) for the ${\bf S}_{1}$ matrix using preconditioned GMRES. The preconditioner is constructed by another Galerkin projection to find the projected SE:
$\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}=\tilde{{\bf B}}_{1},$ (32)
with:
${\tilde{P}_{1}}$ $=\sum_{i=1}^{\ell_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{x}}-\Delta_{t}\alpha_{x,i}\tilde{T}_{1,i}\right)}_{\tilde{A}_{1,i}}+\sum_{k=1}^{k_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{x}}-\Delta_{t}\gamma_{x,k}\tilde{T}_{3,k}\right)}_{\tilde{A}_{3,k}},$
$\tilde{P}_{2}$ $=\sum_{j=1}^{\ell_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{y}}-\Delta_{t}\alpha_{y,j}\tilde{T}_{2,j}\right)}_{\tilde{A}_{2,j}}+\sum_{l=1}^{k_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{y}}-\Delta_{t}\gamma_{y,l}\tilde{T}_{4,l}\right)}_{\tilde{A}_{4,l}}.$
This yields the multi-rank-coefficients ACS preconditioner operator:
$\mathscr{P}:\mathbf{S}\mapsto\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}.$ (33)
We summarize the pseudo-algorithm for multi-rank coefficients in Algorithm 4. The tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, $\epsilon_{\texttt{tol}}$ in the algorithm are set as follows. First, we choose the residual tolerance for xKrylov outer iterations, $\epsilon_{\texttt{tol}}$. To prevent rapid growth of the xKrylov subspaces, we choose the Krylov basis truncation threshold as $\epsilon_{\kappa}\leq\epsilon_{\texttt{tol}}$, typically two to three orders of magnitude smaller. The GMRES-ACS tolerance for the inner reduced system solve, $\epsilon_{\texttt{GMRES}}$, is initially set to be less than $\epsilon_{\texttt{tol}}$ for the first outer-basis iteration. It is subsequently adjusted to be two to three orders of magnitude smaller than the relative residual error from the previous xKrylov outer iteration. This adjustment ensures that errors from the inner solve do not pollute the accuracy of the outer solve. The SVD truncation threshold $\epsilon$ is chosen to be one to three orders of magnitude smaller than the tolerance $\epsilon_{\text{tol}}$. Since solution truncation occurs before residual evaluations, the truncation error in the solution should remain below the residual tolerance to prevent residual pollution by over-truncation.
// This algorithm solves Eq. (25) for $\mathbf{F}_{1}=U_{1}S_{1}V_{1}^{\top}$, where $\mathbf{F}_{0}=U_{0}S_{0}V_{0}^{\top}$
Input: Initial condition matrices $U_{0}$, $V_{0}$, $S_{0}$;
Operators $P_{1}$, $P_{2}$, $\{T_{1,i}\}_{i=1}^{\ell_{x}}$, $\{T_{2,j}\}_{j=1}^{\ell_{y}}$, $\{T_{3,k}\}_{k=1}^{k_{x}}$, $\{T_{4,l}\}_{l=1}^{k_{y}}$;
Tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, $\epsilon_{\texttt{tol}}$;
Output: Updated bases $U_{1}$, $V_{1}$;
Truncated singular values $S_{1}$
1Compute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eqs. (27), (28);
2 Compute $l=2+\ell_{x}+\ell_{y}+k_{x}+k_{y}$;
3 Set $U^{(i)}=U_{0}$ and $V^{(i)}=V_{0}$ for $i=1,\dots,l$;
4 Set $U_{1}=U_{0}\quad V_{1}=V_{0}$;
5 while not converged do
6        // Step K1.
7        Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
8        $U_{1},\{U^{(1)}\cdots U^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}}\};\{U^{(1)}\cdots U^{(l)}\};U_{1};\epsilon_{\kappa}\right)$;
9        $V_{1},\{V^{(1)}\cdots V^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}}\};\{V^{(1)}\cdots V^{(l)}\};V_{1};\epsilon_{\kappa}\right)$;
10
11       // Step K2.
12        Solve the reduced GSE (31) for ${\bf S}_{1}$ using GMRES-ACS (33) to tolerance $\epsilon_{\texttt{GMRES}}$;
13        // Step K3.
14        Truncate $\{U_{1,\epsilon},S_{1,\epsilon},V_{1,\epsilon}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}\})$ to tolerance $\epsilon$ (Alg. 3);
15
16       // Step K4.
17        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
18        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
19
20       Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(-\tilde{B}_{1},S_{1,\epsilon},-\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1,\epsilon}\right)\right)R_{V}^{\top}\right\|$;
21
22       if $\|\mathbf{R}\|/\|\mathbf{F}_{0}\|\geq\epsilon_{\texttt{tol}}$ then
23               Reject solution and cycle to augment bases further;
24
25       else
26               Set $U_{1}\leftarrow U_{1,\epsilon}$; $U_{1}\leftarrow V_{1,\epsilon}$; $S_{1}=S_{1,\epsilon}$;
27               Break;
28               Exit loop;
29
30
Algorithm 4 Backward Euler Adaptive-Rank Integrator for the GSE.
3.6 Computational Complexity
We analyze next the computational complexity of Algorithm 4 for backward Euler. For conciseness, we consider the same spatial resolution per dimension, i.e., $N=N_{1}=N_{2}$. Starting from an initial condition $U_{0}S_{0}V_{0}^{\top}$, where $U_{0}\in\mathbb{R}^{N\times r_{0}}$, $V_{0}\in\mathbb{R}^{N\times r_{0}}$, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$, the $m$-th xKrylov iteration generates $U_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, $V_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, and $S_{1}\in\mathbb{R}^{r_{\kappa}\times r_{\kappa}}$, where $r_{\kappa}=(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0}+r^{(m-1)}$. Here, $r^{(m-1)}$ denotes the basis size from the previous xKrylov iteration (initialized with $r^{(0)}=r_{0}$), and $r^{\ast}_{0}\leq r_{0}$ is the possibly truncated initial rank as produced by line (2) of Algorithm 2. Using SVD-truncated QR, this dimension is then reduced to smaller ranks, $r_{x}$ and $r_{y}$. For simplicity, let $\tilde{r}=\max(r_{x},r_{y})$. The matrix $S_{1}$ from the $m$-th inner iteration solve, initially of size $r_{x}\times r_{y}$, is truncated to rank $r_{1}\leq\min(r_{x},r_{y})$ using Algorithm 3. To simplify the complexity analysis, let $r=\max(r_{0},r^{(m-1)},\tilde{r},(\mathcal{R}_{\text{ranks}}+1)r_{1})$, which we consider independent of the mesh-refinement $N$ for practical purposes. The computational complexity of each step is then estimated as follows:
Step K1
Constructing the generalized Krylov basis as outlined in lines (4) and (4) of Algorithm 4 requires performing the operations $A_{1,1:r_{x}}^{-1}U_{0}$ and $A_{3,1:k_{x}}^{-1}U_{0}$, and $A_{2,1:r_{y}}^{-1}V_{0}$ and $A_{4,1:k_{y}}^{-1}V_{0}$, which require solving the systems $AX=U$ and $AX=V$ (subscripts have been omitted for simplicity). With 1D finite-differences, $A$ has a tridiagonal structure, which can be efficiently factorized using direct-solver techniques such as the Thomas algorithm, with $\mathcal{O}(N)$ complexity per column of $X$ [34]. The same complexity holds for applying $P_{1}^{-1}$ and $P_{2}^{-1}$. Additionally, applying $P_{1}U_{0}$, $\Phi_{1,3}U_{0}$, $\Sigma_{1,3}U_{0}$, $P_{2}V_{0}$, $\Phi_{2,4}V_{0}$, and $\Sigma_{2,4}V_{0}$, can be done with $\mathcal{O}(Nr^{\ast}_{0})$ computational complexity for sparse diagonal and tridiagonal matrices, leading to a total complexity of $\mathcal{O}(N(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0})$ to augment the xKrylov subspaces. The use of Algorithm 2 to compute the orthonormal truncated bases $U_{1}$ and $V_{1}$ incurs a complexity of $\mathcal{O}\left(N\,\max\left[\left(2+\mathcal{R}_{\text{ranks}}\right)\,r_{0}^{\ast}\,r^{(m-1)},\;\left(\left(2+\mathcal{R}_{\text{ranks}}\right)\,r_{0}^{\ast}\right)^{2}\right]\right)\sim\mathcal{O}(Nr^{2})$. Generally, we expect $r$ to be much smaller and independent of $N$ when the solution is sufficiently smooth and exhibits a low-rank structure.
Step K2
The application of the reduced GSE operator using matrix-matrix multiplication and summation in the Krylov method (GMRES) features a complexity of $\mathcal{O}\left(\tilde{r}^{d+1}\right)$. Applying the ACS preconditioner requires solving an SE of size $r_{x}\times r_{y}$, which can be achieved using the Bartels-Stewart algorithm [15] with a complexity of $\mathcal{O}(\tilde{r}^{d+1})$. Hence, the $S$-step has $\mathcal{O}(r^{d+1})$ complexity.
Step K3
The SVD decomposition of ${\widetilde{S}}$ in line (3) of Algorithm 3 is of $\mathcal{O}(\tilde{r}^{d+1})$ complexity. Updating the bases in lines (3) and (3) of Algorithm 3 requires $\mathcal{O}(N\tilde{r}{r_{1}})$ operations, where ${r_{1}}\leq\min(r_{x},r_{y})$ is the post-truncation rank of the new solution from Algorithm 3.
Step K4
The QR decomposition to obtain $R_{U}$ and $R_{V}$ in lines (11)-(4) of Algorithm 4 features a complexity of $\mathcal{O}\left(N\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{2}\right)$. The residual computation in line (4) is of $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d+1}\right)$ due to the matrix multiplications, with an additional $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d}\right)$ for the Frobenius norm.
It follows that, as advertised, the overall complexity of the algorithm scales as $Nr^{2}$ and as $r^{d+1}$, with $d$ the dimensionality, which will be numerically verified in Section 4.
4 Numerical Experiments
We demonstrate the overall efficacy of the proposed algorithm with several 2D and 3D numerical examples. We assume homogeneous Dirichlet boundary conditions for simplicity; other boundary conditions (.e.g, periodic, Neumann, etc.) can be easily incorporated in the formulation through the operators $T_{1:4}$ (10). All simulations are performed using MATLAB. For efficient computational handling in 3D, we employ tools from the Tensor Toolbox [3] and the Htucker package [25]. Throughout this section, we denote the diffusion Courant number and the advection Courant-Friedrichs-Lewy (CFL) number by $\lambda_{D}$ and $\lambda_{A}$, respectively, defined as:
$\lambda_{D}=\frac{\Delta t\phi_{\text{max}}}{\Delta x^{2}},\qquad\lambda_{A}=\frac{\Delta t\sigma_{\text{max}}}{\Delta x},$ (34)
where $\phi_{\text{max}}$ and $\sigma_{\text{max}}$ are the maximum of the diffusion and advection coefficients, respectively. These numbers are measures of the condition number of the associated GSEs. For all simulations presented in this section, we did not need to set a maximum number of outer xKrylov iterations because the algorithm consistently converged within a small number of iterations.
Example 4.1: 3D Advection-Diffusion
We consider a rank-1 variable advection coefficient in a 3D domain $x,y,z\in[-1,1]$:
$\sigma^{x}(x,y,z)$ $=\sigma^{1}_{1}(x)\sigma^{2}_{1}(y)\sigma^{3}_{1}(z),$ $\sigma^{1}_{1}(x)$ $=1-x^{2},$ $\sigma^{2}_{1}(y)$ $=2y,$ $\sigma^{3}_{1}(z)$ $=-2z$
$\sigma^{y}(x,y,z)$ $=\sigma^{1}_{2}(x)\sigma^{2}_{2}(y)\sigma^{3}_{2}(z),$ $\sigma^{1}_{2}(x)$ $=-2x,$ $\sigma^{2}_{2}(y)$ $=1-y^{2},$ $\sigma^{3}_{2}(z)$ $=2z$
$\sigma^{z}(x,y,z)$ $=\sigma^{1}_{3}(x)\sigma^{2}_{3}(y)\sigma^{3}_{3}(z),$ $\sigma^{1}_{2}(x)$ $=4x,$ $\sigma^{2}_{2}(y)$ $=2y,$ $\sigma^{3}_{3}(z)$ $=1-z^{2}.$
and a rank-3 variable diffusion coefficient:
$\phi^{x}(x,y,z)=\phi^{y}(x,y,z)=\sum_{i=1}^{3}\phi^{1}_{i}(x)\phi^{2}_{i}(y)\phi^{3}_{i}(z),$
${\phi}^{1}_{1}(x)$ $=\exp\left(-\left(x-0.3\sin(x)\right)^{2}\right),$ ${\phi}^{2}_{1}(y)$ $=\exp\left(-\left(y-0.3\cos(y)\right)^{2}\right),$ ${\phi}^{3}_{1}(z)$ $=\exp\left(-\left(y-0.3\sin(z)\right)^{2}\right),$
${\phi}^{1}_{2}(x)$ $=\exp\left(-\left(x-0.6\sin(\pi x)\right)^{2}\right),$ ${\phi}^{2}_{2}(y)$ $=\exp\left(-\left(y-0.6\sin(\pi y)\right)^{2}\right),$ ${\phi}^{3}_{2}(z)$ $=\exp\left(-\left(y-0.6\sin(\pi z)\right)^{2}\right),$
${\phi}^{1}_{3}(x)$ $=\exp\left(-\left(x-0.6\sin(2\pi x)\right)^{2}\right),$ ${\phi}^{2}_{3}(y)$ $=\exp\left(-\left(y-0.6\sin(2\pi y)\right)^{2}\right)$ ${\phi}^{3}_{3}(z)$ $=\exp\left(-\left(y-0.6\sin(2\pi z)\right)^{2}\right).$
The initial condition is a rank-2 Gaussian:
$f(x,y,z,t=0)$ $=0.5\exp\left(-400\left((x-0.3)^{2}+(y-0.35)^{2}+(z-0.2)^{2}\right)\right)$
$+0.8\exp\left(-400\left((x-0.65)^{2}+(y-0.5)^{2}+(z-0.55)^{2}\right)\right).$
We use a spatial grid with $N=300$ in all dimensions, and the time-step $\Delta t$ is set to $\lambda_{D}\Delta x^{2}$, with $\lambda_{D}$ ranging from $140$ to $1125$. We test our adaptive-rank algorithm using three different integrators: backward Euler (BE), DIRK2, and DIRK3, as detailed in Tables 3 and 3 in the Appendix. The residual tolerance is $\epsilon_{\texttt{tol}}=10^{-4}$ for BE, DIRK2, and DIRK3, the basis truncation threshold is $\epsilon_{\kappa}=10^{-6}$, the SVD truncation threshold is $\epsilon=10^{-6}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-6}$.
We investigate numerically the following: (1) high-order temporal convergence, (2) effectiveness of the xKrylov iterations in capturing the range of solution, (3) performance of the proposed ACS preconditioner, and (4) verification of the overall computational complexity.
High-Order Temporal Convergence: Figure 1(a) illustrates the $L_{1}$ error norm vs time stepping size (characterized by $\lambda_{D}$) by showcasing the performance of the adaptive-rank integrators for BE, DIRK2, and DIRK3. The expected asymptotic temporal error order of convergence for our proposed adaptive-rank integrators, depicted by lines with markers, is observed. Additionally, the algorithm adeptly captures the rank evolution of the solution in the x, y, and z directions, as illustrated in Figure 1(b).
Figure 1: Example 4.1. (a) Temporal convergence study of BE, DIRK2, and DIRK3 schemes. (b) Rank evolution of the solution in the $x$, $y$, and $z$ directions over time for $\lambda_{D}=140$.
Effectiveness of the xKrylov Iterations: We evaluate the xKrylov subspace dimension required to achieve tolerances of $\epsilon_{\texttt{tol}}=10^{-4},10^{-5},10^{-6}$. We perform a single time-step integration with a fixed step size $\Delta t=10^{-3}$ with grid refinements in $N$. To isolate the effect of $\lambda_{D}$ (a measure of the GSE condition number) on the convergence behavior of xKrylov, we mitigate other potential influencing factors, such as excessive truncation of the Krylov subspace, insufficient convergence of the GMRES-ACS solver, and truncation errors in the constructed solution, by setting related thresholds to $\epsilon=\epsilon_{\texttt{GMRES}}=\epsilon_{\kappa}=10^{-10}$. Figure 2 illustrates the xKrylov subspace dimension required for convergence as a function of the diffusion Courant number $\lambda_{D}$ for a fixed $\Delta t$ and varying spatial resolutions $N$. Our results suggest that, for a fixed tolerance, the convergence of the Krylov subspaces remains largely insensitive to grid refinement across a broad range of $\lambda_{D}$. However, for extremely fine mesh resolutions, beyond typical practical computational needs, the convergence rate of xKrylov scales as $\lambda_{D}^{1/4}$, consistently with the $\kappa^{1/4}$ scaling (with $\kappa$ the condition number) previously reported for the Poisson equation [24, 4].
Figure 2: Example 4.1. Convergence study using the xKrylov subspace method with DIRK3 adaptive-rank integrator for several outer basis tolerances, $\epsilon_{tol}$.
Performance of ACS-Preconditioner: We compare the performance of the GMRES solver with and without the ACS preconditioner for the inner solve. Figure 3 illustrates the GMRES residual history over iterations for ACS-preconditioned and unpreconditioned cases. Results at the first time step ($\Delta t=0.01$) and the third outer-basis augmentation are shown, comparing different mesh sizes. The results confirm that ACS-preconditioned GMRES delivers mesh-independent and super-linear convergence rates, which is not the case for unpreconditioned GMRES.
Figure 3: Example 4.1. Log-log plots of GMRES residuals versus iterations, with and without preconditioning. Simulations are conducted for a single time step with $\Delta t=0.01$ at the third outer-basis augmentation, using grid resolutions $N_{x}=N_{y}=N_{z}=200$ in (a), $1000$ in (b), and $2500$ in (c).
Figure 4 (a) illustrates the average condition number of the Galerkin-projected diffusion and advection operators as the dimension of the proposed xKrylov subspaces increases across 12 outer iterations. For these results, the grid size is fixed at $N=1000$ in 3D, and a single time step of $\Delta t=0.01$ is used. Figures 4 (b) - (d) show the GMRES residual histories for the ACS-preconditioned and unpreconditioned cases, corresponding to the second, fourth, and twelfth outer xKrylov iterations. These correspond to xKrylov subspace dimensions of $r=37$, $r=73$, and $r=217$, respectively. As the dimension of the xKrylov subspace grows, unpreconditioned GMRES convergence for the projected problem becomes progressively slower, as expected. In contrast, the ACS-preconditioner effectively maintains rapid convergence rates independent of subspace size $r$.
Figure 4: Example 4.1. (a) Average condition number for Galerkin-projected advection and diffusion operators. (b)-(d) Log-log plots of GMRES residuals versus iterations, with and without preconditioning for various subspace sizes $r_{\kappa}=37$, $r_{\kappa}=73$, and $r_{\kappa}=217$, respectively. These cases correspond to the highlighted blue data points in (a). Simulations are conducted for a single time step with $\Delta t=0.01$ and a grid of $N_{x}=N_{y}=N_{z}=1000$.
Verification of the Computational Complexity scaling: We verify the $\mathcal{O}(N)$ computational complexity claim next. We vary $N$ from $100$ to $30000$ and the time step is fixed at $\Delta t=10^{-3}$. Such range of mesh sizes is chosen to prevent the $O(r^{4})$ contributions of the ACS-GMRES solver from polluting the mesh-resolution scaling. Figure 5(a) displays $N$ vs. computational time for $10$ time steps of integration. We observe that the complexity remains linear with respect to $N$. The finite y-intercept corresponds to the $\mathcal{O}(r^{4})$ cost of the GMRES-ACS solve in 3D. Figure 5(b) displays the computational time required for the preconditioned GMRES-ACS solver vs. rank for a single time step, using BE with $N=10,000$ and $\epsilon_{\texttt{GMRES}}=10^{-10}$. For this test, we allowed the basis to grow without truncation, recording the simulation time required for the GMRES solve at each basis augmentation iteration. The expected $\mathcal{O}(r^{4})$ scaling in 3D is found. These studies verify that the complexity scaling of our approach is $\mathcal{O}(Nr^{2}+r^{4})$ in 3D.
Figure 5: Example 4.1. (a) Demonstration of ${\cal O}(N)$ computational complexity of DIRK3 adaptive-rank integrator in 3D. (b) Demonstration of ${\cal O}(r^{4})$ computational complexity of the GMRES-ACS solver in 3D. The execution times were recorded using MATLAB’s timeit function.
Example 4.2: 3D Swirling Deformation Flow
We consider next a 3D swirling deformation flow:
$f_{t}+\left[(1-x^{2})\cdot 2y\cdot-2z\cdot f\right]_{x}+\left[2x\cdot(1-y^{2})\cdot 2z\cdot f\right]_{y}+\left[-2x\cdot 2y\cdot 2(1-z^{2})\cdot f\right]_{z}=\nu\Delta f,\quad x,y,z\in[-1,1].$ (35)
We set the diffusion coefficient $\nu=1/500$. The initial condition is a bi-Gaussian function given by:
$f(x,y,0)$ $=0.5\exp\left(-\frac{1}{2\sigma^{2}}\left[(x-0.3)^{2}+(y-0.35)^{2}+(z-0.3)^{2}\right]\right)$
$+0.8\exp\left(-\frac{1}{2\sigma^{2}}\left[(x+0.5)^{2}+(y+0.5)^{2}+(z+0.5)^{2}\right]\right),$
where $\sigma=0.15$. We choose $N_{x}=N_{y}=N_{z}=1000$. The outer basis residual tolerance is $\epsilon_{\texttt{tol}}=10^{-5}$, the basis truncation threshold is $\epsilon_{\kappa}=10^{-5}$, the SVD truncation threshold is $\epsilon=10^{-5}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-10}$.
This example demonstrates the algorithm’s ability to adapt to dynamic rank evolution. The algorithm performs nine xKrylov augmentations at the initial time step, two augmentations at the third and seventh time steps, and a single augmentation otherwise. Figure 6 shows rank evolution. The solution rank rises sharply in the $x$ and $y$ directions [Figures 6 (a)-(b)]. The solution rank in the $z$ direction evolves more gently [Figure 6 (c)] and is roughly one-third of that in the $x$ and $y$ directions. Figures 6 (d)-(e) showcase how the solution deforms according to the prescribed advection coefficients and diffuses slowly over time. Figure 7 shows snapshots of the solution in the $x\text{--}z$ plane at $y=-0.5$. As the flow-driven deformation increases the complexity of the solution structure, the rank grows to capture these features; when the diffusion process starts to dominate, the rank decreases as the solution smooths. This example shows that the low-rank assumption is not intrinsic to our proposed algorithm. Even when the rank becomes moderately large, the algorithm remains capable of capturing the underlying dynamics effectively. Naturally, the benefit of a low-rank approximation decreases as the solution approaches a full-rank structure.
Figure 6: Example 4.2. 3D Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=14$ on a $1000\times 1000\times 1000$ grid. (a)–(c) Evolution of rank, xKrylov subspace before and after truncation, and number of xKrylov iterations over time for the $x$, $y$, and $z$ directions, respectively. (d)–(e) Solution profiles at $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$. The final simulation time is $T_{f}=10$.
Figure 7: Example 4.2. 3D Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=14$ on a $1000\times 1000\times 1000$ grid. (a)–(f) snapshots of the solution in the $x\text{--}z$ plane at $y=-0.5$, corresponding to times $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$, respectively.
Example 4.3: A 2D Steady-State Balanced-Advection-Diffusion Problem
The last example prescribes diffusion and advection coefficients that exactly balance each other, resulting in a non-trivial steady-state solution. We seek an equilibrium solution satisfying the steady-state condition:
$0=\nabla\cdot\left(\bm{\mathit{\Phi}}\cdot\nabla f_{\text{eq}}-\bm{\mathit{\Sigma}}f_{\text{eq}}\right),$ (36)
a solution of which is:
$\bm{\mathit{\Sigma}}=\bm{\mathit{\Phi}}\cdot\nabla\log(|f_{\text{eq}}|).$
This equation defines the steady-state advective flow $\bm{\mathit{\Sigma}}$ for given diffusion coefficient $\bm{\mathit{\Phi}}$ and steady-state solution $f_{\text{eq}}$. We construct a steady-state solution in the domain $0\leq x,y\leq 1$ to be (up to an arbitrary constant):
$f_{\text{eq}}(x,y)=\left(x(1-x)\right)^{2}\left(y(1-y)\right)^{2},$
which satisfies homogeneous Dirichlet boundary conditions. The diffusion coefficients are defined as:
$\phi^{x}(x,y)=\phi^{y}(x,y)$ $=\phi^{1}(x)\phi^{2}(y),\qquad\phi^{1}(x)=x^{2}(1-x)^{2},\qquad\phi^{2}(y)=y^{2}(1-y)^{2}.$
The variable advection coefficients are $\sigma^{x}(x,y)=\sigma_{1}^{1}(x)\sigma_{1}^{2}(y)$, $\sigma^{y}(x,y)=\sigma_{2}^{1}(x)\sigma_{2}^{2}(y),$ where:
$\sigma_{1}^{1}(x)$ $=2x(1-3x+2x^{2}),\qquad\sigma_{1}^{2}(y)=y^{2}(1-y)^{2},\qquad\sigma_{2}^{1}(x)=x^{2}(1-x)^{2},\qquad\sigma_{2}^{2}(y)=2y(1-3y+2y^{2}).$
We select the following rank-1 initial condition: $f(x,y,0)=\left|\sin(2\pi x)\sin(2\pi y)\right|$. The outer basis residual tolerance is $\epsilon_{\texttt{tol}}=10^{-3}$, the basis truncation threshold is $\epsilon_{\kappa}=10^{-8}$, the SVD truncation threshold is $\epsilon=10^{-8}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-8}$.
Figure 8 shows the error of the numerical steady-state solution, reached using ten time steps, relative to the prescribed analytical steady state as a function of mesh size. The advection CFL number, $\lambda_{A}$, for a fixed time step $\Delta t=1000$ ranges from $1,190$ to $36,076$ as the mesh is refined. The results are obtained by rescaling the numerical solution so that its normalization constant is the same as the analytical one. They confirm second-order spatial convergence and demonstrate the ability of our algorithm to take large timesteps to reach the correct steady state with expected asymptotic accuracy.
Figure 8: Example 4.3. Log-log plot of error in the numerical steady-state solution compared to the prescribed analytical steady state as a function of mesh size. The advection CFL number, $\lambda_{A}$, ranges from $1,190$ to $36,076$. The plot shows second-order spatial convergence in the steady state.
5 Conclusion
We have proposed a computationally efficient adaptive-rank implicit algorithm for solving time-dependent advection-diffusion partial differential equations with variable coefficients. Our algorithm is based on several key strategies: constructing dimension-wise basis from xKrylov subspaces, projecting the original matrix equation onto a reduced system to solve for a coefficient matrix, devising an effective Averaged-Coefficient Sylvester (ACS) preconditioner for the reduced system, and a computationally efficient procedure for the residual evaluation. This combined approach achieves a computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$, with $r$ being the rank of the solution during Krylov iteration process and $d$ the dimensionality, effectively mitigating the curse of dimensionality and making it comparable in efficiency to the constant-coefficient case [14]. Such complexity scaling is numerically verified with our extensive tests, with the maximal rank $r$ found to be independent of the 1D grid size $N$ for practical grid resolutions and convergence tolerances. Future research directions include extending the algorithm to higher dimensions and nonlinear PDEs.
Acknowledgements
This work was partially supported by the Multifaceted Mathematics Integrated Capability Centers (MMICCs) program of DOE Office of Applied Scientific Computing Research (ASCR). Los Alamos National Laboratory (LANL) is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). J.Q. and H.E. were supported by NSF grant NSF-DMS-2111253. J.Q. was also supported by Air Force Office of Scientific Research FA9550-22-1-0390, Department of Energy DE-SC0023164 and Air Force Office of Scientific Research (AFOSR) FA9550-24-1-0254 via the Multidisciplinary University Research Initiatives (MURI) Program. H.E. was also supported by MMICCs at LANL during the summer of 2024.
Appendix A Discretization of the Diffusion Operator
We derive the matrix formulation of the discretization of the diffusion operator for the two-dimensional advection-diffusion equation with separable diffusion coefficients. We consider the discretization of the diffusion term in the $x$-direction, with other directions treated similarly:
$\frac{\partial}{\partial x}\left(\phi^{x}(x,y)\frac{\partial u(x,y)}{\partial x}\right),$ (37)
where the diffusion coefficient is separable:
$\phi^{x}(x,y)=\phi^{1}(x)\phi^{2}(y).$ (38)
We discretize this term using a 3-point stencil on a uniform grid with spacing $\Delta x$ in the $x$-direction. Let $u_{i,j}\approx u(x_{i},y_{j})$, with $i=0,\cdots,N_{x}-1$ and $j=0,\cdots,N_{y}-1$. We approximate the spatial derivative using central differences. The discrete approximation at grid point $(i,j)$ is:
$\left[\frac{\partial}{\partial x}\left(\phi^{x}\frac{\partial u}{\partial x}\right)\right]_{i,j}\approx\frac{1}{\Delta x^{2}}\left[\phi^{1}_{i+\frac{1}{2}}(u_{i+1,j}-u_{i,j})-\phi^{1}_{i-\frac{1}{2}}(u_{i,j}-u_{i-1,j})\right]\phi^{2}_{j},$ (39)
with $\phi^{1}_{i+\frac{1}{2}}\phi^{2}_{j}=\phi^{1}(x_{i+\frac{1}{2}})\phi^{2}(y_{j}).$ To express the discretization in matrix form, we define:
  • $\mathbf{F}$: matrix of solution values, $[\mathbf{F}]_{i,j}=f_{i,j}$.
  • $\Phi^{1}_{+}$: diagonal matrix with elements $\left[\Phi^{1}_{+}\right]_{i,i}=\phi^{1}_{i+\frac{1}{2}}$.
  • $\Phi^{1}_{-}$: diagonal matrix with elements $\left[\Phi^{1}_{-}\right]_{i,i}=\phi^{1}_{i-\frac{1}{2}}$.
  • $\Phi_{2}$: diagonal matrix with elements $\left[\Phi_{2}\right]_{j,j}=\phi^{2}_{j}$.
We also define the forward and backward difference matrices in the $x$-direction for the interior nodes $i=1,\cdots,N_{x}-2$. The nodes $i=0$, $i=N_{x}-1$ rest at the boundary and enforce homogeneous Dirichlet boundary conditions (other boundary conditions can be readily considered):
$D_{+}=\frac{1}{\Delta x}\begin{pmatrix}-1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\\ 0&\cdots&0&0&-1\end{pmatrix},D_{-}=\frac{1}{\Delta x}\begin{pmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\end{pmatrix}.$ (40)
Using these definitions, the discrete diffusion term in Eq. (39) can be represented in matrix form as:
$\left(\Phi^{1}_{+}D_{+}-\Phi^{1}_{-}D_{-}\right)\mathbf{F}{\Phi_{2}^{\top}}.$ (41)
We note that the operator multiplying $\mathbf{F}$ from the left is analogous to $D^{\Phi}_{xx}$ in (9). The discretization in the $y$-direction follows the same strategy.
Appendix B Discretization of the Advection Operator
We derive the matrix formulation of the discretization of the advection operator for the two-dimensional advection-diffusion equation with variable, separable advection coefficients. As before, we focus on the discretization of the advection term in the $x$-direction:
$\frac{\partial}{\partial x}\left(\sigma^{x}(x,y)u(x,y)\right),$ (42)
where the advection coefficient is separable:
$\sigma^{x}(x,y)=\sigma^{1}(x)\sigma^{2}(y).$ (43)
We discretize this term on a uniform grid with spacing $\Delta x$ in the $x$-direction and $\Delta y$ in the $y$-direction. To preserve linearity of our discretization without introducing too much numerical dissipation, we employ a central difference scheme to approximate the spatial derivative. This discretization is monotonicity-preserving when the following condition is satisfied:
$\Delta x<2\frac{\phi_{\text{max}}}{\sigma_{\text{max}}},$ (44)
where $\sigma_{\text{max}}$ is the maximum magnitude of the advection velocity and $\phi_{\text{max}}$ is the maximum diffusion coefficient. This condition is practical in our formulation due to the very favorable scaling of our low-rank representation with mesh size and rank.
The discrete approximation at the grid point $(i,j)$ is:
$\left[\frac{\partial}{\partial x}\left(\sigma^{x}(x,y)u(x,y)\right)\right]_{i,j}$ $\approx\frac{1}{2\Delta x}\left[\sigma^{1}_{i+\frac{1}{2}}(u_{i+1,j}+u_{i,j})-\sigma^{1}_{i-\frac{1}{2}}(u_{i,j}+u_{i-1,j})\right]\sigma^{2}_{j},$ (45)
where $\sigma^{1}_{i+\frac{1}{2}}=\sigma^{1}(x_{i+\frac{1}{2}})$ is the advection coefficient evaluated at the midpoint between $x_{i}$ and $x_{i+1}$, and similarly for $\sigma^{1}_{i-\frac{1}{2}}$, and $\sigma^{2}_{j}$ is evaluated at $y_{j}$. To express the discretization in matrix form, we define:
  • $\mathbf{F}$: matrix of solution values, $[\mathbf{F}]_{i,j}=f_{i,j}$.
  • $\Sigma^{1}_{+}$ and $\Sigma^{1}_{-}$: diagonal matrices with elements $\left[\Sigma^{1}_{+}\right]_{i,i}=\sigma^{1}_{i+\frac{1}{2}}$ and $\left[\Sigma^{1}_{-}\right]_{i,i}=\sigma^{1}_{i-\frac{1}{2}}$.
  • $\Sigma_{2}$: diagonal matrix with elements $\left[\Sigma_{2}\right]_{j,j}=\sigma^{2}_{j}$.
We also define the difference matrices in the $x$-direction:
$S_{+}=\frac{1}{2}\begin{pmatrix}1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\\ 0&\cdots&0&0&1\end{pmatrix},S_{-}=\frac{1}{2}\begin{pmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\end{pmatrix}.$ (46)
Using these definitions, the discrete advection term in the $x$-direction can be represented as:
$\frac{1}{\Delta x}\left(\Sigma^{1}_{+}S_{+}-\Sigma^{1}_{-}S_{-}\right)\mathbf{F}{\Sigma_{2}^{\top}}.$ (47)
We note that the operator multiplying $\mathbf{F}$ from the left is analogous to $D^{\Sigma}_{x}$ in Equation (9). The discretization in the $y$-direction follows the same strategy.
Appendix C High-Order Temporal Discretization
We extend the backward Euler implicit integrator to higher-order DIRK time discretizations for the advection-diffusion operator, as compactly written in (11).
C.1 DIRK Scheme
DIRK methods are characterized by the Butcher tableau, Table 1.
Table 1: Butcher tableau for an $s$-stage DIRK scheme. Here, $s$ represents the number of stages in DIRK, $c_{i}$ represents the intermediate stage, $a_{ij}$ is a lower-triangular matrix with coefficients used to approximate the solution at each stage, and $b_{j}$ are the quadrature weights used to update the solution in the final step.
$c_{1}$ $a_{11}$ 0 $\ldots$ 0
$c_{2}$ $a_{21}$ $a_{22}$ldots 0
$\vdots$ $\vdots$ $\vdots$ $\ddots$ $\vdots$
$c_{s}$ $a_{s1}$ $a_{s2}$ldots $a_{ss}$
$b_{1}$ $b_{2}$ldots $b_{s}$
We focus on stiffly accurate DIRK methods, where the coefficients satisfy $b_{i}=a_{si}$ for $i=1,\ldots,s$, ensuring that the final solution update corresponds to the solution at the last DIRK stage. The Butcher tables for DIRK2 and DIRK3 are presented in Tables 3 and 3, respectively.
Table 2: DIRK2 Butcher table, $\gamma=1-\frac{\sqrt{2}}{2}$.
$\gamma$ $\gamma$ 0
$1$ $1-\gamma$ $\gamma$
$1-\gamma$ $\gamma$
Table 3: DIRK3 Butcher table, $x=0.4358665215$.
$x$ $x$ 0 0
$\frac{1+x}{2}$ $\frac{1-x}{2}$ $x$ 0
$1$ $-\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $-\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
$-\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $-\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
Similarly to the backward Euler case, the matrix-based discretization of the advection-diffusion equation using the operator $\mathscr{L}$ defined in (11) results in a GSE that must be solved at each DIRK stage. The corresponding DIRK scheme for the matrix differential equation (11) from $t^{(0)}$ to $t^{(1)}$ is written as:
$\begin{split}\mathbf{F}^{(k)}&=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k}a_{k\ell}\mathbf{Y}_{\ell},\qquad k=1,2,\ldots,s,\end{split}$ (48a)
$\begin{split}\mathbf{Y}_{k}&=\mathscr{L}(\mathbf{F}^{(k)};t^{(k)}),\qquad t^{(k)}=t^{0}+c_{k}\Delta t,\qquad k=1,2,\ldots,s,\end{split}$ (48b)
$\begin{split}\mathbf{F}_{1}&=\mathbf{F}^{(s)}=\mathbf{F}_{0}+\Delta t\sum\limits_{k=1}^{s}b_{k}\mathbf{Y}_{k}.\end{split}$ (48c)
Within each stage $k$ of the DIRK scheme, one must solve the residual equation given by:
${\bf R}^{(k)}_{\mathscr{L}}=\mathbf{F}^{(k)}-a_{kk}\Delta t\mathscr{L}(\mathbf{F}^{(k)})-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell},\quad k=1,2,\ldots,s.$ (49)
Expanding the operator $\mathscr{L}$ in the residual equation above results in the following $k$-th stage GSE:
$\mathbf{F}^{(k)}-a_{k,k}\Delta t\left(T_{1}\mathbf{F}^{(k)}\Phi_{2}{{}^{\top}}+\Phi_{1}\mathbf{F}^{(k)}T_{2}^{\top}+T_{3}\mathbf{F}^{(k)}\Sigma_{2}^{\top}+\Sigma_{1}\mathbf{F}^{(k)}T_{4}^{\top}\right)=\mathbf{B}^{(k)},\quad k=1,2,\ldots,s.$ (50)
For the DIRK update, we adopt the strategy outlined in [14], with minor modifications to improve efficiency. In particular, we: (i) construct the basis using the time-step of the first stage and retain it for subsequent stages, (ii) evolve the projected system for the $S_{1}$ matrix across the DIRK stages, and (iii) truncate and evaluate the residual to determine whether to admit the solution or return to step (i) to further augment the basis. Key modifications vs. [14] include truncating the solution before the residual computation, and only checking the residual at the end of the DIRK step. The reduced matrix is solved using preconditioned GMRES, as before.
C.2 Adaptive-Rank Strategy for High-Order Temporal Discretization via DIRK
[…]
$\mathbf{R}^{(k)}_{\tilde{\mathscr{L}}}=\mathbf{F}_{1}-a_{kk}\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell}.$ (51)
Expanding this residual using $\tilde{\mathscr{L}}$, defined in (16), leads to the $k$th-stage SE:
$\underbrace{\left[\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\alpha_{2}T_{1}\right)}_{A_{1}}+\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\gamma_{2}T_{3}\right)}_{A_{3}}\right]\mathbf{F}^{(k)}}_{P_{1}}+\mathbf{F}^{(k)}\underbrace{\left[\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\alpha_{1}T_{2}\right)^{\top}}_{A^{\top}_{2}}+\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\gamma_{1}T_{4}\right)^{\top}}_{A^{\top}_{4}}\right]}_{P_{2}^{T}}=\mathbf{B}^{(k)}.$ (52)
Note that the operators $A_{1:4}$ are fixed throughout the stages since the DIRK method uses constant diagonal Butcher tableau coefficients.
We follow [14] to perform high-order integration. The key idea is to consider the basis fixed across DIRK stages, and only evolve the matrix of coefficients ${{\bf S}}_{1}$ through the stages. The proposed procedure is as follows:
  1. Step 1.
    Prediction of Krylov basis functions: construct a set of orthonormal bases $U_{1}$ and $V_{1}$ from the Krylov-based low-rank implicit solver at the first DIRK stage, i.e. backward Euler with time stepping size $c_{1}\Delta t$.
  2. Step 2.
    For $k=1:s$ (per DIRK stage)
    1. (a)
      Solve reduced SE for ${\bf S}^{(k)}$ using preconditioned GMRES,
      ${\bf S}^{(k)}-\Delta t\left(\tilde{T}_{1}{\bf S}^{(k)}\tilde{\Phi}^{\top}_{2}+\tilde{\Phi}_{1}{\bf S}^{(k)}\tilde{T}_{2}^{\top}+\tilde{T}_{3}{\bf S}^{(k)}\tilde{\Sigma}^{\top}_{2}+\tilde{\Sigma}_{1}{\bf S}^{(k)}\tilde{T}_{4}^{\top}\right)=\tilde{{\bf B}}^{(k)},$ (53)
      with
      $\tilde{{\bf B}}^{(k)}=U_{1}^{\top}\mathbf{F}_{0}V_{1}+\Delta t\sum\limits_{\ell=1}^{k-1}{a_{k\ell}\tilde{Y_{\ell}}},$
      Here $\tilde{Y_{\ell}}=U_{1}^{\top}{Y}_{\ell}V_{1}\in\mathbb{R}^{r_{x}\times r_{y}}$. The preconditioner used is given by
      ${\bf S}\mapsto\tilde{P}_{1}{\bf S}+{\bf S}\tilde{P}_{2},$ (54)
      where $\tilde{P}_{1}=U_{1}^{T}P_{1}U_{1}$ and $\tilde{P}_{2}=V_{1}^{T}P_{1}V_{1}$, and is valid for all stages since the basis constructed is fixed throught the DIRK stages by assumption. Solve for ${\bf S}^{(k)}$ and obtain the intermediate RK solutions as $U_{1}S^{(k)}V_{1}^{\top}$. To further improve computational efficiency, from (53), we have,
      $\tilde{Y}_{\ell}$ $=\frac{1}{a_{\ell\ell}\Delta t}\left({{\bf S}^{(\ell)}}-{\tilde{{\bf B}}^{(\ell)}}\right),$
      leading to an efficient computation of $\tilde{{\bf B}}^{(k)}$:
      $\tilde{{\bf B}}^{(k)}=\tilde{{\bf B}}_{1}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left({{\bf S}^{(\ell)}}-{\tilde{{\bf B}}^{(\ell)}}\right),$ (55)
      with $\tilde{{\bf B}}_{1}$ defined in (21). This avoids the need to evaluate the full size ${\bf B}^{(k)}$ in (51).
  3. Step 3.
    Perform truncation of evolved solution, i.e. $\{U_{1,\epsilon},S_{1,\epsilon},V_{1,\epsilon}^{\top}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}^{\top}\})$, with $S_{1}:=S_{1}^{(s)}$.
  4. Step 4.
    Efficiently evaluate the residual norm,
    $\left\|{{\bf R}^{(s)}_{\mathscr{L}}}\right\|=\left\|{R}_{U}{\text{diag}\left(-\tilde{B}^{(s)}_{1},S_{1,\epsilon},-a_{ss}\Delta t\left(I_{{\mathcal{R}}_{\text{ranks}}}\otimes S_{1,\epsilon}\right)\right)}{R}_{V}^{\top}\right\|,$
    where $R_{U}$ and $R_{V}$ are upper triangular matrices from a reduced Q-less QR decomposition as described in (24), which can be done once for all DIRK stages.
  5. Step 5.
    Compare the computed residual norm $\|{\bf R}^{(s)}_{\mathscr{L}}\|$ with the given error tolerance $\epsilon_{\texttt{tol}}$. If the tolerance is met, then admit the solution with ${\bf F}_{1}=U_{1,\epsilon}S_{1,\epsilon}V_{1,\epsilon}^{\top}$; otherwise, we go back to Step 1 to further augment the xKrylov subspaces.
C.3 DIRK Generalization to Multi-Rank Advection-Diffusion Coefficients
We now extend the formulation to the multi-rank case, as described in (7). Using a DIRK method for temporal discretization, the $k$-th stage GSE becomes:
$\mathbf{F}^{(k)}-a_{kk}\Delta t\left(\underbrace{\sum_{i=1}^{\ell_{x}}T_{1,i}\mathbf{F}^{(k)}{\Phi_{i}^{2,x}}^{\top}+\sum_{j=1}^{\ell_{y}}\Phi_{j}^{1,y}\mathbf{F}^{(k)}{T_{2,j}}^{\top}+\sum_{k=1}^{k_{x}}T_{3,k}\mathbf{F}^{(k)}{\Sigma_{k}^{2,x}}^{\top}+\sum_{l=1}^{k_{y}}\Sigma_{l}^{1,y}\mathbf{F}^{(k)}{T_{4,l}}^{\top}}_{\mathscr{L}(\mathbf{F}^{(k)})}\right)=\mathbf{B}^{(k)},\quad k=1,2,\ldots,s$ (56)
where $\mathbf{B}^{(k)}$ is as defined in (51). The terms $T_{1:4,i}$ represent compositions of difference operators with diffusion and advection coefficients, similar to (10). Left subscripts 1 and 2 correspond to diffusion, while 3 and 4 correspond to advection in the $x$ and $y$ directions, respectively. The right subscript denotes the rank of the respective diffusion and advection coefficients. This equation generalizes (13), which is recovered for $\ell_{x}=\ell_{y}=k_{x}=k_{y}=1$ and equal diffusion and advection coefficients in both the $x$ and $y$ directions. It also generalizes the backward Euler special case with a single stage and $a_{1,1}=1$, yielding $\mathbf{B}^{(1)}=\mathbf{F}_{0}$, consistently with (13).
Using approximations in (62), we arrive at the $k$th stage approximated SE:
[…]
where
$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{3,k}}=\sum_{i=1}^{\ell_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-a_{kk}\Delta t\alpha_{x,i}T_{1,i}\right)}+\sum_{k=1}^{k_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-a_{kk}\Delta t\gamma_{x,k}T_{3,k}\right)},$
$P_{2}$ $=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{4,l}}=\sum_{j=1}^{\ell_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-a_{kk}\Delta t\alpha_{y,j}T_{2,j}\right)}+\sum_{l=1}^{k_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-a_{kk}\Delta t\gamma_{y,l}T_{4,l}\right)}.$
In order to obtain a reduced equation for the $k$th stage matrix of coefficients ${\bf S}^{(k)}$, we perform a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$ to obtain :
${\bf S}^{(k)}-a_{kk}\Delta_{t}\left(\sum_{i=1}^{\ell_{x}}\tilde{T}_{1,i}{\bf S}^{(k)}\tilde{\Phi}_{i}^{2,x\top}+\sum_{j=1}^{\ell_{y}}\tilde{\Phi}_{j}^{1,y}{\bf S}^{(k)}\tilde{T}_{2,j}^{\top}+\sum_{k=1}^{k_{x}}\tilde{T}_{3,k}{\bf S}^{(k)}\tilde{\Sigma}_{k}^{2,x\top}+\sum_{l=1}^{k_{y}}\tilde{\Sigma}_{l}^{1,y}{\bf S}^{(k)}\tilde{T}_{4,l}^{\top}\right)=\tilde{{\bf B}}^{(k)}$ (58)
where the projected operators are analogous to those defined in the rank-one case.
[…]
$\underbrace{\left[\sum_{i=1}^{\ell_{x}}{\tilde{A}_{1,i}}+\sum_{k=1}^{k_{x}}{\tilde{A}_{3,k}}\right]}_{\tilde{P}_{1}}\mathbf{S}^{(k)}+\mathbf{S}^{(k)}\underbrace{\left[\sum_{j=1}^{\ell_{y}}{\tilde{A}_{2,j}}+\sum_{l=1}^{k_{y}}{\tilde{A}_{4,l}}\right]}_{\tilde{P}_{2}^{\top}}=\tilde{{\bf B}}^{(k)}$ (59)
Using the operators $\tilde{P}_{1}$ and $\tilde{P}_{2}$, we define the ACS-preconditioner mapping:
${\bf S}\mapsto\tilde{P}_{1}{\bf S}+{\bf S}\tilde{P}_{2}^{T},$ (60)
Algorithm 5 solves the time-dependent variable advection-diffusion equation using DIRK methods.
Input: Initial condition matrices $U_{0}$, $V_{0}$, $S_{0}$;
Operators $P_{1}$, $P_{2}$, $\{T_{1,i}\}_{i=1}^{\ell_{x}}$, $\{T_{2,j}\}_{j=1}^{\ell_{y}}$, $\{T_{3,k}\}_{k=1}^{k_{x}}$, $\{T_{4,l}\}_{l=1}^{k_{y}}$;
Butcher table $\{a_{ij}\}$;
Time step size $\Delta t$;
Tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, ${\epsilon}_{\text{tol}}$;
Output: Updated bases $U_{1}$, $V_{1}$;
Truncated singular values $S_{1}$;
1 Compute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eq. (57);
2 Compute $l=2+\ell_{x}+\ell_{y}+k_{x}+k_{y}$;
3 Set $U^{(i)}=U_{0}$ and $V^{(i)}=V_{0}$ for $i=1,\dots,l$;
4 Set $U_{1}=U_{0}\quad V_{1}=V_{0}$;
5 while not converged do
6        // Step 1.
7        Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
8        $U_{1},\{U^{(1)}\cdots U^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}}\};\{U^{(1)}\cdots U^{(l)}\};U_{1};\epsilon_{\kappa}\right)$;
9        $V_{1},\{V^{(1)}\cdots V^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}}\};\{V^{(1)}\cdots V^{(l)}\};V_{1};\epsilon_{\kappa}\right)$;
10        // Step 2.
11        for $k=1$ to $s$ do
12               Compute $\tilde{{\bf B}}^{(k)}=\tilde{{\bf B}}^{(1)}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left({\bf S}^{(\ell)}-\tilde{{\bf B}}^{(\ell)}\right)$;
13               Solve the reduced equation (59) for ${\bf S}^{(k)}$ using GMRES with ACS preconditioner (60) to tolerance $\epsilon_{\texttt{GMRES}}$;
14               Compute and store $\frac{1}{a_{kk}}\left({\bf S}^{(k)}-\tilde{{\bf B}}^{(k)}\right)$ and proceed to the next stage;
15       Truncate $\{U_{1,\epsilon},S_{1,\epsilon},V_{1,\epsilon}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S^{(s)},V_{1}\})$ to tolerance $\epsilon$;
16        // Step 3.
17        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
18        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
19        // Step 4.
20        Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(-\tilde{B}^{(s)},S_{1,\epsilon},-\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1,\epsilon}\right)\right)R_{V}^{\top}\right\|$;
21
22       if $\|\mathbf{R}\|/\|\mathbf{F}_{0}\|\geq\epsilon_{\texttt{tol}}$ then
23               Return to Step 1 to augment bases further;
24
25       else
26               Set $U_{1}\leftarrow U_{1,\epsilon}$; $U_{1}\leftarrow V_{1,\epsilon}$; $S_{1}=S_{1,\epsilon}$;
27               Break;
28               Exit loop;
29
30
Algorithm 5 s-Stages Adaptive-Rank DIRK Integrator for the GSE.
Appendix D Three-Dimensional Formulation
We outline our approach to 3D next. In 3D, we replace the SVD algorithm with a tensor Tucker decomposition, specifically tailored for the three (and higher) dimension tensor-product grids.
D.1 Tucker Tensor Equations
We discretize the 3D time-dependent advection-diffusion equation on a three-dimensional tensor-product spatial grid with $N_{1}$, $N_{2}$, and $N_{3}$ grid points in the $x$, $y$, and $z$ directions, respectively. Here and in the following, we let ${\bf F}\in\mathbb{R}^{N_{1}\times N_{2}\times N_{3}}$ denote a three-dimensional tensor. The mode-$n$ product of ${\bf F}$ with a matrix $A_{n}\in\mathbb{R}^{I\times N_{n}}$, for $n=1,2,3$, is defined as:
${\bf F}\times_{1}A_{1}\in\mathbb{R}^{I\times N_{2}\times N_{3}},\,{\bf F}\times_{2}A_{2}\in\mathbb{R}^{N_{1}\times I\times N_{3}},\,{\bf F}\times_{3}A_{3}\in\mathbb{R}^{N_{1}\times N_{2}\times I}.$
The resulting tensor in each case has its $n$-th dimension replaced by the number of rows of $A_{n}$. The entries of the mode-$n$ product are computed as:
$({\bf F}\times_{n}A_{n})_{i_{1},\ldots,i_{n-1},j,i_{n+1},\ldots,i_{3}}=\sum_{i_{n}=1}^{N_{n}}{\bf F}_{i_{1},\ldots,i_{n},\ldots,i_{3}}\cdot(A_{n})_{j,i_{n}}.$
For a more thorough review of tensor operations, we refer the reader to [22]. The discretization of the 3D advection-diffusion equation using finite-differences in space combined with the backward Euler method in time, yields the following GSE:
${\bf R}_{\mathscr{L}}=\mathbf{F}_{1}-\Delta t\mathscr{L}({\bf F}_{1})-\mathbf{F}_{0}=\mathbf{0},$ (61)
where ${\bf F}_{0}$ denotes the initial tensor condition at time $t_{0}$, and ${\bf F}_{1}$ is the sought tensor solution at time $t_{1}$, and,
$\mathscr{L}:\mathbf{F}\mapsto$ $\sum_{i=1}^{\ell_{x}}\mathbf{F}\times_{1}T_{1,i}\times_{2}\Phi_{i}^{2,x}\times_{3}\Phi_{i}^{3,x}+\sum_{i=1}^{\ell_{y}}\mathbf{F}\times_{1}\Phi_{i}^{1,y}\times_{2}T_{2,i}\times_{3}\Phi_{i}^{3,y}+\sum_{i=1}^{\ell_{z}}\mathbf{F}\times_{1}\Phi_{i}^{1,z}\times_{2}\Phi_{i}^{2,z}\times_{3}T_{3,i}$
$+$ $\sum_{i=1}^{k_{x}}\mathbf{F}\times_{1}T_{4,i}\times_{2}\Sigma_{i}^{2,x}\times_{3}\Sigma_{i}^{3,x}+\sum_{i=1}^{k_{y}}\mathbf{F}\times_{1}\Sigma_{i}^{1,y}\times_{2}T_{5,i}\times_{3}\Sigma_{i}^{3,y}+\sum_{i=1}^{k_{z}}\mathbf{F}\times_{1}\Sigma_{i}^{1,z}\times_{2}\Sigma_{i}^{2,z}\times_{3}T_{6,i}.$
The operators $T$ correspond to the composition of advection and diffusion difference operators with diagonal matrices, analogous to those defined in Section 2. Similarly, the operators $\Phi$ and $\Sigma$ are analogous to the diagonal matrices for diffusion and advection coefficients, respectively, used for pointwise multiplication, as defined in Section 2. To obtain an averaged (approximate) mapping for preconditioning purposes, we define the average coefficients:
$\alpha_{i}^{2,x}$ $=\text{avg}(\Phi_{i}^{2,x}),\,\alpha_{i}^{3,x}$ $=\text{avg}(\Phi_{i}^{3,x}),\,\alpha_{i}^{1,y}$ $=\text{avg}(\Phi_{i}^{1,y}),\,\alpha_{i}^{3,y}$ $=\text{avg}(\Phi_{i}^{3,y}),\,\alpha_{i}^{1,z}$ $=\text{avg}(\Phi_{i}^{1,z}),\,\alpha_{i}^{2,z}$ $=\text{avg}(\Phi_{i}^{2,z}),$ (62)
$\gamma_{i}^{2,x}$ $=\text{avg}(\Sigma_{i}^{2,x}),\,\gamma_{i}^{3,x}$ $=\text{avg}(\Sigma_{i}^{3,x}),\,\gamma_{i}^{1,y}$ $=\text{avg}(\Sigma_{i}^{1,y}),\,\gamma_{i}^{3,y}$ $=\text{avg}(\Sigma_{i}^{3,y}),\,\gamma_{i}^{1,z}$ $=\text{avg}(\Sigma_{i}^{1,z}),\,\gamma_{i}^{2,z}$ $=\text{avg}(\Sigma_{i}^{2,z}).$
which yields the three-dimensional approximate operator $\tilde{\mathscr{L}}$ as:
[…]
$+\mathbf{F}\times_{2}\bigg{(}\sum_{j=1}^{\ell_{y}}\alpha_{j}^{1,y}\alpha_{j}^{3,y}T_{2,j}+\sum_{l=1}^{k_{y}}\gamma_{l}^{1,y}\gamma_{l}^{3,y}T_{5,l}\bigg{)}$
$+\mathbf{F}\times_{3}\bigg{(}\sum_{k=1}^{\ell_{z}}\alpha_{k}^{1,z}\alpha_{k}^{2,z}T_{3,k}+\sum_{l=1}^{k_{z}}\gamma_{l}^{1,z}\gamma_{l}^{2,z}T_{6,l}\bigg{)}.$ (63)
Using this approximate operator, and letting ${\mathcal{R}_{\text{ranks}}}=r_{x}+r_{y}+r_{z}+\ell_{x}+\ell_{y}+\ell_{z}$, we then define the averaged three-dimensional approximate SE of the form:
${\bf F}_{1}\times_{1}P_{1}+{\bf F}_{1}\times_{2}P_{2}+{\bf F}_{1}\times_{3}P_{3}={\bf F}_{0},$ (64)
with
$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\alpha_{i}^{2,x}\alpha_{i}^{3,x}T_{1,i}\right)}_{{A_{1,i}}}+\sum_{k=1}^{k_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\gamma_{k}^{2,x}\gamma_{k}^{3,x}T_{4,k}\right)}_{{A_{4,k}}}=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{4,k}},$
$P_{2}$ $=\sum_{j=1}^{\ell_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\alpha_{j}^{1,y}\alpha_{j}^{3,y}T_{2,j}\right)}_{A_{2,i}}+\sum_{l=1}^{k_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{l}^{1,y}\gamma_{l}^{3,y}T_{5,l}\right)}_{A_{5,i}}=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{5,l}},$
$P_{3}$ $=\sum_{j=1}^{\ell_{z}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{3}}-\Delta t\alpha_{k}^{1,z}\alpha_{k}^{2,z}T_{3,k}\right)}_{A_{3,i}}+\sum_{l=1}^{k_{z}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{l}^{1,z}\gamma_{l}^{2,z}T_{6,l}\right)}_{A_{6,i}}=\sum_{j=1}^{\ell_{z}}{A_{3,j}}+\sum_{l=1}^{k_{z}}{A_{6,l}}.$
These one-dimensional matrix operators, analogous to those in (26), are used (i) to construct xKrylov subspaces for the Tucker basis factors and (ii) to precondition the solution of the three-dimensional GSE arising from the Galerkin condition on the residual in Eq. (61).
D.2 GMRES-ACS Solution of the Reduced GSE in 3D
We assume the existence of a low-rank Tucker factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}\times N_{3}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}\times N_{3}}$ at time $t^{(1)}=t^{(0)}+\Delta t$, also in the low-rank Tucker format. This is represented as:
[…]
where $U_{0}\in\mathbb{R}^{N_{1}\times r^{x}_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r^{y}_{0}}$, $W_{0}\in\mathbb{R}^{N_{3}\times r^{z}_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r^{x}_{1}}$, $V_{1}\in\mathbb{R}^{N_{2}\times r^{y}_{1}}$, and $W_{1}\in\mathbb{R}^{N_{3}\times r^{z}_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r^{x}_{0}\times r^{y}_{0}\times r^{z}_{0}}$ and $S_{1}\in\mathbb{R}^{r^{x}_{1}\times r^{y}_{1}\times r^{z}_{1}}$ are core tensors at the corresponding times. The bases $U_{1}$, $V_{1}$, and $W_{1}$ are constructed from orthonormalization of dimension-wise xKrylov, as discussed in 3.2
Using a Galerkin projection on the residual, $\mathbf{R}_{\mathscr{L}}\times_{1}U_{1}^{\top}\times_{2}V_{1}^{\top}\times_{3}W_{1}^{\top}$, we obtain a reduced-size GSE for the core tensor ${\bf S}_{1}$ as follows:
$\mathbf{S}_{1}-\Delta t\mathscr{L}(\mathbf{S}_{1})\times_{1}U_{1}^{\top}\times_{2}V_{1}^{\top}\times_{3}W_{1}^{\top}=\mathbf{S}_{0}\times_{1}U_{1}^{\top}U_{0}\times_{2}V_{1}^{\top}V_{0}\times_{3}W_{1}^{\top}W_{0}.$ (66)
This projected system has dimensions $r_{x}\times r_{y}\times r_{z}$, where $r_{x}$, $r_{y}$, and $r_{z}$ denote the dimensions of the truncated xKrylov subspaces in the $x$, $y$, and $z$ directions, respectively, and will be solved using the GMRES-ACS preconditioner.
Eq. (66) is solved iteratively using GMRES, with $\mathscr{L}$-operator application cast as a series of tensor-matrix multiplications, along with the ACS preconditioning operator given by:
$\mathcal{P}:\mathbf{S}\mapsto\mathbf{S}\times_{1}\tilde{P}_{1}+\mathbf{S}\times_{2}\tilde{P}_{2}+\mathbf{S}\times_{3}\tilde{P}_{3},$ (67)
where $\tilde{P}_{i}=P_{i}\times_{1}U_{1}^{\top}\times_{2}V_{1}^{\top}\times_{3}W_{1}^{\top}$ denotes the projection of operators in Eq. (64), analogous to those introduced in Section 3. This operator is used to precondition each GMRES iteration of the inner system solve. The solution of the projected SE has a computational cost of $\mathcal{O}(r^{4})$ [9, 41]. The tensor-times-matrix operations, executed using the ttm command in the Tensor Toolbox, incur a cost of $\mathcal{O}(r^{4})$. The total complexity in three dimensions is therefore $\mathcal{O}(Nr^{2}+r^{4})$.