### v3

# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Variable Coefficients

Abstract
We consider the adaptive-rank integration of multi-dimensional time-dependent advection-diffusion partial differential equations (PDEs) with variable coefficients. We employ a standard finite-difference method for spatial discretization coupled with high-order diagonally implicit Runge-Kutta temporal schemes. The discrete equation is a generalized Sylvester equation (GSE), which we solve with a projection-based adaptive-rank algorithm structured around two key strategies: (i) constructing dimension-wise subspaces using a novel atypical extended Krylov strategy, and (ii) efficiently solving the basis coefficient matrix with a preconditioned GMRES solver. The low-rank decomposition is performed in 2D using SVD and with high-order SVD (HOSVD) in 3D to represent the tensor in a compressed Tucker format. For $d$-dimensional problems (here, $d=2$ or $3$), the computational complexity and memory storage of the approach are found numerically to scale as $\mathcal{O}(N{r^{2}})+\mathcal{O}({r^{d+1}})$ and $\mathcal{O}(Nr)+\mathcal{O}(r^{d})$, respectively, with $N$ the one-dimensional resolution and $r$ the maximal rank during the Krylov iteration (which we find to be largely independent of $N$ on our numerical examples). We present numerical examples that illustrate the advertised properties of the algorithm.
Keywords: Adaptive-rank, extended Krylov subspaces, implicit Runge-Kutta integrators, advection-diffusion equations, Sylvester preconditioning, matrix equations.
1 Introduction
We propose an adaptive-rank numerical scheme for approximating solutions of multi-dimensional time-dependent variable-coefficient advection-diffusion equations. The goal is to reduce the computational complexity of high-dimensional PDE solvers by dynamically exploiting the low-rank structure of the time-dependent solution when it exists, building upon existing literature on low-rank PDE solvers [16, 13] and model order reduction (MOR) for solving high-dimensional linear and nonlinear systems [25, 16, 5, 43].
Two primary strategies in the literature exploit evolving low-rank structures for time-dependent PDEs: the dynamical low-rank (DLR) approximation [22, 28, 8, 11] and the step-and-truncate (SAT) method, which has both explicit [24, 10, 18, 17, 19] and implicit [38, 31, 14, 29, 30] variants. Recently, [13] provided an overview of these two approaches with a focus on time-dependent kinetic simulations. Our study employs the SAT approach, which is known for preserving the high-order spatial and temporal accuracy of traditional mesh-based solvers for time-dependent problems. Specifically, to capture the solution dynamics accurately, we apply a high-order diagonally implicit Runge-Kutta (DIRK) time discretization to the matrix differential equation arising from the semi-discretization of advection-diffusion equations on a tensor-product of 1D spatial grids. The backward Euler time discretization of a 2D constant-coefficient advection-diffusion equation yields a Sylvester equation (SE) of the form:
$A\mathbf{X}_{1}+\mathbf{X}_{1}B^{\top}=\mathbf{X}_{0},$ (1)
where $\mathbf{X}_{0},\mathbf{X}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ represent the solution snapshots at times $t_{0}$ and $t_{1}$, respectively, and $A$ and $B$ are coefficient matrices derived from the discretizations of advection and diffusion differential operators in two dimensions. Here $N_{1}$ and $N_{2}$ denote the number of grid points along each coordinate direction of the tensor product grid. Generalization to problems with variable advection and/or diffusion coefficients leads to the so-called generalized Sylvester equation (GSE) in matrix and Kronecker forms, respectively:
$\sum_{i=1}^{l}A_{i}\mathbf{X}_{1}{B}^{\top}_{i}=\mathbf{X}_{0}\qquad\iff\qquad\underbrace{\left(\sum_{i=1}^{l}{B}_{i}\otimes A_{i}\right)}_{\mathcal{A}}\text{vec}(\mathbf{X}_{1})=\text{vec}(\mathbf{X_{0}}),$ (2)
where $A_{i}$ and $B_{i}$ are coefficient matrices from discretization of the PDE operators, leveraging the separability of variable advection-diffusion coefficients. Higher-order DIRK schemes lead to SE and GSE in forms similar to (1)-(2), but with the right-hand side (RHS) term, $\mathbf{X}_{0}$, constructed from the previous DIRK stages. To exploit low-rank structure in matrices (for 2D) and tensors (for 3D and beyond), such RHS term can be efficiently compressed using SVD rounding in 2D, Tucker decomposition in 3D, and hierarchical Tucker or tensor train decomposition for problems beyond 3D [32, 27]. Ultimately, a successful development of efficient and robust solvers for the GSE (2) is of critical importance to enable adaptive-rank integrators that accurately capture the transient states of time-dependent problems.
Two broad families of algorithms are commonly employed to tackle GSEs (and SEs as a special case): projection-type methods [5, 20, 45, 36, 39, 25, 43, 41] and matrix-oriented classical Krylov methods [45, 35, 5, 34, 33, 26]. Projection-type schemes first construct low-dimensional subspaces that capture the dominant features of the solution. The GSE is then projected dimension-by-dimension via a Galerkin condition, yielding a much smaller GSE for the expansion coefficients that compose the bases together. Matrix-oriented Krylov iterations, in contrast, construct an approximation subspace by applying powers of the operator $\mathcal{A}$ in (2) to the vectorized RHS, but exploit the Kronecker structure so that the action of $\mathcal{A}$ is performed at the matrix level, and the full Kroneckerized operator is never formed explicitly. Inner products and other Krylov routines are likewise executed at the matrix-level, with low-rank truncation applied throughout to control memory. Although an appropriately chosen projection space can make a projection-type method formally equivalent to a matrix-oriented Krylov scheme, the projection-based framework offers additional freedom to design nonstandard subspaces that may accelerate convergence [45]. On the downside, projection methods are less amenable to precondition: their efficiency hinges on the efficacy of the chosen subspaces and on the accuracy with which the reduced problem is solved. Classical Krylov methods, in contrast, benefit from well-established preconditioning strategies but offer no flexibility to alter the search space, as the basis is typically generated by an Arnoldi or Lanczos-type process applied directly to $\mathcal{A}$. In summary, each approach has complementary strengths and weaknesses, and the choice between them should be guided by the structure of the particular GSE under investigation. Below we survey representative work in both directions.
In the context of solving SEs, projection-type methods based on Krylov subspace techniques have proven highly effective, particularly by exploiting the low-rank structure of PDE solutions [25, 42, 38, 6, 14]. These methods typically construct dimension-wise solution bases from Krylov or Krylov-like subspaces and apply a Galerkin projection onto the reduced subspace spanned by the outer product of dimension-wise Krylov subspaces [39, 25, 43]. To accelerate convergence, extended Krylov subspace methods (first proposed in [12] for the approximation of matrix functions) that incorporate $A$ and $A^{-1}$, as well as $B$ and $B^{-1}$, in each direction, respectively, were successfully applied to matrix equations in [42, 14] within a projection-based scheme termed Krylov-Plus-Inverted-Krylov (KPIK).
Applying projection-type methods to GSEs is much more challenging. The major difficulty lies in preventing the rapid growth of the solution approximation spaces due to the (possibly large) multi-term structure of the equation, which otherwise results in large reduced systems (of size $r^{d}$, where $r$ is the dimension of the per-direction Krylov subspace and $d$ is the number of dimensions). Such large reduced system demands efficient, tailored solution strategies. Generalized Lyapunov equations of the form ($A\mathbf{X}_{1}+\mathbf{X}_{1}A+\sum_{i=1}^{l}M_{i}X_{1}M_{i}=\mathbf{X}_{0}$), a special case of the GSE matrix equation (2), have received significant attention  [5, 20] exemplifying such challenges. In [5], projection spaces are built from extended Krylov subspaces from the operator $A$ and standard Krylov subspaces from the remaining operators ($M_{1},\cdots M_{l}$), but do not address how to solve the reduced system efficiently. In [20], for the case $l=1$, projection spaces are built from a rational Krylov from $A$ and inverted Krylov from $A+M_{1}$. The authors then use the Sherman-Woodbury formula to express the projected equation as a series of Lyapunov solves, relying on the availability of a low-rank decomposition of the Galerkin projected generalized operators in the system. The resulting complexity for the reduced system is $\mathcal{O}(r^{d+1})$ under the assumption of fast decaying singular values of the operators, but deteriorates to $\mathcal{O}(r^{3d})$ otherwise.
Low-rank matrix-oriented Krylov methods such as Conjugate Gradient (CG) or GMRES [5, 26, 33, 46] provide an alternative to projection-type solvers. In [5], several low-rank matrix-oriented Krylov methods, including CG and BiCGSTAB, are surveyed, and a comparative numerical study against projection-based methods is presented. The study shows that, across a few benchmark problems, projection methods (for the specific approximations spaces utilized) underperforms preconditioned BiCGSTAB Krylov method in terms of both convergence rate and memory usage. In contrast, Reference  [45] shows that projection-based methods can be computationally superior to the classical conjugate gradient method when an appropriate solution subspace is chosen, and that classical Krylov methods suffer from performance degradation due to loss of orthogonality from the low-rank truncation of the subspaces. Reference [34] confirms the loss of orthogonality issues in low-rank Krylov subspace methods, and proposes techniques to mitigate it. They further find that preconditioning is key to avoid substantial subspace growth in matrix-oriented low-rank Krylov methods.
Preconditioning is indeed key to low-rank Krylov solvers for GSEs. Reference [34] proposes an interesting preconditioner for GSEs stemming from variable advection–diffusion equations, which will be of particular relevance to the present study. By averaging the PDE coefficients—a strategy we term here the average-coefficient-Sylvester (ACS) approximation—they obtain a constant-coefficient preconditioner of an SE form. The GSE system is then solved using matrix-oriented GMRES (but without low-rank truncation), where the operator application exploits the Kronecker product structure of the GSE, and the preconditioner is applied using KPIK [42]. In another related work on anisotropic diffusion problems [29], a basis-update-and-Galerkin (BUG) DLR integrator is employed as a preconditioner for a low-rank matrix-oriented GMRES solver for the associated GSE. The BUG preconditioner evolves the basis and the coefficient matrix to provide effective initial guesses for both the basis vectors and coefficient updates in the GMRES iteration, demonstrating significant potential in reducing the GMRES iteration count and the maximal Krylov rank. Therefore, the performance of (low-rank) matrix-oriented Krylov methods critically depends on the availability of a cheap and suitable preconditioner and an efficient truncation mechanism to limit the iteration count and control subspace growth.
In this study, we build upon the advances above for SE and GSE solvers within the projection-based framework and develop an adaptive-rank implicit integrator for 2D and 3D time-dependent advection-diffusion equations with variable coefficients. The main innovation of the algorithm lies in the specific treatment of both the basis construction and coefficient-finding steps within a standard projection-based Krylov framework, in which we combine solution elements proposed elsewhere to produce a novel algorithm. Notably, from the projection-based framework, we borrow the idea of enriched atypical Krylov subspaces to ensure an efficient representation of the solution bases. From classical matrix-oriented Krylov methods, we adopt the idea of truncating the subspaces along the iterations to keep it as compact as possible, and the idea of preconditioning the GSE with a simpler equation that is spectrally close. We assemble these ingredients in the following way: For the basis-building step (which we will refer to as the outer iteration), we introduce a new Krylov subspace construction strategy that combines (i) inverted Krylov subspaces for each differentiation coefficient matrix, (ii) standard Krylov subspaces for diagonal matrices arising from separability of advection-diffusion coefficients, and (iii) extended Krylov subspaces for averaged-coefficient matrices formed from the ACS approximation. We control excessive growth of the rank of the so-constructed Krylov subspaces bases by a targeted truncation procedure using SVD-based QR and modified Gram-Schmidt orthogonalization. In the coefficient-finding step (which we refer to as the inner iteration), we follow a standard Galerkin projection and solve the associated system with GMRES, preconditioned by the SE resulting from the Galerkin projection of the ACS approximation. This preconditioner is expected to perform optimally because it is compact equivalent to the original system [1, 2] and therefore spectrally close. The key distinction of this approach from others in the literature is applying the ACS SE preconditioner to the reduced-system of the projection-based framework with $\mathcal{O}(r^{d+1})$, rather than to the full-scale system (as was done for instance in the matrix-oriented GMRES framework proposed in [35]).
These algorithmic innovations enable the method to achieve linear scaling with the one-dimensional resolution $N$, both for computational complexity [as $\mathcal{O}(Nr^{2})+\mathcal{O}(r^{d+1})$] and for memory storage [as $\mathcal{O}(Nr)+\mathcal{O}(r^{d})$], and allow us to perform 3D simulations of advection-diffusion problems with equivalent resolutions of $\mathcal{O}\left(10^{12}\right)$ degrees of freedom on a standard laptop, orders of magnitude faster than an optimally performing multigrid method scaling as $N^{d}\log(N)$ [21]. Additionally, as in [14], the adaptive low-rank integrator is amenable for use with DIRK methods for high-order temporal accuracy (up to third-order is considered in this study). All these advertised properties will be demonstrated numerically in the test examples.
The rest of this paper is organized as follows. In Section 2, we consider the case of rank-one advection and diffusion coefficients and derive the GSE that results from a finite-difference discretization in space and backward Euler in time. In Section 3, we introduce an adaptive-rank solution strategy for 2D problems, including the coefficient-averaged approximation operator, the construction of extended Krylov subspaces (which we abbreviate as xKrylov in the sequel), the preconditioning technique in solving the reduced system, and the low-rank computation of the residual norm. The extension of the technique to high-order temporal discretizations with multi-rank coefficients is provided in Appendix C, and for 3D problems in Appendix D. In Section 4, numerical results are presented to demonstrate the efficacy of the algorithm. Finally, we conclude and provide an outlook for future research in Section 5.
2 Problem Description and Generalized Sylvester Equation
For simplicity, we focus the discussion in this section on 2D, with the extension to 3D elaborated in Appendix D. We establish the following notation conventions. Lowercase Greek letters, such as $\phi(x,y)$ or $\sigma(x,y)$, denote continuous scalar fields. Uppercase bold italic Greek letters, such as $\bm{\mathit{\Sigma}}(x,y)$ or $\bm{\mathit{\Phi}}(x,y)$, are used for vector-valued and tensor-valued functions. Uppercase Greek letters, like $\Phi$ or $\Sigma$, represent discrete matrices arising from the discretization of continuous scalar functions or differential operators on a grid. Boldface uppercase letters, such as $\mathbf{F}$ or $\mathbf{S}$, represent unknown matrix solutions or matrix right-hand sides in matrix equations. The identity matrix of size $n$ is denoted by ${I}_{n}$.
We aim to solve the time-dependent advection-diffusion equation:
$\frac{\partial f}{\partial t}=\nabla_{v}\cdot(\bm{\mathit{\Phi}}\cdot\nabla f)-\nabla\cdot(\bm{\mathit{\Sigma}}f),$ (3)
where $(x,y)\in\Omega=[a_{1},b_{1}]\times[a_{2},b_{2}]$ and $\bm{\mathit{\Phi}}\in\mathbb{R}^{2\times 2}$ is an isotropic diffusion tensor:
$\bm{\mathit{\Phi}}(x,y)=\begin{bmatrix}\phi^{x}(x,y)&0\\ 0&\phi^{y}(x,y)\end{bmatrix}$,
and $\bm{\mathit{\Sigma}}\in\mathbb{R}^{2}$ is the advection vector, given by:
$\bm{\mathit{\Sigma}}(x,y)=\begin{bmatrix}\sigma^{x}(x,y)\\ \sigma^{y}(x,y)\end{bmatrix}$.
We assume that the diffusion and advection coefficients are separable with finite rank. Specifically, the diagonal diffusion coefficients, $\phi^{x}(x,y)$ and $\phi^{y}(x,y)$, are given by:
$\phi^{x}(x,y)=\sum_{i=1}^{\ell_{x}}\phi^{1,x}_{i}(x)\phi^{2,x}_{i}(y),\qquad\phi^{y}(x,y)=\sum_{i=1}^{\ell_{y}}\phi^{1,y}_{i}(x)\phi^{2,y}_{i}(y).$ (4)
Similarly, for the advection coefficients we have:
$\sigma^{x}(x,y)=\sum_{i=1}^{k_{x}}\sigma^{1,x}_{i}(x)\sigma^{2,x}_{i}(y),\qquad\sigma^{y}(x,y)=\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\sigma^{2,y}_{i}(y).$ (5)
Substituting these coefficients in their separable forms into Equation (3), we obtain:
$\frac{\partial f}{\partial t}$ $=\sum_{i=1}^{\ell_{x}}\frac{\partial}{\partial x}\left(\phi^{1,x}_{i}(x)\frac{\partial f(x,y)}{\partial x}\right)\phi^{2,x}_{i}(y)+\sum_{i=1}^{\ell_{y}}\phi^{1,y}_{i}(x)\frac{\partial}{\partial y}\left(\phi^{2,y}_{i}(y)\frac{\partial f(x,y)}{\partial y}\right)$ (6)
$-\sum_{i=1}^{k_{x}}\frac{\partial}{\partial x}\left(\sigma^{1,x}_{i}(x)f(x,y)\right)\sigma^{2,x}_{i}(y)-\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\frac{\partial}{\partial y}\left(\sigma^{2,y}_{i}(y)f(x,y)\right),$
where, in each term, the factor independent of the differentiation variable has been pulled out of the derivative. To simplify the presentation, we assume for now rank-one advection and diffusion coefficients:
$\phi^{x}=\phi^{1,x}(x)\phi^{2,x}(y),\qquad\phi^{y}=\phi^{1,y}(x)\phi^{2,y}(y),$
$\sigma^{x}=\sigma^{1,x}(x)\sigma^{2,x}(y),\qquad\sigma^{y}=\sigma^{1,y}(x)\sigma^{2,y}(y).$
This will be generalized to multiple ranks below. Equation (6) then simplifies to:
$\frac{\partial f}{\partial t}$ $={\frac{\partial}{\partial x}\left(\phi^{1,x}(x)\frac{\partial f(x,y)}{\partial x}\right)\phi^{2,x}(y)+\phi^{1,y}(x)\frac{\partial}{\partial y}\left(\phi^{2,y}(y)\frac{\partial f(x,y)}{\partial y}\right)}$ (7)
${-}{\frac{\partial}{\partial x}\left(\sigma^{1,x}(x)f(x,y)\right)\sigma^{2,x}(y){-}\sigma^{1,y}(x)\frac{\partial}{\partial y}\left(\sigma^{2,y}(y)f(x,y)\right)}.$
We discretize this equation in a two-dimensional tensor-product spatial grid with ${N}_{1}$ and ${N}_{2}$ grid points in the $x$ and $y$ directions, respectively. The grid points in the $x$ direction are denoted as $x_{i}$ for $i=0,\dots,N_{1}-1$, and similarly for the $y$ direction. We consider the point-wise multiplication of the advection and diffusion coefficients in the discrete by setting:
$\Phi_{1}$ $=\text{diag}(\phi^{1,y}(x_{i})),\Sigma_{1}=\text{diag}(\sigma^{1,y}(x_{i})),\qquad i=0,\cdots,N_{1}-1,$
$\Phi_{2}$ $=\text{diag}(\phi^{2,x}(y_{j})),\Sigma_{2}=\text{diag}(\sigma^{2,x}(y_{j})),\qquad j=0,\cdots,N_{2}-1.$
We approximate the spatial derivative with a second-order finite-difference scheme using a three-point stencil, and obtain the following matrix differential equation for $\mathbf{F}$, for which its $(i,j)$ entry $F_{i,j}$ corresponds to the approximation of $f(x_{i},y_{j})$:
$\frac{\partial\mathbf{F}}{\partial t}={D_{xx}^{\Phi}\mathbf{F}\Phi_{2}^{\top}+\Phi_{1}\mathbf{F}{D_{yy}^{\Phi}}^{\top}}-{D_{x}^{\Sigma}\mathbf{F}\Sigma_{2}^{\top}-\Sigma_{1}\mathbf{F}{D_{y}^{\Sigma}}^{\top}}.$ (8)
Here, $D_{xx}^{\Phi},D_{yy}^{\Phi},D_{x}^{\Sigma},D_{y}^{\Sigma}$ represent compositions of difference operators with pointwise multiplication operators. See Appendices A and B for more details. For notational simplicity, let:
$T_{1}=D_{xx}^{\Phi},\qquad T_{2}=D_{yy}^{\Phi},\qquad T_{3}=-D_{x}^{\Sigma},\qquad T_{4}=-D_{y}^{\Sigma}.$ (9)
Equation (8) can be written compactly as:
$\frac{\partial\mathbf{F}}{\partial t}=\mathscr{L}(\mathbf{F}),$ (10)
where the right-hand side induces the following mapping:
${\mathscr{L}}:\mathbf{F}\mapsto T_{1}\mathbf{F}\Phi_{2}^{\top}+\Phi_{1}\mathbf{F}T_{2}^{\top}+T_{3}\mathbf{F}\Sigma_{2}^{\top}+\Sigma_{1}\mathbf{F}T_{4}^{\top}.$ (11)
We consider first a backward Euler implicit treatment of the temporal component of the matrix differential equation (10), which gives the following discrete GSE:
$\mathbf{F}_{1}-\Delta t\mathscr{L}({\bf F}_{1})=\mathbf{F}_{0},$ (12)
and associated residual equation:
$\mathbf{R_{\mathscr{L}}}=\mathbf{F}_{1}-\Delta t\,\mathscr{L}(\mathbf{F}_{1})-\mathbf{F}_{0}.$ (13)
Iterative solvers update the solution $\mathbf{F}_{1}$ until the residual is sufficiently small under a suitable norm. In this work, we opt for Galerkin-projection-based iterative solvers, where the bases spanning $\mathbf{F}_{1}$ are updated by augmenting Krylov-like subspaces, as shown next.
3 Krylov-Based Adaptive-Rank Solver for Generalized Sylvester Equation
We describe next the projection-based framework for solving the GSE (12) [5, 20, 45, 36]. We assume a low-rank factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ at time $t^{(1)}=t^{(0)}+\Delta t$ also in the low-rank format, i.e.:
$\mathbf{F}_{0}=U_{0}S_{0}V_{0}^{\top}\xrightarrow{\text{evolve in time}}\mathbf{F}_{1}=U_{1}S_{1}V_{1}^{\top},$ (14)
where $U_{0}\in\mathbb{R}^{N_{1}\times r_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r_{1}}$, and $V_{1}\in\mathbb{R}^{N_{2}\times r_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$ and $S_{1}\in\mathbb{R}^{r_{1}\times r_{1}}$ are diagonal matrices with singular values at the corresponding times. Projection methods involve constructing bases $U_{1}$ and $V_{1}$ and then projecting the original equation onto them to determine the coefficients of $S_{1}$. Efficiently solving the GSE requires choosing suitable solution subspaces effectively to build $U_{1}$ and $V_{1}$ from $U_{0}$ and $V_{0}$, and solving the reduced equation for $S_{1}$ efficiently. To accomplish both, in the spirit of the work in [35] we leverage an approximate operator $\tilde{\mathscr{L}}$ (which is of the Sylvester form) to the original operator $\mathscr{L}$} in (11). In the following, we describe the construction of the approximated operator $\tilde{\mathscr{L}}$ in Section 3.1; we discuss the dimension-wise basis construction in Section 3.2; we present the reduced equation for the matrix of coefficients $S_{1}$ and introduce our particular implementation of the ACS preconditioner in Section 3.3, and outline the residual computation in Section 3.4. Finally, Section 3.5 extends the algorithm to the case of multi-rank coefficients.
3.1 Averaged-Coefficient Approximate Operator $\tilde{\mathscr{L}}$
We incorporate ideas from [35] and [14] to construct an approximate operator $\tilde{\mathscr{L}}$ and efficiently solve the time-dependent inhomogeneous advection-diffusion equation using a projection-based adaptive-rank integration framework. The approximated operator $\tilde{\mathscr{L}}$ serves two purposes: (i) contribute to constructing the solution bases $U_{1}$ and $V_{1}$ from xKrylov subspaces, and (ii) preconditioning the solution of the coefficient matrix $S_{1}$.
We seek an approximation to $\mathscr{L}$} that captures its spectral content and is easier to invert. As proposed in [35], the advection diagonal operators can be approximated by scaling an identity matrix, where the scaling factor is the average of the diagonal entries. In this work, we approximate not only the advection coefficients by a suitable constant multiple of the identity, but extend this strategy to the scalar diffusion coefficients by averaging them across the computational domain. Let $\Phi_{i}\approx{\alpha}_{i}I$ and $\Sigma_{i}\approx\gamma_{i}I$, where $\alpha_{i}$ and $\gamma_{i}$ are the averages of the diagonal entries of $\Phi_{i}$ and $\Sigma_{i}$, respectively. This approximation yields the following mapping:
${\tilde{\mathscr{L}}}:\mathbf{F}\mapsto\left(\alpha_{2}T_{1}+\gamma_{2}T_{3}\right)\mathbf{F}+\mathbf{F}\left(\alpha_{1}T_{2}+\gamma_{1}T_{4}\right)^{\top}.$ (15)
A backward Euler temporal discretization using (15) leads to the SE:
$(A_{1}+A_{3})\mathbf{F}_{1}+\mathbf{F}_{1}(A_{2}+A_{4})^{\top}=\mathbf{F}_{0},$ (16)
with:
$A_{1}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{2}T_{1}\right)\qquad A_{3}=\left(\frac{1}{4}I-\Delta t\,\gamma_{2}T_{3}\right),\qquad P_{1}=A_{1}+A_{3},$ (17)
$A_{2}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{1}T_{2}\right)\qquad A_{4}=\left(\frac{1}{4}I-\Delta t\,\gamma_{1}T_{4}\right),\qquad P_{2}=A_{2}+A_{4}.$ (18)
Equation (16) is an SE approximation to the GSE (12), which can be efficiently solved using low-rank approximations [14, 42, 40, 39]. It can be exploited as a preconditioner for a GMRES iterative solution of the GSE (12). This was done in [35] with $\mathcal{O}(N^{3})$ computational complexity. In this work, we extend this approach by performing a Galerkin projection of both the original operator $\mathscr{L}$ and its approximation $\tilde{\mathscr{L}}$ on the low-rank bases before solving the GSE with GMRES. This reduces the problem to a lower-dimensional subspace, where preconditioning can be applied more effectively on a smaller $d$-dimensional problem with $\mathcal{O}(r^{d+1})$ complexity.
3.2 Basis Construction
The construction of the the xKrylov subspaces per dimension at the $m$-th outer iteration is performed as:
$\kappa_{m}(M_{1},\cdots,M_{l},U)=\text{{\color[rgb]{0,0,0}Range}}\Big\{U,\underbrace{M_{1}U,\cdots,M_{l}U}_{\text{1st augmentation}},\ldots,\underbrace{M_{1}^{m+1}U,\cdots,M_{l}^{m+1}U}_{\text{mth augmentation}}\Big\},$ (19)
where $M_{1},\cdots,M_{l}$ are operators of interest and their inverses arising in the GSE, and $l$ is the number of operators used for each basis augmentation step. For the $M_{l}$ operators, we include $P_{1}$ and its inverse from the SE approximation to the GSE (16), the individual operators $A_{1}$ and $A_{3}$ and their inverses, as they may contain information not captured by the inverse of $P_{1}$, and the diagonal operators $\Phi_{1}$, $\Sigma_{1}$, $\Phi_{2}$, and $\Sigma_{2}$, which are averaged to obtain $\tilde{\mathscr{L}}$. Using these operators, the generalized Krylov subspace (for instance in the $x$-direction) considered is as follows:
$\kappa_{m}\left(M_{1},\cdots,M_{6},U_{0}\right),\qquad\text{where}$ $M_{1}=P_{1},$ $M_{2}=P_{1}^{-1},$ $M_{3}=A_{1}^{-1},$ $M_{4}=A_{3}^{-1},$ $M_{5}=\Phi_{1},$ $M_{6}=\Sigma_{1}.$
We have found that including all these matrices is important to capture the range of the solution.
Depending on the structure of the GSE (2), $l$ in (19) may be large, and the generated subspace vectors at each augmentation may exhibit linear dependencies. To limit subspace inflation, we employ an SVD-based truncation strategy (Algorithm 1), such that only basis vectors corresponding to significant singular values are retained. With this in mind, the construction of the dimension-wise xKrylov subspaces (19), following [42], is as follows. At each iteration $m$, Algorithm 2 augments the previously constructed basis $\mathcal{V}_{(m-1)}$ by applying powers of operators $M_{1}$ through $M_{l}$ to their corresponding basis blocks $V^{(1)},\dots,V^{(l)}$. Each block matrix $V^{(i)}$ may have a rank of at most $r_{0}$. We apply Algorithm 2 to orthogonalize the subspace and obtain a basis $\mathcal{U}_{(m)}\in\mathbb{R}^{N_{1}\times r^{(m)}_{x}}$ and $\mathcal{V}_{(m)}\in\mathbb{R}^{N_{2}\times r^{(m)}_{y}}$ in the $x$ and $y$ directions, respectively. For notational simplicity, we omit the calligraphic notation and subscript $(m)$ from now on, understanding that $U_{1}\in\mathbb{R}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{R}^{N_{2}\times r_{y}}$ refer to the bases of the Krylov subspaces at the $m$th iteration. The truncation step in line 2 of Algorithm 2 removes any linear dependencies within each block during the power iterations; line 2 applies a global truncation across all newly generated blocks, eliminating redundancies among them. As a result, the subspace is kept optimal in size.
Input: Krylov Subspace $\kappa_{m}$; Tolerance $\epsilon_{\kappa}$.
Output: Truncated bases Q{Q}.
1 Compute the reduced QR decomposition: $\{Q,R\}=\texttt{QR}\left(\kappa_{m}\right)$;
2 Perform reduced SVD decomposition: $Z_{1}\tilde{R}Z_{2}=\texttt{SVD}(R)$, where $\tilde{R}=\text{diag}(\sigma_{j})$ ;
Identify the last index $r$ such that $\sigma_{r}>\epsilon_{\kappa}$ ;
// Optional: Set Maximum Allowable Rank
3 Update the basis vectors: $Q\leftarrow QZ_{1}(:,1:r)$;
Algorithm 1 SVD-truncated QR, $\mathscr{T}_{\epsilon_{\kappa}}$
Input: Operators $M_{1},\dots,M_{l}$; augmentation basis blocks $\{V^{(1)},\dots,V^{(l)}\}$; previous Krylov basis ${\mathcal{V}}_{(m-1)}$; tolerance $\epsilon_{\kappa}$.
Output: Updated Krylov basis $\mathcal{V}_{(m)}$; basis blocks $\{V^{(1)},\dots,V^{(l)}\}$ for the next iteration.
1
2for $i=1,\dots,l$ do
3    $V^{(i)}\leftarrow M_{i}V^{(i)}$;
4    $\widetilde{V}^{(i)}\leftarrow\text{orthogonalize }V^{(i)}\text{ w.r.t. }V_{(m-1)}$ (e.g., modified Gram-Schmidt);
    $V^{(i)}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}(\widetilde{V}^{(i)})$ ;
    // truncate to $\epsilon_{\kappa}$ for next iteration using Algorithm 1
5
[…]
$V^{\prime}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}(V^{\prime})$ ;
// final orthogonalization and truncation using Algorithm 1
7 $\mathcal{V}_{(m)}\leftarrow[\mathcal{V}_{(m-1)},V^{\prime}]$;
Algorithm 2 Construction of xKrylov Subspace $\kappa_{m}^{\texttt{trunc}}$
We comment next on the expected performance of the xKrylov iteration. In principle, the number of xKrylov iterations may depend on the mesh size, $N$. For instance, [25, 4] theoretically show for the 2D and 3D Poisson equations that the dimension of the extended Krylov subspace (denoted by $r^{(m)}$ in the basis construction above) scales as $\mathcal{O}(\kappa^{1/4})\sim\mathcal{O}(N^{1/2})$ in 2D, and $(\kappa^{2/3})\sim\mathcal{O}(N^{4/3})$ in 3D, with $\kappa\sim N^{2}$ the condition number. While this provides an upper bound of the convergence rate (without considering solution smoothness), our numerical results suggest that it may be pessimistic in practice, with Krylov subspace dimensions remaining largely independent of grid resolution for moderate convergence tolerances and mesh refinements.
3.3 Solving for the Matrix of Coefficients, $S_{1}$: ACS-Preconditioned GMRES
With the orthonormal bases $U_{1}\in\mathbb{R}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{R}^{N_{2}\times r_{y}}$ at hand, we perform a Galerkin projection to derive a reduced GSE for the matrix $S_{1}\in\mathbb{R}^{r_{x}\times r_{y}}$. The reduced GSE for ${\bf S}_{1}$ arises from projecting the residual via the Galerkin condition $U_{1}^{\top}\mathbf{R}_{\mathscr{L}}V_{1}=0$, yielding:
${\bf S}_{1}-\Delta t\left(\tilde{T}_{1}{\bf S}_{1}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}{\bf S}_{1}\tilde{T}_{2}^{\top}+\tilde{T}_{3}{\bf S}_{1}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}{\bf S}_{1}\tilde{T}_{4}^{\top}\right)=\tilde{{\bf B}}_{1},\hskip 18.49988pt\text{with}\hskip 18.49988pt\tilde{{\bf B}}_{1}=(U_{1}^{\top}U_{0})S_{0}(V_{1}^{\top}V_{0})^{\top}.$ (20)
This matrix-based equation is equivalent to the classical Kronecker formulation:
$\underbrace{\left[I_{r_{x}r_{y}\times r_{x}r_{y}}-\Delta t\left(\tilde{\Phi}_{2}\otimes\tilde{T}_{1}+\tilde{T}_{2}\otimes\tilde{\Phi}_{1}+\tilde{\Sigma}_{2}\otimes\tilde{T}_{3}+\tilde{T}_{4}\otimes\tilde{\Sigma}_{1}\right)\right]}_{\mathcal{A}}\text{vec}({\bf S}_{1})=\text{vec}(\tilde{{\bf B}}_{1}).$ (21)
The projected operators are defined as follows:
$\tilde{T}_{1}$ $=U_{1}^{\top}T_{1}U_{1},\qquad\tilde{T}_{3}=U_{1}^{\top}T_{3}U_{1},\qquad\tilde{\Phi}_{1}=U_{1}^{\top}\Phi_{1}U_{1},\qquad\tilde{\Sigma}_{1}=U_{1}^{\top}\Sigma_{1}U_{1},$
$\tilde{T}_{2}$ $=V_{1}^{\top}T_{2}V_{1},\qquad\tilde{T}_{4}=V_{1}^{\top}T_{4}V_{1},\qquad\tilde{\Phi}_{2}=V_{1}^{\top}\Phi_{2}V_{1},\qquad\tilde{\Sigma}_{2}=V_{1}^{\top}\Sigma_{2}V_{1}.$
Solving the Kronecker formulation $\mathcal{A}\text{vec}({{\bf S}_{1}})=\text{vec}({\tilde{{\bf B}}_{1}})$ in (21) with a direct solver scales as $\mathcal{O}(r^{3d})$, with $r$ the maximum rank of all dimensions and $d$ the dimensionality, which may be too expensive even for moderate $r$-values. Instead, we solve this equation iteratively using preconditioned GMRES. We wrap standard GMRES around the Kronecker form of the GSE, but exploiting the matrix-equation form (20) for both the $\mathcal{A}$-times-vector operation and the preconditioning step as indicated below. The resulting overall complexity is $\mathcal{O}(r^{d+1})$. For the GMRES operator-vector product, given a vector $s_{k}=\text{vec}(\mathbf{S}_{k})$, the product of the matrix $\mathcal{A}$ times a vector $s_{k}$ is performed as follows:
  1. 1.
    Given $s_{k}$, reshape to $\mathbf{S}_{k}\in\mathbb{R}^{r_{x}\times r_{y}}$, such that $\text{vec}(\mathbf{S}_{k})=s_{k}$.
  2. 2.
    Compute $\hat{\mathbf{Z}}_{k}\leftarrow\mathbf{S}_{k}-\Delta t\left(\tilde{T}_{1}\mathbf{S}_{k}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}\mathbf{S}_{k}\tilde{T}_{2}^{\top}+\tilde{T}_{3}\mathbf{S}_{k}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}\mathbf{S}_{k}\tilde{T}_{4}^{\top}\right)$, with complexity $\mathcal{O}(r^{d+1})$.
  3. 3.
    Reshape $\hat{z}_{k}=\text{vec}(\hat{\mathbf{Z}}_{k})$.
For efficiency, GMRES requires preconditioning so that the eigenvalues of the preconditioned matrix are sufficiently clustered to deliver $r$-independent convergence rates  [7]. Here, we employ the ACS preconditioner, with $\mathcal{P}=I_{r_{y}}\otimes\tilde{P}_{1}+\tilde{P}_{2}\otimes I_{r_{x}}$ arising from the Galerkin projection of (16), with $\tilde{P}_{1}=U_{1}^{\top}P_{1}U_{1}$ and $\tilde{P}_{2}=V_{1}^{\top}P_{2}V_{1}$. The ACS preconditioner delivers optimal, rank-independent, superlinear convergence (all observed in our numerical experiments), owing to its connection to the theory of compact-equivalent operators for diffusion-like equations. [^1] The averaged-coefficient advection-diffusion equation is compact-equivalent to the variable-coefficient one, which are known to be effective preconditioners for Krylov methods [1, 2]. In particular, we consider left preconditioning as:
$\mathcal{P}^{-1}\mathcal{A}\text{vec}({\bf S}_{1})=\mathcal{P}^{-1}\text{vec}(\tilde{{\bf B}}_{1}).$ (22)
To compute $\mathcal{P}^{-1}\hat{z}_{k}=z_{k}$ for an arbitrary vector $\hat{z}_{k}$ during GMRES, we solve the corresponding SE:
$\tilde{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\tilde{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k},$
where $z_{k}=\text{vec}(\mathbf{Z}_{k})\in\mathbb{R}^{r_{x}r_{y}}$. Specific steps are:
  1. 1.
    Given $\hat{z}_{k}$, reshape to $\hat{\mathbf{Z}}_{k}\in\mathbb{R}^{r_{x}\times r_{y}}$, such that $\text{vec}(\hat{\mathbf{Z}}_{k})=\hat{z}_{k}$.
  2. 2.
    Solve $\mathcal{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\mathcal{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k}$ with complexity $\mathcal{O}(r^{d+1})$, e.g. with the Bartels-Stewart algorithm [15].
  3. 3.
    Reshape $z_{k}=\text{vec}(\mathbf{Z}_{k})$.
As a result, the $S$-matrix solve scales as $O(r^{d+1})$ overall, a great reduction in complexity compared to the $O(r^{3d})$ scaling of a naïve implementation of GMRES with the Kronecker form.
3.4 Truncation and Residual Computation
Once the solution ${\bf F}_{1}=U_{1}S_{1}V_{1}^{\top}$ is obtained, we apply truncation to control rank growth and reduce the residual computational load. Since $S_{1}$ resulting from the GMRES solve is not necessarily diagonal, we diagonalize it with SVD [19] [also scaling as $\mathcal{O}(r^{d+1})$] and discard singular values below a threshold $\epsilon$. The SVD truncation gives ${\bf F}_{1}=U_{1,\epsilon}S_{1,\epsilon}V_{1,\epsilon}^{\top}$, with $U_{1,\epsilon}\in\mathbb{R}^{N_{1}\times r_{1}}$, $V_{1,\epsilon}\in\mathbb{R}^{N_{2}\times r_{1}}$, and $S_{1,\epsilon}\in\mathbb{R}^{r_{1}\times r_{1}}$, with $r_{1}\leq\min(r_{x},r_{y})$ (Algorithm 3).
Following [14], we compute the residual using the identity:
$\|\mathbf{R_{{\mathscr{L}}}}\|$ $=\|\mathbf{F}_{1}-\Delta t\,\mathscr{L}(\mathbf{F}_{1})-\mathbf{F}_{0}\|$
$=\underbrace{\left\|R_{U}\begin{bmatrix}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}\\ \end{bmatrix}R_{V}^{\top}\right\|}_{\in\mathbb{R}^{(r_{x}+5r_{1})\times(r_{y}+5r_{1})}}$
$={\left\|R_{U}\begin{bmatrix}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&{S_{1,\epsilon}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,(I_{\mathcal{R}_{\text{ranks}}}\otimes{S_{1,\epsilon}})\end{bmatrix}R_{V}^{\top}\right\|}\leq\epsilon_{\texttt{tol}},$ (23)
where $\mathcal{R}_{\text{ranks}}={\ell_{x}+k_{x}+\ell_{y}+k_{y}}$ is the total rank number of the advection and diffusion coefficients, and $R_{U}$ and $R_{V}$ are the upper triangular matrices arising from the following Q-less QR decomposition:
[…]
The above residual result follows from the property that $U_{1}(U_{1})^{\top}{\bf F}_{0}V_{1}(V_{1})^{\top}={\bf F}_{0}$, with ${\bf F}_{0}=U_{0}S_{0}V_{0}^{\top}$ (where $U_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ and $V_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ have orthonormal columns), since $U_{1}$ and $V_{1}$ contain $U_{0}$ and $V_{0}$, respectively.
We note that, as discussed in [33] and analyzed in detail later in this study, forming, storing, and evaluating the norm of the residual tensor in (23) can be a very expensive operation when the number of GSE terms, $\mathcal{R}_{\text{ranks}}$, becomes large. This is so because the complexity of a naïve implementation of the residual evaluation scales as $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d+1})$ and the memory storage scales as $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d})$ (although the latter has been reduced in our implementation to $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{2})$ by careful algorithmic choreography). While this is not a pressing issue in our numerical experiments (which feature small-to-moderate values of $\mathcal{R}_{\text{ranks}}$), it could become a limiting factor in applications where $\mathcal{R}_{\text{ranks}}$ becomes large. In such instances, it may be necessary to use alternate strategies as proxy measures of convergence such as the difference between consecutive iterates [33], which is commensurate in cost with the rest of our algorithm.
Input: Bases $U$, $V$; matrix of coefficients $\bf S$}; Tolerance $\epsilon$.
Output: Truncated bases $\tilde{U}$, $\tilde{V}$; truncated singular values $\tilde{S}$.
1
2Perform reduced SVD decomposition: $T_{1}\tilde{S}T_{2}=\texttt{SVD}(S)$, where $\tilde{S}=\text{diag}(\sigma_{j})$;
3
4Identify the last index $r_{1}$ such that $\sigma_{r_{1}}/\|S\|>\epsilon$;
5
6Update matrix of singular values: $\tilde{S}\leftarrow\tilde{S}(1:r_{1},1:r_{1})$;
7
8Update left singular vectors: $\tilde{U}\leftarrow UT_{1}(:,1:r_{1})$;
9
10Update the right singular vectors: $\tilde{V}\leftarrow VT_{2}(:,1:r_{1})$;
11
Algorithm 3 Truncated SVD, $\mathcal{T}_{\epsilon}$
3.5 Generalization to Multi-rank Coefficients
We generalize next the adaptive-rank approach to the multi-rank coefficient case in (4) and (5), using backward Euler for now. Discretization of the advection-diffusion equation in this case yields a multi-term GSE of the form:
$\mathbf{F}_{1}-\Delta t\left(\underbrace{\sum_{i=1}^{\ell_{x}}T_{1,i}\mathbf{F}_{1}{\Phi_{i}^{2,x}}^{\top}+\sum_{j=1}^{\ell_{y}}\Phi_{j}^{1,y}\mathbf{F}_{1}T_{2,j}^{\top}+\sum_{k=1}^{k_{x}}T_{3,k}\mathbf{F}_{1}{\Sigma_{k}^{2,x}}^{\top}+\sum_{l=1}^{k_{y}}\Sigma_{l}^{1,y}\mathbf{F}_{1}T_{4,l}^{\top}}_{\mathscr{L}(\mathbf{F}_{1})}\right)={\bf F}_{0}.$ (25)
As before, we find $\mathscr$$\tilde${L}} by averaging the diagonal terms $(\Phi_{i}^{2,x},\Phi_{j}^{1,y},\Sigma_{k}^{2,x},\Sigma_{l}^{1,y})$ from (25) to arrive at the approximated SE:
${P_{1}}\mathbf{F}_{1}+\mathbf{F}_{1}{P_{2}^{\top}}=\mathbf{F}_{0},$ (26)
where
$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{3,k}}=\sum_{i=1}^{\ell_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\alpha_{x,i}T_{1,i}\right)}+\sum_{k=1}^{k_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\gamma_{x,k}T_{3,k}\right)},$ (27)
$P_{2}$ $=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{4,l}}=\sum_{j=1}^{\ell_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\alpha_{y,j}T_{2,j}\right)}+\sum_{l=1}^{k_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{y,l}T_{4,l}\right)}.$ (28)
Also as before, the bases $U_{1}$ and $V_{1}$ are constructed from orthonormalization of dimension-wise xKrylov from equations (25) and (26) as:
$U_{1}$ $=\kappa^{\texttt{trunc}}_{m}\left(P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}},U_{0}\right),$ (29)
$V_{1}$ $=\kappa^{\texttt{trunc}}_{m}\left(P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}},V_{0}\right).$ (30)
In the multi-rank setting, after a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$, we obtain a GSE similar to (20) for the matrix of coefficients ${\bf S}_{1}$:
${\bf S}_{1}-\Delta_{t}\left(\sum_{i=1}^{\ell_{x}}\tilde{T}_{1,i}{\bf S}_{1}\tilde{\Phi}_{i}^{2,x\top}+\sum_{j=1}^{\ell_{y}}\tilde{\Phi}_{j}^{1,y}{\bf S}_{1}\tilde{T}_{2,j}^{\top}+\sum_{k=1}^{k_{x}}\tilde{T}_{3,k}{\bf S}_{1}\tilde{\Sigma}_{k}^{2,x\top}+\sum_{l=1}^{k_{y}}\tilde{\Sigma}_{l}^{1,y}{\bf S}_{1}\tilde{T}_{4,l}^{\top}\right)=\tilde{{\bf B}}_{1},$ (31)
where the projected operators are analogous to those defined in the rank-one case. We solve (31) for the ${\bf S}_{1}$ matrix using preconditioned GMRES. The preconditioner is constructed by another Galerkin projection to find the projected SE:
$\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}=\tilde{{\bf B}}_{1},$ (32)
with:
${\tilde{P}_{1}}$ $=\sum_{i=1}^{\ell_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{x}}-\Delta_{t}\alpha_{x,i}\tilde{T}_{1,i}\right)}_{\tilde{A}_{1,i}}+\sum_{k=1}^{k_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{x}}-\Delta_{t}\gamma_{x,k}\tilde{T}_{3,k}\right)}_{\tilde{A}_{3,k}},$
$\tilde{P}_{2}$ $=\sum_{j=1}^{\ell_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{y}}-\Delta_{t}\alpha_{y,j}\tilde{T}_{2,j}\right)}_{\tilde{A}_{2,j}}+\sum_{l=1}^{k_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{y}}-\Delta_{t}\gamma_{y,l}\tilde{T}_{4,l}\right)}_{\tilde{A}_{4,l}}.$
This yields the multi-rank-coefficients ACS preconditioner operator:
$\mathscr{P}:\mathbf{S}\mapsto\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}.$ (33)
We summarize the pseudo-algorithm for multi-rank coefficients in Algorithm 4. The tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, $\epsilon_{\texttt{tol}}$ in the algorithm are set as follows. First, we choose the residual tolerance for xKrylov outer iterations, $\epsilon_{\texttt{tol}}$. To prevent rapid growth of the xKrylov subspaces, we choose the Krylov basis truncation threshold as $\epsilon_{\kappa}\leq\epsilon_{\texttt{tol}}$, typically two to three orders of magnitude smaller. The GMRES-ACS tolerance for the inner reduced system solve, $\epsilon_{\texttt{GMRES}}$, is initially set to be less than $\epsilon_{\texttt{tol}}$ for the first outer-basis iteration. It is subsequently adjusted to be two to three orders of magnitude smaller than the relative residual error from the previous xKrylov outer iteration. This adjustment ensures that errors from the inner solve do not pollute the accuracy of the outer solve. The SVD truncation threshold $\epsilon$ is chosen to be one to three orders of magnitude smaller than the tolerance $\epsilon_{\texttt{tol}}$. Since solution truncation occurs before residual evaluations, the truncation error in the solution should remain below the residual tolerance to prevent residual pollution by over-truncation.
// This algorithm solves Eq. (25) for $\mathbf{F}_{1}=U_{1}S_{1}V_{1}^{\top}$, where $\mathbf{F}_{0}=U_{0}S_{0}V_{0}^{\top}$
Input: Initial condition matrices $U_{0}$, $V_{0}$, $S_{0}$;
Operators $P_{1}$, $P_{2}$, $\{T_{1,i}\}_{i=1}^{\ell_{x}}$, $\{T_{2,j}\}_{j=1}^{\ell_{y}}$, $\{T_{3,k}\}_{k=1}^{k_{x}}$, $\{T_{4,l}\}_{l=1}^{k_{y}}$;
Tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, $\epsilon_{\texttt{tol}}$;
Output: Updated bases $U_{1}$, $V_{1}$;
Truncated singular values $S_{1}$
1ex
1Compute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eqs. (27), (28);
2 Compute $l=2+\ell_{x}+\ell_{y}+k_{x}+k_{y}$;
3 Set $U^{(i)}=U_{0}$ and $V^{(i)}=V_{0}$ for $i=1,\dots,l$;
4 Set $U_{1}=U_{0}\qquad V_{1}=V_{0}$;
5 while not converged do
6    // Step K1.
7    Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
8    $U_{1},\{U^{(1)}\cdots U^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}}\};\{U^{(1)}\cdots U^{(l)}\};U_{1};\epsilon_{\kappa}\right)$;
9    $V_{1},\{V^{(1)}\cdots V^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}}\};\{V^{(1)}\cdots V^{(l)}\};V_{1};\epsilon_{\kappa}\right)$;
10
11   // Step K2.
12    Solve the reduced GSE (31) for ${\bf S}_{1}$ using GMRES-ACS (33) to tolerance $\epsilon_{\texttt{GMRES}}$;
13    // Step K3.
14    Truncate $\{U_{1,\epsilon},S_{1,\epsilon},V_{1,\epsilon}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}\})$ to tolerance $\epsilon$ (Alg. 3);
15
16   // Step K4.
17    Compute $\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
18    $\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
19
20   Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(-\tilde{B}_{1},S_{1,\epsilon},-\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1,\epsilon}\right)\right)R_{V}^{\top}\right\|$;
21
22   if $\|\mathbf{R}\|/\|\mathbf{F}_{0}\|\geq\epsilon_{\texttt{tol}}$ then
23        Reject solution and cycle to augment bases further;
24
25   else
26        Set $U_{1}\leftarrow U_{1,\epsilon}$; $U_{1}\leftarrow V_{1,\epsilon}$; $S_{1}=S_{1,\epsilon}$;
27        Break;
28        Exit loop;
29
30
Algorithm 4 Backward Euler Adaptive-Rank Integrator for the GSE.
3.6 Analysis of Computational and Memory Complexity
We analyze next the computational complexity and memory storage of Algorithm 4 for backward Euler. For conciseness, we consider the same spatial resolution per dimension, i.e., $N=N_{1}=N_{2}$. Starting from an initial condition $U_{0}S_{0}V_{0}^{\top}$, where $U_{0}\in\mathbb{R}^{N\times r_{0}}$, $V_{0}\in\mathbb{R}^{N\times r_{0}}$, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$, the $m$-th xKrylov iteration generates $U_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, $V_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, and $S_{1}\in\mathbb{R}^{r_{\kappa}\times r_{\kappa}}$, where $r_{\kappa}=(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0}+r^{(m-1)}$. Here, $r^{(m-1)}$ denotes the basis size from the previous xKrylov iteration (with $r^{(0)}=r_{0}$), and $r_{0}^{\ast}\leq r_{0}$ is the (possibly truncated) initial rank produced at line (2) of Algorithm 2. While not used in this work, one may impose a maximum allowable rank $r_{\max}$ in the orthogonalization step of Algorithm 1 to prevent the subspace from exceeding available computational resources. Using SVD-truncated QR, this dimension is then reduced to smaller ranks, $r_{x}$ and $r_{y}$. Let $\tilde{r}=\max(r_{x},r_{y})$. The matrix $S_{1}$ from the $m$-th inner iteration solve, initially of size $r_{x}\times r_{y}$, is truncated to rank $r_{1}\leq\min(r_{x},r_{y})$ using Algorithm 3. To simplify the complexity analysis, let $r=\max(r_{0},r^{(m-1)},\tilde{r},(\mathcal{R}_{\text{ranks}}+1)r_{1})$, which we consider independent of the mesh-refinement $N$ for practical purposes. The computational complexity of each step is then estimated as follows:
Step K1
Constructing the generalized Krylov basis as outlined in lines (4) and (4) of Algorithm 4 requires performing the operations $A_{1,1:r_{x}}^{-1}U_{0}$ and $A_{3,1:k_{x}}^{-1}U_{0}$, and $A_{2,1:r_{y}}^{-1}V_{0}$ and $A_{4,1:k_{y}}^{-1}V_{0}$, which require solving the systems $AX=U$ and $AX=V$ (subscripts have been omitted for simplicity). With 1D finite-differences, $A$ has a tridiagonal structure, which can be efficiently factorized using direct-solver techniques such as the Thomas algorithm, with $\mathcal{O}(N)$ complexity per column of $X$ [37]. The same complexity holds for applying $P_{1}^{-1}$ and $P_{2}^{-1}$. Additionally, applying $P_{1}U_{0}$, $\Phi_{1,3}U_{0}$, $\Sigma_{1,3}U_{0}$, $P_{2}V_{0}$, $\Phi_{2,4}V_{0}$, and $\Sigma_{2,4}V_{0}$, can be done with $\mathcal{O}(Nr^{\ast}_{0})$ computational complexity for sparse diagonal and tridiagonal matrices, leading to a total complexity of $\mathcal{O}(N(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0})$ to augment the xKrylov subspaces. The use of Algorithm 2 to compute the orthonormal truncated bases $U_{1}$ and $V_{1}$ incurs a complexity of $\mathcal{O}\left(N\,\max\left(r_{\kappa}\,r^{(m-1)},\;r_{\kappa}^{2}\right)\right)\sim\mathcal{O}(Nr^{2})$. The memory complexity to store the truncated bases is $\mathcal{O}\left(N\,\left(r^{(m-1)}+r_{\kappa}\right)\right)\sim\mathcal{O}(Nr)$. Generally, we expect $r$ to be much smaller and independent of $N$ when the solution is sufficiently smooth and exhibits a low-rank structure.
Step K2
The application of the reduced GSE operator using matrix-matrix multiplication and summation in the Krylov method (GMRES) features a complexity of $\mathcal{O}\left(\tilde{r}^{d+1}\right)$. Applying the ACS preconditioner requires solving an SE of size $r_{x}\times r_{y}$, which can be achieved using the Bartels-Stewart algorithm [15] with a complexity of $\mathcal{O}(\tilde{r}^{d+1})\sim\mathcal{O}(r^{d+1})$. Hence, the computation of the $S$-step has $\mathcal{O}(r^{d+1})$   complexity. The memory complexity to store the S matrix/tensor is $\mathcal{O}\left(\tilde{r}^{d}\right)\sim\mathcal{O}(r^{d})$.
Step K3
The SVD decomposition of $\widetilde{S}$} in line (3) of Algorithm 3 is of $\mathcal{O}(\tilde{r}^{d+1})$ complexity. Updating the bases in lines (3) and (3) of Algorithm 3 requires $\mathcal{O}(N\tilde{r}{r_{1}})$ operations, where ${r_{1}}\leq\min(r_{x},r_{y})$ is the post-truncation rank of the new solution from Algorithm 3.
Step K4
The QR decomposition to obtain $R_{U}$ and $R_{V}$ in lines (11)–(4) of Algorithm 4 has a complexity of $\mathcal{O}\left(N\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{2}\right)\sim\mathcal{O}(Nr^{2})$. The residual computation in line (4) requires $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d+1}\right)\sim\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d+1}r^{d+1})$ due to matrix multiplications, with an additional $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d}\right)\sim\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d}r^{d})$ for the Frobenius norm. The memory storage requirements for the residual tensor may be quite prohibitive if stored fully [it would scale as $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d}r^{d})$]. Instead, we store only the $R_{U}$ and $R_{V}$ factor matrices and the $S$ matrix, and build the residual for the norm calculation on the fly still with computational complexity $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d+1}r^{d+1})$. Storing $R_{U}$ and $R_{V}$ requires $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{2}\right)\sim\mathcal{O}(\mathcal{R}_{\text{ranks}}^{2}r^{2})$, and the residual tensor blocks in Eq. 23 require $\mathcal{O}(r^{d})$, resulting in a total storage of $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{2}r^{2})+\mathcal{O}(r^{d})$. Nevertheless, it is clear from this analysis that the residual evaluation may become prohibitively expensive when $\mathcal{R}_{\text{ranks}}\gg 1$. In such situations, the difference between successive solution iterates may be used as a alternate measure of convergence [33], commensurate in cost to the rest of the algorithm.
It follows that the overall complexity of the algorithm scales as $\mathcal{O}(Nr^{2})+\mathcal{O}(r^{d+1})$, and the memory complexity scales to leading order as $\mathcal{O}(Nr)+\mathcal{O}(r^{d})$, with $d$ the dimensionality. Both claims will be numerically verified in Section 4.
4 Numerical Experiments
We demonstrate the overall efficacy of the algorithm with several 2D and 3D numerical examples. We assume homogeneous Dirichlet boundary conditions for simplicity; other boundary conditions (e.g., periodic, Neumann, etc.) can be easily incorporated in the formulation through the operators $T_{1:4}$ (9). All simulations are performed using MATLAB. For efficient computational handling in 3D, we employ tools from the Tensor Toolbox [3] and the Htucker package [27]. High-order temporal integration using DIRK is used throughout the numerical section. Details on the the DIRK scheme implementation in our adaptive-rank algorithm are provided in Appendix C. Throughout this section, we denote the diffusion Courant number and the advection Courant-Friedrichs-Lewy (CFL) number by $\lambda_{D}$ and $\lambda_{A}$, respectively, defined as:
$\lambda_{D}=\frac{\Delta t\phi_{\text{max}}}{\Delta x^{2}},\hskip 18.49988pt\lambda_{A}=\frac{\Delta t\sigma_{\text{max}}}{\Delta x},$ (34)
where $\phi_{\text{max}}$ and $\sigma_{\text{max}}$ are the maximum of the diffusion and advection coefficients, respectively. The numbers $\lambda_{D}$ and $\lambda_{A}$ are appropriate measures of the maximum eigenvalue of the associated GSE, and by extension of its condition number since the minimum eigenvalue is of $\mathcal{O}(1)$. For all simulations presented in this section, we did not need to set a maximum number of outer xKrylov iterations because the algorithm consistently converged within a small number of iterations.
Example 4.1: 3D Advection-Diffusion
We consider a rank-1 variable advection coefficient in a 3D domain $x,y,z\in[-1,1]$:
$\sigma^{x}(x,y,z)$ $=\sigma^{1}_{1}(x)\sigma^{2}_{1}(y)\sigma^{3}_{1}(z),$ $\sigma^{1}_{1}(x)$ $=1-x^{2},$ $\sigma^{2}_{1}(y)$ $=2y,$ $\sigma^{3}_{1}(z)$ $=-2z$
$\sigma^{y}(x,y,z)$ $=\sigma^{1}_{2}(x)\sigma^{2}_{2}(y)\sigma^{3}_{2}(z),$ $\sigma^{1}_{2}(x)$ $=-2x,$ $\sigma^{2}_{2}(y)$ $=1-y^{2},$ $\sigma^{3}_{2}(z)$ $=2z$
$\sigma^{z}(x,y,z)$ $=\sigma^{1}_{3}(x)\sigma^{2}_{3}(y)\sigma^{3}_{3}(z),$ $\sigma^{1}_{2}(x)$ $=4x,$ $\sigma^{2}_{2}(y)$ $=2y,$ $\sigma^{3}_{3}(z)$ $=1-z^{2}.$
Note that this flow is divergence-free and therefore a pure vorticity flow, which results in interesting flow dynamics such as differential rotation. For diffusion, we specify a rank-3 variable coefficient as follows:
$\phi^{x}(x,y,z)=\phi^{y}(x,y,z)=\sum_{i=1}^{3}\phi^{1}_{i}(x)\phi^{2}_{i}(y)\phi^{3}_{i}(z),$
${\phi}^{1}_{1}(x)$ $=\exp\left(-\left(x-0.3\sin(x)\right)^{2}\right),$ ${\phi}^{2}_{1}(y)$ $=\exp\left(-\left(y-0.3\cos(y)\right)^{2}\right),$ ${\phi}^{3}_{1}(z)$ $=\exp\left(-\left(y-0.3\sin(z)\right)^{2}\right),$
${\phi}^{1}_{2}(x)$ $=\exp\left(-\left(x-0.6\sin(\pi x)\right)^{2}\right),$ ${\phi}^{2}_{2}(y)$ $=\exp\left(-\left(y-0.6\sin(\pi y)\right)^{2}\right),$ ${\phi}^{3}_{2}(z)$ $=\exp\left(-\left(y-0.6\sin(\pi z)\right)^{2}\right),$
${\phi}^{1}_{3}(x)$ $=\exp\left(-\left(x-0.6\sin(2\pi x)\right)^{2}\right),$ ${\phi}^{2}_{3}(y)$ $=\exp\left(-\left(y-0.6\sin(2\pi y)\right)^{2}\right)$ ${\phi}^{3}_{3}(z)$ $=\exp\left(-\left(y-0.6\sin(2\pi z)\right)^{2}\right).$
These diffusion coefficients are relatively balanced and therefore mostly isotropic. We will investigate anisotropic diffusion in future work. The initial condition is a rank-2 Gaussian:
$f(x,y,z,t=0)$ $=0.5\exp\left(-400\left((x-0.3)^{2}+(y-0.35)^{2}+(z-0.2)^{2}\right)\right)$
$+0.8\exp\left(-400\left((x-0.65)^{2}+(y-0.5)^{2}+(z-0.55)^{2}\right)\right).$
We use a spatial grid with $N=300$ in all dimensions, and the time-step $\Delta t$ is set to $\lambda_{D}\Delta x^{2}$, with $\lambda_{D}$ ranging from $140$ to $1125$. We test the adaptive-rank algorithm using three different integrators: backward Euler (BE), DIRK2, and DIRK3, as detailed in Tables 4 and 4 in Appendix C. The residual tolerance is $\epsilon_{\texttt{tol}}=10^{-4}$ for BE, DIRK2, and DIRK3, the basis truncation threshold is $\epsilon_{\kappa}=10^{-6}$, the SVD truncation threshold is $\epsilon=10^{-6}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-6}$.
We investigate numerically the following: (1) high-order temporal convergence, (2) effectiveness of the xKrylov iterations in capturing the range of solution, (3) performance of the ACS preconditioner to solve the reduced GSE, and (4) verification of the overall computational complexity.
High-Order Temporal Convergence: Figure 1(a) illustrates the $L_{1}$ error norm vs time stepping size (characterized by $\lambda_{D}$) by showcasing the performance of the adaptive-rank integrators for BE, DIRK2, and DIRK3. The expected asymptotic order of convergence of the temporal error for the adaptive-rank integrators, depicted by lines with markers, is observed. Additionally, the algorithm adeptly captures the rank evolution of the solution in the $x$, $y$, and $z$ directions, as illustrated in Figure 1(b).
Figure 1: Example 4.1. (a) Temporal convergence study of BE, DIRK2, and DIRK3 schemes. (b) Rank evolution of the solution in the $x$, $y$, and $z$ directions over time for $\lambda_{D}=140$.
Effectiveness of the xKrylov Iterations: We evaluate the xKrylov subspace dimension required to achieve tolerances of $\epsilon_{\texttt{tol}}=10^{-4},10^{-5},10^{-6}$. We perform a single time-step integration with a fixed step size $\Delta t=10^{-3}$ with grid refinements in $N$. The convergence bound for the extended Krylov method depends on the condition number of the advection-diffusion operator, $\kappa$. Assuming that the diffusion operator dominates the condition number, and that the variable coefficients perturb its eigenvalues only by factors of order unity, $\lambda_{D}$ provides a good estimate of the condition number (i.e., $\kappa=\mathcal{O}(\lambda_{D})$ in the case of the diffusion equation), and is therefore a good figure of merit to assess the convergence of the xKrylov subspaces. As reported in [25], the convergence factor for the extended Krylov subspace method applied to an equation with Kronecker product structure is $\kappa^{1/4}$ in 2D and $\kappa^{2/3}$ in 3D.
To isolate the effect of $\lambda_{D}$ on the convergence of xKrylov, we mitigate other potential influencing factors, such as excessive truncation of the Krylov subspace, insufficient convergence of the GMRES-ACS solver, and truncation errors in the constructed solution, by setting related thresholds to $\epsilon=\epsilon_{\texttt{GMRES}}=\epsilon_{\kappa}=10^{-10}$. Figure 2 illustrates the xKrylov subspace dimension required for convergence as a function of the diffusion Courant number $\lambda_{D}$ for a fixed $\Delta t$ and varying the spatial 1D resolution, $N$. Our results suggest that, for a fixed tolerance, the convergence of the Krylov subspaces remains largely insensitive to grid refinement across a broad range of $\lambda_{D}$. For extremely fine 1D mesh resolutions, $N\sim 10^{4}$, beyond typical practical computational needs, the convergence rate of xKrylov scales as $\lambda_{D}^{2/3}$, consistently with the theoretically expected $\kappa^{2/3}$ scaling in 3D [25, 4] .
Figure 2: Example 4.1. Convergence study using the xKrylov subspace method with DIRK3 adaptive-rank integrator for several outer basis tolerances, $\epsilon_{tol}$.
Performance of ACS-Preconditioner: Ill-conditioning for the projected system stems from both grid resolution (the projected system becomes more ill-conditioned as we refine the mesh) and rank growth (the projected system becomes larger in size). For effectiveness, the preconditioner must be able to deal with both sources of ill-conditioning. We consider first the scaling of the ACS-preconditioned GMRES inner solve with respect to the 1D resolution $N$. For this, we compare performance with and without the ACS preconditioner at the first time step ($\Delta t=0.01$) and the third outer-basis augmentation (i.e., keeping the subspace size constant, and therefore the rank), varying the mesh size. Figure 3 illustrates the GMRES residual history over iterations for ACS-preconditioned and unpreconditioned cases. The results confirm that ACS-preconditioned GMRES delivers mesh-independent and super-linear convergence rates, which is not the case for unpreconditioned GMRES.
Figure 3: Example 4.1. Log-log plots of GMRES residuals versus iterations, with and without preconditioning. Simulations are conducted for a single time step with $\Delta t=0.01$ at the third outer-basis augmentation, using grid resolutions $N_{x}=N_{y}=N_{z}=200$ in (a), $1000$ in (b), and $2500$ in (c).
We consider next the performance of the GMRES inner solve with respect to the solution rank $r$. For this test, the grid size is fixed at $N=1000$ in 3D, and a single time step $\Delta t=0.01$ is used. Figure 4 (a) illustrates the average condition number of the Galerkin-projected diffusion and advection operators as the dimension of the xKrylov subspaces increases across 12 outer iterations, confirming that the projected system gets more ill-conditioned as the rank increases. Figures 4 (b) - (d) show the GMRES residual histories for the ACS-preconditioned and unpreconditioned cases, corresponding to the second, fourth, and twelfth outer xKrylov iterations. These correspond to xKrylov subspace dimensions of $r=37$, $r=73$, and $r=217$, respectively. As $r$ grows, unpreconditioned GMRES converges progressively slower, as expected. In contrast, the ACS-preconditioner maintains rapid convergence rates independent of subspace size $r$.
To assess the relative cost of the preconditioner, we measure the ratio of the wall-clock timings for a single GMRES iteration with and without preconditioning for the different solution ranks considered in Fig. 4, yielding $1.64$, $2.09$, and $1.85$, respectively. A preconditioned GMRES iteration is only about twice as expensive as an unpreconditioned one, but results in a dramatic improvement on GMRES’ convergence rate.
Figure 4: Example 4.1. (a) Average condition number for Galerkin-projected advection and diffusion operators. (b)-(d) Log-log plots of GMRES residuals versus iterations, with and without preconditioning for various subspace sizes $r_{\kappa}=37$, $r_{\kappa}=73$, and $r_{\kappa}=217$, respectively. These cases correspond to the highlighted blue data points in (a). Simulations are conducted for a single time step with $\Delta t=0.01$ and a grid of $N_{x}=N_{y}=N_{z}=1000$.
Verification of the Computational Complexity and Memory Scalings: We verify the $\mathcal{O}(N)$ computational complexity and memory storage claims next. We vary $N$ from $100$ to $30000$ and the time step is fixed at $\Delta t=10^{-3}$. Figure 5(a) displays a profiling study of the algorithm, depicting the computational time of various components vs. $N$ obtained for $10$ time steps. We observe that only the xKrylov component scales with mesh refinement $N$, and does so linearly. Other components do not scale with $N$, consistently with the rank of the solution being bounded. Furthermore, the inner solve remains the dominant computational cost of the algorithm until, for grids of size $N\approx 20{,}000$, the xKrylov routine begins to dominate. This, in turn, highlights the importance of an efficient reduced-system solution strategy. Figure 5(b) displays the computational time required for the preconditioned GMRES-ACS solver vs. rank for a single time step, using BE with $N=10^{4}$ and $\epsilon_{\texttt{GMRES}}=10^{-10}$. For this test, the xKrylov basis was allowed to continue to grow boundlessly, albeit with truncation applied at each augmentation step. The simulation time required for the GMRES solve was recorded at each basis augmentation. The expected $\mathcal{O}(r^{4})$ scaling in 3D is found. These studies verify that the complexity scaling of the approach is $\mathcal{O}(Nr^{2})+\mathcal{O}(r^{4})$ in 3D. Figure 5(c) depicts the results of a memory scaling study demonstrating that the memory requirements also scale as $\mathcal{O}(N)$. The memory usage reported corresponds to the total size of all variables allocated within the solver function, as measured by MATLAB’s whos command at the end of the function call. A similar trend is expected (and observed in our simulations) in terms of memory storage: the inner tensor dominates the storage for coarser grids, while storing the xKrylov subspace dominates for finer grids.
Figure 5: Example 4.1. (a) Profiling study depicting the computational cost of various components of the algorithm, demonstrating overall ${\cal O}(N)$ computational complexity of DIRK3 adaptive-rank integrator in 3D. (b) Demonstration of ${\cal O}(r^{4})$ computational complexity of the GMRES-ACS solver in 3D. For this test only, we allow the rank to grow indefinitely by setting the outer residual tolerance, $\epsilon_{\texttt{tol}}$, to machine precision. (c) Demonstration of ${\cal O}(N)$ memory storage complexity of DIRK3 adaptive-rank integrator in 3D. The execution times were recorded using MATLAB’s timeit function. The memory usage was recorded using MATLAB’s whos command at the end of each time step, and the values reported correspond to the average memory usage across all time steps in megabytes.
To provide additional perspective on our results, we compare the performance of the adaptive-rank integrator for this example with a hypothetical optimal full-rank multigrid solver, which at best would scale as $\mathcal{O}(N^{d}\log N)$ [21]. Assuming that the low-rank and multigrid solvers are at parity for $N=100$ (Appendix E), Table 1 reports the theoretical wall-clock-time speedup of the adaptive-rank integrator vs. the full-rank multigrid solver. The table suggests vast efficiency gains from our approach vs. optimal multigrid for this problem, scaling as 1D with grid refinement and effectively mitigating the curse of dimensionality.
Table 1: Theoretical speedup of the adaptive-rank solver versus a hypothetical optimal multigrid solver for various 1D grid resolutions NN in a 3D domain.
1D grid size NN Wall-clock-time speedup in 3D
$1000$ $1446.1$
$4000$ $99229.3$
$10000$ $1418193.2$
Example 4.2: 3D Swirling Deformation Flow
We consider next a 3D swirling deformation flow:
$f_{t}+\left[-4(1-x^{2})yzf\right]_{x}+\left[-4x(1-y^{2})zf\right]_{y}+\left[8xy(1-z^{2})f\right]_{z}=\nu\Delta f,\qquad x,y,z\in[-1,1].$ (35)
We set the diffusion coefficient $\nu=1/500$. The flow is solenoidal, and therefore a pure-vorticity flow. The initial condition is a bi-Gaussian function given by:
$f(x,y,0)$ $=0.5\exp\left(-\frac{1}{2\sigma^{2}}\left[(x-0.3)^{2}+(y-0.35)^{2}+(z-0.3)^{2}\right]\right)$
$+0.8\exp\left(-\frac{1}{2\sigma^{2}}\left[(x+0.5)^{2}+(y+0.5)^{2}+(z+0.5)^{2}\right]\right),$
where $\sigma=0.15$. We choose $N_{x}=N_{y}=N_{z}=1000$. The outer basis residual tolerance is $\epsilon_{\texttt{tol}}=10^{-5}$, the basis truncation threshold is $\epsilon_{\kappa}=10^{-5}$, the SVD truncation threshold is $\epsilon=10^{-5}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-10}$.
This example demonstrates the algorithm’s ability to adapt to dynamic rank evolution. The algorithm performs nine xKrylov augmentations at the initial time step, two augmentations at the third and seventh time steps, and a single augmentation otherwise. Figure 6 shows rank evolution. The solution rank rises sharply in the $x$- and $y$-directions [Figures 6 (a)-(b)]. The solution rank in the $z$-direction evolves more gently [Figure 6 (c)] and is roughly one-third of that in the $x$ and $y$ directions. Figures 6 (d)-(e) showcase how the solution deforms according to the prescribed advection coefficients and diffuses slowly over time. Figure 7 shows snapshots of the solution in the $x\text{--}z$ plane at y=0.5y=-0.5. As the flow-driven deformation increases the complexity of the solution structure, the rank grows to capture these features; when the diffusion process starts to dominate, the rank decreases as the solution smooths. This example shows that the low-rank assumption is not intrinsic to the algorithm. Even when the rank becomes moderately large, the algorithm remains capable of capturing the underlying dynamics effectively. Naturally, the benefit of a low-rank approximation decreases as the solution approaches a full-rank structure.
From a computational perspective, the per-time-step cost breakdown is as follows: xKrylov construction and orthogonalization (10%), ACS-GMRES inner solve (47.5%), and residual evaluation (42.5%). The relatively small contribution of the cost of xKrylov construction is partly a consequence of the truncation mechanism, together with the effectiveness of the chosen projection spaces, which collectively keep the approximation spaces compact. On average, in the $x$-$y$ directions, the truncation reduces the subspace dimension by a factor of about $4.5$, from roughly $450$ to $100$. In the $z$ direction, the reduction amounts to a factor of approximately $3$, from about $150$ to $50$. On the other hand, the relatively large contribution of the residual evaluation to the simulation time originates in the poor scaling of the computational complexity with the number of GSE terms (eight in this example, including the RHS) and suggests that a more economical stopping criterion (as discussed in Section 3.4) should be used.
Figure 6: Example 4.2. 3D Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=14$ on a $1000\times 1000\times 1000$ grid. (a)–(c) Evolution of rank, xKrylov subspace before and after truncation, and number of xKrylov iterations over time for the $x$, $y$, and $z$ directions, respectively. (d)–(e) Solution profiles at $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$. The red dots in (a)–(c) mark these time instants and correspond to the profiles shown in (d)–(e). The final simulation time is $T_{f}=10$.
Figure 7: Example 4.2. 3D Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=14$ on a $1000\times 1000\times 1000$ grid. (a)–(f) snapshots of the solution in the $x\text{--}z$ plane at y=0.5y=-0.5, corresponding to times $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$, respectively.
Example 4.3: A 2D Steady-State Balanced-Advection-Diffusion Problem
The last example prescribes diffusion and advection coefficients that exactly balance each other, resulting in a non-trivial steady-state solution. This test demonstrates the ability of the algorithm to take large timesteps stably to solve directly for the steady state with expected spatial asymptotic accuracy.
We seek an equilibrium solution satisfying the steady-state condition:
$0=\nabla\cdot\left(\bm{\mathit{\Phi}}\cdot\nabla f_{\text{eq}}-\bm{\mathit{\Sigma}}f_{\text{eq}}\right),$ (36)
a solution of which is:
$\bm{\mathit{\Sigma}}=\bm{\mathit{\Phi}}\cdot\nabla\log(|f_{\text{eq}}|).$
This equation defines the steady-state advective flow $\bm{\mathit{\Sigma}}$ for given diffusion coefficient $\bm{\mathit{\Phi}}$ and steady-state solution $f_{\text{eq}}$. We construct a steady-state solution in the domain $0\leq x,y\leq 1$ to be (up to an arbitrary constant):
$f_{\text{eq}}(x,y)=\left(x(1-x)\right)^{2}\left(y(1-y)\right)^{2},$
which satisfies homogeneous Dirichlet boundary conditions. The diffusion coefficients are defined as:
$\phi^{x}(x,y)=\phi^{y}(x,y)$ $=\phi^{1}(x)\phi^{2}(y),\hskip 18.49988pt\phi^{1}(x)=x^{2}(1-x)^{2},\hskip 18.49988pt\phi^{2}(y)=y^{2}(1-y)^{2}.$
The variable advection coefficients are $\sigma^{x}(x,y)=\sigma_{1}^{1}(x)\sigma_{1}^{2}(y)$, $\sigma^{y}(x,y)=\sigma_{2}^{1}(x)\sigma_{2}^{2}(y),$ where:
$\sigma_{1}^{1}(x)$ $=2x(1-3x+2x^{2}),\hskip 18.49988pt\sigma_{1}^{2}(y)=y^{2}(1-y)^{2},\hskip 18.49988pt\sigma_{2}^{1}(x)=x^{2}(1-x)^{2},\hskip 18.49988pt\sigma_{2}^{2}(y)=2y(1-3y+2y^{2}).$
We select the following rank-1 initial condition: $f(x,y,0)=\left|\sin(2\pi x)\sin(2\pi y)\right|$. The outer basis residual tolerance is $\epsilon_{\texttt{tol}}=10^{-3}$, the basis truncation threshold is $\epsilon_{\kappa}=10^{-8}$, the SVD truncation threshold is $\epsilon=10^{-8}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-8}$.
Figure 8 shows the error of the numerical steady-state solution relative to the prescribed analytical steady state as a function of mesh size. The solution is obtained using ten time steps of size $\Delta t=1000$. The advection CFL number $\lambda_{A}$ ranges from $1,190$ to $36,076$ as the mesh is refined. The error is obtained by comparing the numerical and analytical solutions, with the numerical solution rescaled so that its normalization constant is the same as the analytical one. The results confirm second-order spatial convergence and the ability of the algorithm to solve directly for the correct steady state.
Figure 8: Example 4.3. Log-log plot of error in the numerical steady-state solution compared to the prescribed analytical steady state as a function of mesh size. The advection CFL number, $\lambda_{A}$, ranges from $1,190$ to $36,076$. The plot shows second-order spatial convergence in the steady state.
5 Conclusion
We propose a computationally efficient adaptive-rank implicit algorithm for solving time-dependent advection-diffusion partial differential equations with variable coefficients. We formulate the corresponding implicit system as a GSE, for which we follow a projection type approach, namely: constructing dimension-wise basis from xKrylov subspaces, projecting the original matrix equation onto a reduced system, and solving for the coefficient matrix. Our approach innovates in the way we construct the Krylov subspaces, and in the solution of the reduced system, which leverages GMRES effectively preconditioned with an effective Averaged-Coefficient Sylvester (ACS) approximation. The approach achieves a 1D-like computational complexity (of $\mathcal{O}(Nr^{2})+\mathcal{O}(r^{d+1})$, with $r$ being the rank of the solution during Krylov iteration process and $d$ the dimensionality) and memory storage [of $\mathcal{O}(Nr)+\mathcal{O}(r^{d})$], effectively mitigating the curse of dimensionality when the solution rank is bounded and making it comparable in efficiency to the constant-coefficient case [14]. Such complexity scaling is numerically verified with extensive tests, with the maximal rank $r$ found to be independent of the 1D grid size $N$ for practical grid resolutions and convergence tolerances. Future research directions include extending the algorithm to higher dimensions and nonlinear PDEs.
Acknowledgements
This work was partially supported by the Multifaceted Mathematics Integrated Capability Centers (MMICCs) program of DOE Office of Applied Scientific Computing Research (ASCR). Los Alamos National Laboratory (LANL) is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). J.Q. and H.E. were supported by NSF grant NSF-DMS-2111253. J.Q. was also supported by Air Force Office of Scientific Research FA9550-22-1-0390, Department of Energy DE-SC0023164 and Air Force Office of Scientific Research (AFOSR) FA9550-24-1-0254 via the Multidisciplinary University Research Initiatives (MURI) Program. H.E. was also supported by MMICCs at LANL during the summer of 2024.
Appendix A Discretization of the Diffusion Operator
We derive the matrix formulation of the discretization of the diffusion operator for the two-dimensional advection-diffusion equation with separable diffusion coefficients. We consider the discretization of the diffusion term in the $x$-direction, with other directions treated similarly:
$\frac{\partial}{\partial x}\left(\phi^{x}(x,y)\frac{\partial u(x,y)}{\partial x}\right),$ (37)
where the diffusion coefficient is separable:
$\phi^{x}(x,y)=\phi^{1}(x)\phi^{2}(y).$ (38)
We discretize this term using a 3-point stencil on a uniform grid with spacing $\Delta x$ in the $x$-direction. Let $u_{i,j}\approx u(x_{i},y_{j})$, with $i=0,\cdots,N_{x}-1$ and $j=0,\cdots,N_{y}-1$. We approximate the spatial derivative using central differences. The discrete approximation at grid point $(i,j)$ is:
$\left[\frac{\partial}{\partial x}\left(\phi^{x}\frac{\partial u}{\partial x}\right)\right]_{i,j}\approx\frac{1}{\Delta x^{2}}\left[\phi^{1}_{i+\frac{1}{2}}(u_{i+1,j}-u_{i,j})-\phi^{1}_{i-\frac{1}{2}}(u_{i,j}-u_{i-1,j})\right]\phi^{2}_{j},$ (39)
with $\phi^{1}_{i+\frac{1}{2}}\phi^{2}_{j}=\phi^{1}(x_{i+\frac{1}{2}})\phi^{2}(y_{j}).$ To express the discretization in matrix form, we define:
  • $\mathbf{F}$: matrix of solution values, $\mathbf{F}]_{i,j}=f_{i,j}$.
  • $\Phi^{1}_{+}$: diagonal matrix with elements $\left[\Phi^{1}_{+}\right]_{i,i}=\phi^{1}_{i+\frac{1}{2}}$.
  • $\Phi^{1}_{-}$: diagonal matrix with elements $\left[\Phi^{1}_{-}\right]_{i,i}=\phi^{1}_{i-\frac{1}{2}}$.
  • $\Phi_{2}$: diagonal matrix with elements $\left[\Phi_{2}\right]_{j,j}=\phi^{2}_{j}$.
We also define the forward and backward difference matrices in the $x$-direction for the interior nodes $i=1,\cdots,N_{x}-2$. The nodes $i=0$, $i=N_{x}-1$ rest at the boundary and enforce homogeneous Dirichlet boundary conditions (other boundary conditions can be readily considered):
$D_{+}=\frac{1}{\Delta x}\begin{pmatrix}-1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\\ 0&\cdots&0&0&-1\end{pmatrix},D_{-}=\frac{1}{\Delta x}\begin{pmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\end{pmatrix}$. (40)
Using these definitions, the discrete diffusion term in Eq. (39) can be represented in matrix form as:
$\left(\Phi^{1}_{+}D_{+}-\Phi^{1}_{-}D_{-}\right)\mathbf{F}{\Phi_{2}^{\top}}.$ (41)
We note that the operator multiplying $\mathbf{F}$ from the left is analogous to $D^{\Phi}_{xx}$ in (8). The discretization in the $y$-direction follows the same strategy.
Appendix B Discretization of the Advection Operator
We derive the matrix formulation of the discretization of the advection operator for the two-dimensional advection-diffusion equation with variable, separable advection coefficients. As before, we focus on the discretization of the advection term in the $x$-direction:
$\frac{\partial}{\partial x}\left(\sigma^{x}(x,y)u(x,y)\right),$ (42)
where the advection coefficient is separable:
$\sigma^{x}(x,y)=\sigma^{1}(x)\sigma^{2}(y).$ (43)
We discretize this term on a uniform grid with spacing $\Delta x$ in the $x$-direction and $\Delta y$ in the $y$-direction. To preserve linearity of the discretization without introducing too much numerical dissipation, we employ a central difference scheme to approximate the spatial derivative. This discretization is monotonicity-preserving when the following condition is satisfied:
$\Delta x<2\frac{\phi_{\text{max}}}{\sigma_{\text{max}}},$ (44)
where $\sigma_{\text{max}}$ is the maximum magnitude of the advection velocity and $\phi_{\text{max}}$ is the maximum diffusion coefficient. This condition is practical in our formulation due to the very favorable scaling of the low-rank representation with mesh size and rank.
The discrete approximation at the grid point $(i,j)$ is:
$\left[\frac{\partial}{\partial x}\left(\sigma^{x}(x,y)u(x,y)\right)\right]_{i,j}$ $\approx\frac{1}{2\Delta x}\left[\sigma^{1}_{i+\frac{1}{2}}(u_{i+1,j}+u_{i,j})-\sigma^{1}_{i-\frac{1}{2}}(u_{i,j}+u_{i-1,j})\right]\sigma^{2}_{j},$ (45)
where $\sigma^{1}_{i+\frac{1}{2}}=\sigma^{1}(x_{i+\frac{1}{2}})$ is the advection coefficient evaluated at the midpoint between $x_{i}$ and $x_{i+1}$, and similarly for $\sigma^{1}_{i-\frac{1}{2}}$, and $\sigma^{2}_{j}$ is evaluated at $y_{j}$. To express the discretization in matrix form, we define:
  • $\mathbf{F}$: matrix of solution values, $\mathbf{F}]_{i,j}=f_{i,j}$.
  • $\Sigma^{1}_{+}$ and $\Sigma^{1}_{-}$: diagonal matrices with elements $\left[\Sigma^{1}_{+}\right]_{i,i}=\sigma^{1}_{i+\frac{1}{2}}$ and $\left[\Sigma^{1}_{-}\right]_{i,i}=\sigma^{1}_{i-\frac{1}{2}}$.
  • $\Sigma_{2}$: diagonal matrix with elements $\left[\Sigma_{2}\right]_{j,j}=\sigma^{2}_{j}$.
We also define the difference matrices in the $x$-direction:
$S_{+}=\frac{1}{2}\begin{pmatrix}1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\\ 0&\cdots&0&0&1\end{pmatrix},S_{-}=\frac{1}{2}\begin{pmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\end{pmatrix}$. (46)
Using these definitions, the discrete advection term in the $x$-direction can be represented as:
$\frac{1}{\Delta x}\left(\Sigma^{1}_{+}S_{+}-\Sigma^{1}_{-}S_{-}\right)\mathbf{F}{\Sigma_{2}^{\top}}.$ (47)
We note that the operator multiplying $\mathbf{F}$ from the left is analogous to $D^{\Sigma}_{x}$ in Equation (8). The discretization in the $y$-direction follows the same strategy.
Appendix C High-Order Temporal Discretization
We extend the backward Euler implicit integrator to higher-order DIRK time discretizations for the advection-diffusion operator, as compactly written in (10).
C.1 DIRK Scheme
DIRK methods are characterized by the Butcher tableau, Table 2.
Table 2: Butcher tableau for an $s$-stage DIRK scheme. Here, $s$ represents the number of stages in DIRK, $c_{i}$ represents the intermediate stage, $a_{ij}$ is a lower-triangular matrix with coefficients used to approximate the solution at each stage, and $b_{j}$ are the quadrature weights used to update the solution in the final step.
$c_{1}$ $a_{11}$ 0 $\ldots$ 0
$c_{2}$ $a_{21}$ $a_{22}$ $\ldots$ 0
$\vdots$ $\vdots$ $\vdots$ $\ddots$ $\vdots$
$c_{s}$ $a_{s1}$ $a_{s2}$ $\ldots$ $a_{ss}$
$b_{1}$ $b_{2}$ $\ldots$ $b_{s}$
We focus on stiffly accurate DIRK methods, where the coefficients satisfy $b_{i}=a_{si}$ for $i=1,\ldots,s$, ensuring that the final solution update corresponds to the solution at the last DIRK stage. The Butcher tables for DIRK2 and DIRK3 are presented in Tables 4 and 4, respectively.
Table 3: DIRK2 Butcher table, $\gamma=1-\frac{\sqrt{2}}{2}$.
$\gamma$ $\gamma$ 0
11 $1-\gamma$ $\gamma$
$1-\gamma$ $\gamma$
Table 4: DIRK3 Butcher table, $x=0.4358665215$.
$x$ $x$ 0 0
$\frac{1+x}{2}$ $\frac{1-x}{2}$ $x$ 0
11 $\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
$\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
Similarly to the backward Euler case, the matrix-based discretization of the advection-diffusion equation using the operator $\mathscr{L}$ defined in (10) results in a GSE that must be solved at each DIRK stage. The corresponding DIRK scheme for the matrix differential equation (10) from $t^{(0)}$ to $t^{(1)}$ is written as:
$\begin{split}\mathbf{F}^{(k)}&=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k}a_{k\ell}\mathbf{Y}_{\ell},\hskip 18.49988ptk=1,2,\ldots,s,\end{split}$ (48a)
$\begin{split}\mathbf{Y}_{k}&=\mathscr{L}(\mathbf{F}^{(k)};t^{(k)}),\hskip 18.49988ptt^{(k)}=t^{0}+c_{k}\Delta t,\hskip 18.49988ptk=1,2,\ldots,s,\end{split}$ (48b)
$\begin{split}\mathbf{F}_{1}&=\mathbf{F}^{(s)}=\mathbf{F}_{0}+\Delta t\sum\limits_{k=1}^{s}b_{k}\mathbf{Y}_{k}.\end{split}$ (48c)
Within each stage $k$ of the DIRK scheme, one must solve the residual equation given by:
${\bf R}^{(k)}_{\mathscr{L}}=\mathbf{F}^{(k)}-a_{kk}\Delta t\mathscr{L}(\mathbf{F}^{(k)})-\mathbf{B}^{(k)}=\mathbf{0},\qquad\text{with}\qquad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell},\qquad k=1,2,\ldots,s.$ (49)
Expanding the operator $\mathscr{L}$ in the residual equation above results in the following $k$-th stage GSE:
$\mathbf{F}^{(k)}-a_{k,k}\Delta t\left(T_{1}\mathbf{F}^{(k)}\Phi_{2}{{}^{\top}}+\Phi_{1}\mathbf{F}^{(k)}T_{2}^{\top}+T_{3}\mathbf{F}^{(k)}\Sigma_{2}^{\top}+\Sigma_{1}\mathbf{F}^{(k)}T_{4}^{\top}\right)=\mathbf{B}^{(k)},\qquad k=1,2,\ldots,s.$ (50)
For the DIRK update, we adopt the strategy outlined in [14], with minor modifications to improve efficiency. In particular, we: (i) construct the basis using the time-step of the first stage and retain it for subsequent stages, (ii) evolve the projected system for the $S_{1}$ matrix across the DIRK stages, and (iii) truncate and evaluate the residual to determine whether to admit the solution or return to step (i) to further augment the basis. Key modifications vs. [14] include truncating the solution before the residual computation, and only checking the residual at the end of the DIRK step. The reduced matrix is solved using preconditioned GMRES, as before.
C.2 Adaptive-Rank Strategy for High-Order Temporal Discretization via DIRK
The high-order DIRK method features a stage-by-stage generalized-Sylvester-equation solve. The strategy follows closely that of the backward Euler adaptive-rank treatment, i.e., we construct an approximate $\mathscr$$\tilde${L}} used for both the basis construction of $U_{1}$ and $V_{1}$ and the preconditioning of the reduced system to solve $S_{1}$.
At each stage $k$ of the DIRK scheme, the corresponding DIRK formulation for the matrix differential equation from $t^{(0)}$ to $t^{(1)}$ requires solving an SE for the approximated operator $\mathscr$$\tilde${L}} by solving the $k$th-stage residual:
$\mathbf{R}^{(k)}_{\tilde{\mathscr{L}}}=\mathbf{F}_{1}-a_{kk}\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}-\mathbf{B}^{(k)}=\mathbf{0},\qquad\text{with}\qquad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell}.$ (51)
Expanding this residual using $\tilde{\mathscr{L}}$, defined in (15), leads to the $k$th-stage SE:
$\underbrace{\left[\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\alpha_{2}T_{1}\right)}_{A_{1}}+\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\gamma_{2}T_{3}\right)}_{A_{3}}\right]\mathbf{F}^{(k)}}_{P_{1}}+\mathbf{F}^{(k)}\underbrace{\left[\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\alpha_{1}T_{2}\right)^{\top}}_{A^{\top}_{2}}+\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\gamma_{1}T_{4}\right)^{\top}}_{A^{\top}_{4}}\right]}_{P_{2}^{T}}=\mathbf{B}^{(k)}.$ (52)
Note that the operators $A_{1:4}$ are fixed throughout the stages since the DIRK method uses constant diagonal Butcher tableau coefficients.
We follow [14] to perform high-order integration. The key idea is to consider the basis fixed across DIRK stages, and only evolve the matrix of coefficients ${{\bf S}}_{1}$ through the stages. The procedure is as follows:
  1. Step 1.
    Prediction of Krylov basis functions: construct a set of orthonormal bases $U_{1}$ and $V_{1}$ from the Krylov-based low-rank implicit solver at the first DIRK stage, i.e. backward Euler with time stepping size $c_{1}\Delta t$.
  2. Step 2.
    For $k=1:s$ (per DIRK stage)
    1. (a)
      Solve reduced SE for ${\bf S}^{(k)}$ using preconditioned GMRES,
      ${\bf S}^{(k)}-\Delta t\left(\tilde{T}_{1}{\bf S}^{(k)}\tilde{\Phi}^{\top}_{2}+\tilde{\Phi}_{1}{\bf S}^{(k)}\tilde{T}_{2}^{\top}+\tilde{T}_{3}{\bf S}^{(k)}\tilde{\Sigma}^{\top}_{2}+\tilde{\Sigma}_{1}{\bf S}^{(k)}\tilde{T}_{4}^{\top}\right)=\tilde{{\bf B}}^{(k)},$ (53)
      with
      $\tilde{{\bf B}}^{(k)}=U_{1}^{\top}\mathbf{F}_{0}V_{1}+\Delta t\sum\limits_{\ell=1}^{k-1}{a_{k\ell}\tilde{Y_{\ell}}}$,
      Here $\tilde{Y_{\ell}}=U_{1}^{\top}{Y}_{\ell}V_{1}\in\mathbb{R}^{r_{x}\times r_{y}}$. The preconditioner used is given by
      ${\bf S}\mapsto\tilde{P}_{1}{\bf S}+{\bf S}\tilde{P}_{2},$ (54)
      where $\tilde{P}_{1}=U_{1}^{T}P_{1}U_{1}$ and $\tilde{P}_{2}=V_{1}^{T}P_{1}V_{1}$, and is valid for all stages since the basis constructed is fixed throught the DIRK stages by assumption. Solve for ${\bf S}^{(k)}$ and obtain the intermediate RK solutions as $U_{1}S^{(k)}V_{1}^{\top}$. To further improve computational efficiency, from (53), we have,
[…]
      leading to an efficient computation of $\tilde{{\bf B}}^{(k)}$:
[…]
      with $\tilde{{\bf B}}_{1}$ defined in (20). This avoids the need to evaluate the full size ${\bf B}^{(k)}$ in (51).
  3. Step 3.
    Perform truncation of evolved solution, i.e. $\{U_{1,\epsilon},S_{1,\epsilon},V_{1,\epsilon}^{\top}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}^{\top}\})$, with $S_{1}:=S_{1}^{(s)}$.
  4. Step 4.
    Efficiently evaluate the residual norm,
    $\left\|{{\bf R}^{(s)}_{\mathscr{L}}}\right\|=\left\|{R}_{U}{\text{diag}\left(-\tilde{B}^{(s)}_{1},S_{1,\epsilon},-a_{ss}\Delta t\left(I_{{\mathcal{R}}_{\text{ranks}}}\otimes S_{1,\epsilon}\right)\right)}{R}_{V}^{\top}\right\|,$
    where $R_{U}$ and $R_{V}$ are upper triangular matrices from a reduced Q-less QR decomposition as described in (24), which can be done once for all DIRK stages.
  5. Step 5.
    Compare the computed residual norm $\|{\bf R}^{(s)}_{\mathscr{L}}\|$ with the given error tolerance $\epsilon_{\texttt{tol}}$. If the tolerance is met, then admit the solution with ${\bf F}_{1}=U_{1,\epsilon}S_{1,\epsilon}V_{1,\epsilon}^{\top}$; otherwise, we go back to Step 1 to further augment the xKrylov subspaces.
C.3 DIRK Generalization to Multi-Rank Advection-Diffusion Coefficients
We now extend the formulation to the multi-rank case, as described in (6). Using a DIRK method for temporal discretization, the $k$-th stage GSE becomes:
$\mathbf{F}^{(k)}-a_{kk}\Delta t\left(\underbrace{\sum_{i=1}^{\ell_{x}}T_{1,i}\mathbf{F}^{(k)}{\Phi_{i}^{2,x}}^{\top}+\sum_{j=1}^{\ell_{y}}\Phi_{j}^{1,y}\mathbf{F}^{(k)}{T_{2,j}}^{\top}+\sum_{k=1}^{k_{x}}T_{3,k}\mathbf{F}^{(k)}{\Sigma_{k}^{2,x}}^{\top}+\sum_{l=1}^{k_{y}}\Sigma_{l}^{1,y}\mathbf{F}^{(k)}{T_{4,l}}^{\top}}_{\mathscr{L}(\mathbf{F}^{(k)})}\right)=\mathbf{B}^{(k)},\qquad k=1,2,\ldots,s$ (56)
where $\mathbf{B}^{(k)}$ is as defined in (51). The terms $T_{1:4,i}$ represent compositions of difference operators with diffusion and advection coefficients, similar to (9). Left subscripts 1 and 2 correspond to diffusion, while 3 and 4 correspond to advection in the $x$ and $y$ directions, respectively. The right subscript denotes the rank of the respective diffusion and advection coefficients. This equation generalizes (12), which is recovered for $\ell_{x}=\ell_{y}=k_{x}=k_{y}=1$ and equal diffusion and advection coefficients in both the $x$ and $y$ directions. It also generalizes the backward Euler special case with a single stage and $a_{1,1}=1$, yielding $\mathbf{B}^{(1)}=\mathbf{F}_{0}$, consistently with (12).
Using approximations in (62), we arrive at the $k$th stage approximated SE:
[…]
where
$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{3,k}}=\sum_{i=1}^{\ell_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-a_{kk}\Delta t\alpha_{x,i}T_{1,i}\right)}+\sum_{k=1}^{k_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-a_{kk}\Delta t\gamma_{x,k}T_{3,k}\right)},$
$P_{2}$ $=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{4,l}}=\sum_{j=1}^{\ell_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-a_{kk}\Delta t\alpha_{y,j}T_{2,j}\right)}+\sum_{l=1}^{k_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-a_{kk}\Delta t\gamma_{y,l}T_{4,l}\right)}.$
In order to obtain a reduced equation for the $k$th stage matrix of coefficients ${\bf S}^{(k)}$, we perform a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$ to obtain :
${\bf S}^{(k)}-a_{kk}\Delta_{t}\left(\sum_{i=1}^{\ell_{x}}\tilde{T}_{1,i}{\bf S}^{(k)}\tilde{\Phi}_{i}^{2,x\top}+\sum_{j=1}^{\ell_{y}}\tilde{\Phi}_{j}^{1,y}{\bf S}^{(k)}\tilde{T}_{2,j}^{\top}+\sum_{k=1}^{k_{x}}\tilde{T}_{3,k}{\bf S}^{(k)}\tilde{\Sigma}_{k}^{2,x\top}+\sum_{l=1}^{k_{y}}\tilde{\Sigma}_{l}^{1,y}{\bf S}^{(k)}\tilde{T}_{4,l}^{\top}\right)=\tilde{{\bf B}}^{(k)}$ (58)
where the projected operators are analogous to those defined in the rank-one case.
[…]
$\underbrace{\left[\sum_{i=1}^{\ell_{x}}{\tilde{A}_{1,i}}+\sum_{k=1}^{k_{x}}{\tilde{A}_{3,k}}\right]}_{\tilde{P}_{1}}\mathbf{S}^{(k)}+\mathbf{S}^{(k)}\underbrace{\left[\sum_{j=1}^{\ell_{y}}{\tilde{A}_{2,j}}+\sum_{l=1}^{k_{y}}{\tilde{A}_{4,l}}\right]}_{\tilde{P}_{2}^{\top}}=\tilde{{\bf B}}^{(k)}$ (59)
Using the operators $\tilde{P}_{1}$ and $\tilde{P}_{2}$, we define the ACS-preconditioner mapping:
${\bf S}\mapsto\tilde{P}_{1}{\bf S}+{\bf S}\tilde{P}_{2}^{T},$ (60)
Algorithm 5 solves the time-dependent variable advection-diffusion equation using DIRK methods.
Input: Initial condition matrices $U_{0}$, $V_{0}$, $S_{0}$;
Operators $P_{1}$, $P_{2}$, $\{T_{1,i}\}_{i=1}^{\ell_{x}}$, $\{T_{2,j}\}_{j=1}^{\ell_{y}}$, $\{T_{3,k}\}_{k=1}^{k_{x}}$, $\{T_{4,l}\}_{l=1}^{k_{y}}$;
Butcher table $\{a_{ij}\}$;
Time step size $\Delta t$;
Tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, ${\epsilon}_{\texttt{tol}}$;
Output: Updated bases $U_{1}$, $V_{1}$;
Truncated singular values $S_{1}$;
1
21exCompute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eq. (57);
3 Compute $l=2+\ell_{x}+\ell_{y}+k_{x}+k_{y}$;
4 Set $U^{(i)}=U_{0}$ and $V^{(i)}=V_{0}$ for $i=1,\dots,l$;
5 Set $U_{1}=U_{0}\qquad V_{1}=V_{0}$;
6 while not converged do
7    // Step 1.
8    Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
9    $U_{1},\{U^{(1)}\cdots U^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}}\};\{U^{(1)}\cdots U^{(l)}\};U_{1};\epsilon_{\kappa}\right)$;
10    $V_{1},\{V^{(1)}\cdots V^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}}\};\{V^{(1)}\cdots V^{(l)}\};V_{1};\epsilon_{\kappa}\right)$;
11    // Step 2.
12    for $k=1$ to $s$ do
[…]
14        Solve the reduced equation (59) for ${\bf S}^{(k)}$ using GMRES with ACS preconditioner (60) to tolerance $\epsilon_{\texttt{GMRES}}$;
15        Compute and store $\frac{1}{a_{kk}}\left({\bf S}^{(k)}-\tilde{{\bf B}}^{(k)}\right)$ and proceed to the next stage;
16   Truncate $\{U_{1,\epsilon},S_{1,\epsilon},V_{1,\epsilon}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S^{(s)},V_{1}\})$ to tolerance $\epsilon$;
17    // Step 3.
18    Compute $\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
19    $\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
20    // Step 4.
21    Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(-\tilde{B}^{(s)},S_{1,\epsilon},-\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1,\epsilon}\right)\right)R_{V}^{\top}\right\|$;
22
23   if $\|\mathbf{R}\|/\|\mathbf{F}_{0}\|\geq\epsilon_{\texttt{tol}}$ then
24        Return to Step 1 to augment bases further;
25
26   else
27        Set $U_{1}\leftarrow U_{1,\epsilon}$; $U_{1}\leftarrow V_{1,\epsilon}$; $S_{1}=S_{1,\epsilon}$;
28        Break;
29        Exit loop;
30
31
Algorithm 5 s-Stages Adaptive-Rank DIRK Integrator for the GSE.
Appendix D Three-Dimensional Formulation
We outline the approach to 3D next. In 3D, we replace the SVD algorithm with a tensor Tucker decomposition, specifically tailored for the three (and higher) dimension tensor-product grids.
D.1 Tucker Tensor Equations
We discretize the 3D time-dependent advection-diffusion equation on a three-dimensional tensor-product spatial grid with $N_{1}$, $N_{2}$, and $N_{3}$ grid points in the $x$, $y$, and $z$ directions, respectively. Here and in the following, we let ${\bf F}\in\mathbb{R}^{N_{1}\times N_{2}\times N_{3}}$ denote a three-dimensional tensor. The mode-$n$ product of $\bf F$} with a matrix $A_{n}\in\mathbb{R}^{I\times N_{n}}$, for $n=1,2,3$, is defined as:
${\bf F}\times_{1}A_{1}\in\mathbb{R}^{I\times N_{2}\times N_{3}},\,{\bf F}\times_{2}A_{2}\in\mathbb{R}^{N_{1}\times I\times N_{3}},\,{\bf F}\times_{3}A_{3}\in\mathbb{R}^{N_{1}\times N_{2}\times I}.$
The resulting tensor in each case has its $n$-th dimension replaced by the number of rows of $A_{n}$. The entries of the mode-$n$ product are computed as:
$({\bf F}\times_{n}A_{n})_{i_{1},\ldots,i_{n-1},j,i_{n+1},\ldots,i_{3}}=\sum_{i_{n}=1}^{N_{n}}{\bf F}_{i_{1},\ldots,i_{n},\ldots,i_{3}}\cdot(A_{n})_{j,i_{n}}.$
For a more thorough review of tensor operations, we refer the reader to [23]. The discretization of the 3D advection-diffusion equation using finite-differences in space combined with the backward Euler method in time, yields the following GSE:
${\bf R}_{\mathscr{L}}=\mathbf{F}_{1}-\Delta t\mathscr{L}({\bf F}_{1})-\mathbf{F}_{0}=\mathbf{0},$ (61)
where ${\bf F}_{0}$ denotes the initial tensor condition at time $t_{0}$, and ${\bf F}_{1}$ is the sought tensor solution at time $t_{1}$, and,
$\mathscr{L}:\mathbf{F}\mapsto$ $\sum_{i=1}^{\ell_{x}}\mathbf{F}\times_{1}T_{1,i}\times_{2}\Phi_{i}^{2,x}\times_{3}\Phi_{i}^{3,x}+\sum_{i=1}^{\ell_{y}}\mathbf{F}\times_{1}\Phi_{i}^{1,y}\times_{2}T_{2,i}\times_{3}\Phi_{i}^{3,y}+\sum_{i=1}^{\ell_{z}}\mathbf{F}\times_{1}\Phi_{i}^{1,z}\times_{2}\Phi_{i}^{2,z}\times_{3}T_{3,i}$
$+$ $\sum_{i=1}^{k_{x}}\mathbf{F}\times_{1}T_{4,i}\times_{2}\Sigma_{i}^{2,x}\times_{3}\Sigma_{i}^{3,x}+\sum_{i=1}^{k_{y}}\mathbf{F}\times_{1}\Sigma_{i}^{1,y}\times_{2}T_{5,i}\times_{3}\Sigma_{i}^{3,y}+\sum_{i=1}^{k_{z}}\mathbf{F}\times_{1}\Sigma_{i}^{1,z}\times_{2}\Sigma_{i}^{2,z}\times_{3}T_{6,i}.$
The operators $T$ correspond to the composition of advection and diffusion difference operators with diagonal matrices, analogous to those defined in Section 2. Similarly, the operators $\Phi$ and $\Sigma$ are analogous to the diagonal matrices for diffusion and advection coefficients, respectively, used for pointwise multiplication, as defined in Section 2. To obtain an averaged (approximate) mapping for preconditioning purposes, we define the average coefficients:
$\alpha_{i}^{2,x}$ $=\text{avg}(\Phi_{i}^{2,x}),\,\alpha_{i}^{3,x}$ $=\text{avg}(\Phi_{i}^{3,x}),\,\alpha_{i}^{1,y}$ $=\text{avg}(\Phi_{i}^{1,y}),\,\alpha_{i}^{3,y}$ $=\text{avg}(\Phi_{i}^{3,y}),\,\alpha_{i}^{1,z}$ $=\text{avg}(\Phi_{i}^{1,z}),\,\alpha_{i}^{2,z}$ $=\text{avg}(\Phi_{i}^{2,z}),$ (62)
$\gamma_{i}^{2,x}$ $=\text{avg}(\Sigma_{i}^{2,x}),\,\gamma_{i}^{3,x}$ $=\text{avg}(\Sigma_{i}^{3,x}),\,\gamma_{i}^{1,y}$ $=\text{avg}(\Sigma_{i}^{1,y}),\,\gamma_{i}^{3,y}$ $=\text{avg}(\Sigma_{i}^{3,y}),\,\gamma_{i}^{1,z}$ $=\text{avg}(\Sigma_{i}^{1,z}),\,\gamma_{i}^{2,z}$ $=\text{avg}(\Sigma_{i}^{2,z}).$
which yields the three-dimensional approximate operator $\tilde{\mathscr{L}}$ as:
[…]
$+\mathbf{F}\times_{2}\bigg(\sum_{j=1}^{\ell_{y}}\alpha_{j}^{1,y}\alpha_{j}^{3,y}T_{2,j}+\sum_{l=1}^{k_{y}}\gamma_{l}^{1,y}\gamma_{l}^{3,y}T_{5,l}\bigg)$
$+\mathbf{F}\times_{3}\bigg(\sum_{k=1}^{\ell_{z}}\alpha_{k}^{1,z}\alpha_{k}^{2,z}T_{3,k}+\sum_{l=1}^{k_{z}}\gamma_{l}^{1,z}\gamma_{l}^{2,z}T_{6,l}\bigg).$ (63)
Using this approximate operator, and letting ${\mathcal{R}_{\text{ranks}}}=r_{x}+r_{y}+r_{z}+\ell_{x}+\ell_{y}+\ell_{z}$, we then define the averaged three-dimensional approximate SE of the form:
${\bf F}_{1}\times_{1}P_{1}+{\bf F}_{1}\times_{2}P_{2}+{\bf F}_{1}\times_{3}P_{3}={\bf F}_{0},$ (64)
with
$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\alpha_{i}^{2,x}\alpha_{i}^{3,x}T_{1,i}\right)}_{{A_{1,i}}}+\sum_{k=1}^{k_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\gamma_{k}^{2,x}\gamma_{k}^{3,x}T_{4,k}\right)}_{{A_{4,k}}}=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{4,k}},$
$P_{2}$ $=\sum_{j=1}^{\ell_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\alpha_{j}^{1,y}\alpha_{j}^{3,y}T_{2,j}\right)}_{A_{2,i}}+\sum_{l=1}^{k_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{l}^{1,y}\gamma_{l}^{3,y}T_{5,l}\right)}_{A_{5,i}}=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{5,l}},$
$P_{3}$ $=\sum_{j=1}^{\ell_{z}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{3}}-\Delta t\alpha_{k}^{1,z}\alpha_{k}^{2,z}T_{3,k}\right)}_{A_{3,i}}+\sum_{l=1}^{k_{z}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{l}^{1,z}\gamma_{l}^{2,z}T_{6,l}\right)}_{A_{6,i}}=\sum_{j=1}^{\ell_{z}}{A_{3,j}}+\sum_{l=1}^{k_{z}}{A_{6,l}}.$
These one-dimensional matrix operators, analogous to those in (26), are used (i) to construct xKrylov subspaces for the Tucker basis factors and (ii) to precondition the solution of the three-dimensional GSE arising from the Galerkin condition on the residual in Eq. (61).
D.2 GMRES-ACS Solution of the Reduced GSE in 3D
We assume the existence of a low-rank Tucker factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}\times N_{3}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}\times N_{3}}$ at time $t^{(1)}=t^{(0)}+\Delta t$, also in the low-rank Tucker format. This is represented as:
$\mathbf{F}_{0}=S_{0}\times_{1}U_{0}\times_{2}V_{0}\times_{3}W_{0}\xrightarrow{\text{evolve in time}}\mathbf{F}_{1}=S_{1}\times_{1}U_{1}\times_{2}V_{1}\times_{3}W_{1},$ (65)
where $U_{0}\in\mathbb{R}^{N_{1}\times r^{x}_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r^{y}_{0}}$, $W_{0}\in\mathbb{R}^{N_{3}\times r^{z}_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r^{x}_{1}}$, $V_{1}\in\mathbb{R}^{N_{2}\times r^{y}_{1}}$, and $W_{1}\in\mathbb{R}^{N_{3}\times r^{z}_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r^{x}_{0}\times r^{y}_{0}\times r^{z}_{0}}$ and $S_{1}\in\mathbb{R}^{r^{x}_{1}\times r^{y}_{1}\times r^{z}_{1}}$ are core tensors at the corresponding times. The bases $U_{1}$, $V_{1}$, and $W_{1}$ are constructed from orthonormalization of dimension-wise xKrylov, as discussed in 3.2
Using a Galerkin projection on the residual, $\mathbf{R}_{\mathscr{L}}\times_{1}U_{1}^{\top}\times_{2}V_{1}^{\top}\times_{3}W_{1}^{\top}$, we obtain a reduced-size GSE for the core tensor ${\bf S}_{1}$ as follows:
$\mathbf{S}_{1}-\Delta t\mathscr{L}(\mathbf{S}_{1})\times_{1}U_{1}^{\top}\times_{2}V_{1}^{\top}\times_{3}W_{1}^{\top}=\mathbf{S}_{0}\times_{1}U_{1}^{\top}U_{0}\times_{2}V_{1}^{\top}V_{0}\times_{3}W_{1}^{\top}W_{0}.$ (66)
This projected system has dimensions $r_{x}\times r_{y}\times r_{z}$, where $r_{x}$, $r_{y}$, and $r_{z}$ denote the dimensions of the truncated xKrylov subspaces in the $x$, $y$, and $z$ directions, respectively, and will be solved using the GMRES-ACS preconditioner.
Eq. (66) is solved iteratively using GMRES, with $\mathscr{L}$-operator application cast as a series of tensor-matrix multiplications, along with the ACS preconditioning operator given by:
$\mathcal{P}:\mathbf{S}\mapsto\mathbf{S}\times_{1}\tilde{P}_{1}+\mathbf{S}\times_{2}\tilde{P}_{2}+\mathbf{S}\times_{3}\tilde{P}_{3},$ (67)
where $\tilde{P}_{i}=P_{i}\times_{1}U_{1}^{\top}\times_{2}V_{1}^{\top}\times_{3}W_{1}^{\top}$ denotes the projection of operators in Eq. (64), analogous to those introduced in Section 3. This operator is used to precondition each GMRES iteration of the inner system solve. The solution of the projected SE has a computational cost of $\mathcal{O}(r^{4})$ [9, 44]. The tensor-times-matrix operations, executed using the ttm command in the Tensor Toolbox, incur a cost of $\mathcal{O}(r^{4})$. The total complexity in three dimensions is therefore $\mathcal{O}(Nr^{2}+r^{4})$.
Appendix E Technical Details of the Hypothetical Multigrid Comparison
We provide here a brief explanation of the assumptions involved in the performance comparison between a hypothetical optimally scaling full-rank multigrid method and the adaptive-rank approach. Optimal multigrid methods exhibit a computational scaling of $\mathcal{O}(N^{d}\log N)$ [21]. Let us denote the wall-clock time (WCT) of a multigrid method as:
$WCT_{\mathrm{MG}}=C_{\mathrm{MG}}N^{d}\log N.$
From Fig. 1-a, the wall-clock time of the adaptive-rank method may be written as:
$WCT_{\mathrm{LR}}=C_{\mathrm{LR}}(0.0023\,N+55.2734).$
The speedup of the adaptive-rank method vs. multigrid is given by the ratio:
$\frac{WCT_{\mathrm{MG}}}{WCT_{\mathrm{LR}}}=C_{\mathrm{WCT}}\frac{N^{3}\log N}{0.0023\,N+55.2734},$ (68)
where $C_{\mathrm{WCT}}=1.2\times 10^{-5}$ is found assuming that both methods achieve comparable wall-clock times for $N=100$. This formula is used to populate Table 1.